### v1

# Gauge Symmetries, Contact Reduction, and Singular Field Theories

Abstract
The reduction of dynamical systems which are invariant under changes of global scale is well-understood, for classical theories of particles, and fields. The excision of the superfluous degree of freedom describing such a scale leads to a dynamically-equivalent theory, which is frictional in nature. In this article, we extend the formalism to physical models, of both particles and fields, described by singular Lagrangians. Our treatment of classical field theory is based on the manifestly covariant Hamilton De-Donder Weyl formalism, in which the Lagrangian density is introduced as a bundle morphism on the pre-multisymplectic velocity phase space $J^{1}E$. The results obtained are subsequently applied to a number of physically-motivated examples, as well as a discussion presented on the implications of our work for classical General Relativity.
I Introduction
Singular field theories are, in many cases, amongst the most interesting class of models encountered in the description of natural phenomena [1, 2, 3]. Indeed, the most precisely tested mathematical theory to date, the Standard Model of Particle Physics, is a chiral quantum field theory, gauged under local transformations of the group $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ [4, 5, 6]. Further, it is postulated that any UV-complete theory of quantum gravity may possess no global symmetries, and that all symmetries must therefore be either spontaneously broken, or gauged [7, 8, 9]. It is thus clear that singular theories are of central importance to the advancement of modern theoretical physics.
Any attempt to rigorously develop a quantum theory requires a firm understanding of the classical counterpart upon which it is based. As such, in this article, we shall present a framework in which singular theories that possess scaling symmetries may be reduced to a lower-dimensional, dynamically-equivalent description, that is frictional in nature [10]. For theories of particles, we take, as our starting point, the geometrical constraint algorithm for pre-symplectic systems, developed by Nester et. al. in [11, 12, 13]. This is presented in section (III.1), following a brief review of elementary pre-symplectic geometry in section (II).
Having introduced this preliminary contextual framework, we provide an overview of contact and pre-contact geometry, and how this is used to describe the dynamical evolution of non-conservative systems. Section (VI) is then dedicated to the constraint algorithm for pre-contact Hamiltonian systems, which closely mirrors our presentation of the pre-symplectic constraint procedure.
Following this, we illustrate how the formalism of contact reduction, described in [14] and [15], may be extended to cases in which degeneracies require constraints to be placed upon the dynamical system of interest. We also consider the commutative relationship between excising redundant scaling degrees of freedom, and the restriction of a system’s phase space, finding that the order of operations is inconsequential for the final dynamics. Our formalism is then applied to a complete example, in which the symmetry-reduced description is deduced in each of the two possible orders. We have structured things in such a way so as to allow the reader familiar with constrained geometrical mechanics to omit sections (II) - (VI).
A finite-dimensional, manifestly covariant description of classical field theory is best formulated on fibred manifolds, within the context of multisymplectic geometry [16, 17, 18, 19, 20]. A constraint algorithm for treating singular field theories has been developed precisely in [21] and [22], and implemented in a number of interesting cases in [23] and [24]. Following a succinct review of multisymplectic geometry, and its implementation in the Lagrangian description of classical field theory, we dedicate section (X) to a discussion of non-conservative systems; action-dependent field theories are significantly less well-understood than their multisymplectic counterparts [25, 26]; however, due to their capacity for describing dissipative phenomena, such models have recently been the focus of heightened research interest. The study of non-conservative field theories is grounded in a rigorous mathematical framework, known as multicontact geometry [27, 28]; in [29], it was demonstrated that contact reduction may be extended to a field-theoretic context, and that the resulting reduced space is in fact a multicontact manifold. Drawing upon the formalism developed in this work, the latter part of the present article provides a generalisation of these ideas, applicable to singular field theories.
Finally, in a similar spirit to our treatment of the particle case, we conclude with a second complete example, in which we analyse a string-inspired, low-energy effective non-Abelian gauge theory [30, 31]; such a model is of particular interest, as it is found that the redundant degree of freedom coincides precisely with the dilaton field, the excision of which is, a priori, alarming, since it is generally known that the expectation value of this field sets the strength of the string coupling [32, 33].
The ideas developed over the course of the article constitute a foundational framework, within which to analyse singular (field) theories from a systematic perspective; it has recently been shown that the classical Einstein-Hilbert action possesses a redundant scaling degree of freedom, which is made manifest upon decomposing the spacetime metric into the product of a conformal factor and a symmetric rank-two tensor of fixed determinant [34]. Further comments on the application of our work to General Relativity, together with a number of open questions, and lines of future investigation, are presented in section (XIII).
II Geometrical Preliminaries
Suppose that $Q$ is an $n$-dimensional smooth manifold, corresponding to the configuration space of some mechanical system. Local coordinates on $Q$ are taken to be $q^{i}$, and we introduce a Lagrangian function $L:TQ\rightarrow\mathbb{R}$ on the tangent bundle [35, 36, 37]. Fibre coordinates of $TQ$ are denoted $v^{i}$, and we have the bundle projection $\pi:TQ\rightarrow Q$. The tangent bundle is endowed with a number of canonical geometrical features, including the vertical endomorphism $J:T(TQ)\rightarrow T(TQ)$, satisfying $J^{2}=0$, and $\textrm{rank}\,J=n$, which constitutes an almost-tangent structure on $TQ$ [38, 39]. Additionally, we have a distinguished Liouville vector field $\Delta\in\mathfrak{X}^{\infty}(TQ)$, defined such that its one-parameter family of transformations generates dilatations of the fibres of $TQ$. In local bundle coordinates $(q^{i},v^{i})$, we have
$J=\frac{\partial}{\partial v^{i}}\otimes dq^{i}\quad\quad\quad\quad\Delta=v^{i}\,\frac{\partial}{\partial v^{i}}$ (1)
Given a Lagrangian function $L$, the canonical structures $J$ and $\Delta$ are used to define a (pre-)symplectic Lagrangian system. In particular, the function $E_{L}:=\Delta(L)-L$ corresponds to the scalar energy of the mechanical system described by $L$, whilst $\theta_{L}:=\,^{t}\hskip-1.70717ptJ\circ dL$ is used to construct, the Poincaré-Cartan form $\omega_{L}:=-\,d\theta_{L}$. Clearly, $\omega_{L}$ is closed; non-degeneracy, on the other hand, is not assured, and throughout subsequent sections, we shall be interested in those systems for which $\omega_{L}$ is degenerate. In such cases, we refer to the triple $(TQ,\omega_{L},E_{L})$ as a pre-symplectic Lagrangian system [40], and from (1) it follows that
$E_{L}=\frac{\partial L}{\partial v^{i}}v^{i}-L\quad\quad\quad\quad\omega_{L}=-\,d\left(\frac{\partial L}{\partial v^{i}}dq^{i}\right)$ (2)
The degeneracy of the 2-form $\omega_{L}$ may be expressed as the requirement that the matrix of second derivatives
$W_{ij}:=\left(\frac{\partial^{2}L}{\partial v^{i}\partial v^{j}}\right)$
be of non-maximal rank, and thus singular. The equations of motion for a Lagrangian system $(TQ,\omega_{L},E_{L})$ are formulated geometrically via the following vector bundle morphism
$\begin{split}\flat_{\scriptscriptstyle L}\;:\;T(TQ)\,&\longrightarrow\,T^{*}(TQ)\\ X\;&\longmapsto\;\iota_{X}\omega_{L}\end{split}$ (3)
It should be noted that when $\omega_{L}$ is degenerate, $\flat_{\scriptscriptstyle L}$ is not an isomorphism. The dynamics are deduced seeking a vector field $X_{L}\in\mathfrak{X}^{\infty}(TQ)$, such that
$\flat_{\scriptscriptstyle L}(X_{L})=\iota_{X_{L}}\omega_{L}=dE_{L}$ (4)
When $(TQ,\omega_{L},E_{L})$ is a symplectic system, such vector fields are guaranteed to exist, and are unique. In addition, integral curves of $X_{L}$ project down to curves on the configuration space $Q$, and these latter curves correspond to physical trajectories of the mechanical system. Such vector fields are interchangeably referred to as holonomic or second-order differential equations (SODE) [41].
When the system of interest possesses a singular Lagrangian, not only are vector field solutions of (4) not guaranteed to exist, but since $\textrm{ker}\,\omega_{L}$ is non-trivial, any solution that is found is no longer unique [40]. In the most favourable of cases, there will exist a submanifold $P_{f}\hookrightarrow TQ$, upon which (4) admits solutions, and this space is deduced through use of a geometrical constraint algorithm [12, 13].
For the purposes of studying the contact reduction of singular systems, we shall work almost exclusively within the Hamiltonian formalism. Our motivation for this is twofold: on the one hand, the Hamiltonian framework provides a well-defined bracket structure, which readily allows us to classify the constraint functions, and deduce the dynamical evolution of any phase space variable. Further, not only does the Lagrangian formalism lack such a bracket, but the geometrical constraint algorithm introduces a myriad of additional complications, which arise owing to a need to impose that, at each step, the vector field solution satisfy the second-order problem. Since the dynamical content of the two formalisms is identical, we shall evade the undesirable features of the Lagrangian description by using it only as a means to obtain the corresponding Hamiltonian function.
For pre-symplectic Lagrangian systems, the Legendre map $FL:TQ\rightarrow T^{*}Q$ is not a diffeomorphism [42]; in practice, this translates into the more intuitive statement that one cannot ‘invert’ the momenta to solve for the $v^{i}$, and write down a Hamiltonian in a straightforward fashion. If we wish to consider symmetry reductions of singular theories, it will be necessary to restrict our attention to those cases in which the Lagrangian is almost-regular; such systems are characterised as follows
  • $\star$
    $M_{0}:=FL(TQ)$ is a closed submanifold of $T^{*}Q$
  • $\star$
    FLFL is a submersion onto its image
  • $\star$
    For every $p\in M_{0}$, the fibres $FL^{-1}(p)$ are connected submanifolds of TQTQ
Given an almost-regular Lagrangian system $(TQ,\omega_{L},E_{L})$, we denote the restriction of the Legendre map to its image via $FL_{0}$; since this is a submersion, it follows that there exists a unique function $H_{0}:M_{0}\rightarrow\mathbb{R}$, such that $FL_{0}^{*}\hskip 0.85358ptH_{0}=E_{L}$. On the cotangent bundle, there exists a canonical symplectic form $\omega$, expressed in local Darboux coordinates as $\omega=dq^{i}\wedge dp_{i}$; if $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$ denotes the inclusion map, then $M_{0}$ inherits a pre-symplectic form $\omega_{0}:=\jmath_{0}^{*}\,\omega$. We shall refer to the triple $(M_{0},\omega_{0},H_{0})$ as a pre-symplectic Hamiltonian system, defined on the primary constraint manifold $M_{0}$.
In order to formulate the dynamical problem, the vector bundle morphism $\flat:T(T^{*}Q)\rightarrow T^{*}(T^{*}Q)$ is restricted to $M_{0}$, and the resulting map denoted $\flat_{0}$. We then seek a vector field $X_{H}$, such that
$\flat_{0}(X_{H})=\iota_{X_{H}}\omega_{0}=dH_{0}$ (5)
As in the Lagrangian case, solutions generally do not exist on the entirety of $M_{0}$, thereby necessitating the implementation of a constraint algorithm, which is described in detail in section (III.1). The final constraint manifold $M_{f}$ is then the maximal subspace of the cotangent bundle upon which the dynamical problem (5) possesses well-defined (albeit non-unique) solutions.
III The Pre-symplectic Constraint Algorithm
Before reviewing how one uses geometrical principles to systematically restrict the phase space of a singular system, we begin by establishing a number of notational conventions used throughout. Firstly, we denote by $\langle\,\cdot\,$$\cdot\,\rangle$ the natural pairing between a vector space and its dual, writing $\langle dH_{0},TM_{0}\rangle$ to refer to the contraction of a particular object, such as $dH_{0}\in T^{*}M_{0}$, with all elements of the space $TM_{0}$.
In general, a (pre-)symplectic form $\omega$ on a manifold $M$ allows us to introduce a notion of orthogonality; in particular, for any subspace $N\subset M$, we define the symplectic orthogonal (or symplectic complement) of $TN$, denoted $TN^{\perp}$, as follows
$TN^{\perp}:=\{X\in TM|_{\scriptscriptstyle N}\;|\;\omega|_{\scriptscriptstyle N}(X,Y)=0\;\,\textrm{for all}\;Y\in TN\,\}$ (6)
In addition to the symplectic orthogonal, we also introduce the annihilator of a subspace $S\subset TM$ as
$S^{o}:=\{\alpha\in T^{*}M\;|\;\langle\alpha,v\rangle=0\;\;\textrm{for all}\;v\in S\,\}$ (7)
The manifold $M$ is said to be reflexive if $(M^{*})^{*}=M$, and topologically closed if $\flat:TM\rightarrow T^{*}M$ maps closed sets of $TM$ onto closed sets of $T^{*}M$ [11]; for this class of space, we find that, for a submanifold $N$ of $M$, the annihilator of the symplectic orthogonal of $TN$ is precisely the image of $TN$ under $\flat$: that is to say $(TN^{\perp})^{o}=\flat(TN)$.
III.1 The Hamiltonian Algorithm
When analysing the dynamics of singular systems, we shall use the almost-regular Lagrangian $L:TQ\rightarrow\mathbb{R}$ to obtain the pre-symplectic Hamiltonian system $(M_{0},\omega_{0},H_{0})$, introducing the embedding $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$. It will often be convenient to describe $M_{0}$ as the zero-set of a collection of primary constraint functions $\{\phi^{a}\}$, with $a=1,\,\cdots,n-k$, in which $n=\textrm{dim}\,Q$, and $k=\textrm{rank}\,W_{ij}$ [43]. We shall assume throughout that all subspaces generated are regular, closed submanifolds of $T^{*}Q$, whose natural injections are embeddings. We therefore begin by seeking a vector field $X_{H}$ which satisfies
$\flat_{0}(X_{H})=\iota_{X_{H}}\omega_{0}=dH_{0}$ (8)
Since $\omega_{0}$ is pre-symplectic, $\flat_{0}$ is not an isomorphism; consequently, an initial prerequisite for our geometrical equation to possess solutions is that, at every point, $dH_{0}$ must lie in the range of the map $\flat_{0}$. Given that such a condition is not guaranteed to be satisfied a priori, the first step of the constraint algorithm requires us to restrict $M_{0}$ to the following subset
$M_{1}:=\{x\in M_{0}\;|\;(dH_{0})_{x}\in\,\flat_{0}(T_{x}M_{0})\,\}$ (9)
From the discussion above, under the assumptions of reflexivity and topological closedness, we recall that the image $\flat_{0}(T_{x}M_{0})$ coincides with the annihilator of $T_{x}M_{0}^{\perp}$; this is particularly noteworthy, as $T_{x}M_{0}^{\perp}$ is precisely $\textrm{ker}\,(\omega_{0})_{x}$, which is readily deduced in local coordinates. Thus, from a pragmatic standpoint, it is preferable to characterise $M_{1}$ as follows
$M_{1}=\{x\in M_{0}\;|\;\langle(dH_{0})_{x},\textrm{ker}\,(\omega_{0})_{x}\rangle=0\,\}$ (10)
If $\{X^{A}\}$ constitutes a basis of $\textrm{ker}\,\omega_{0}$, then $M_{1}$ may be described locally as the zero-set of the functions
$\psi^{A}:=\langle dH_{0},X^{A}\rangle$ (11)
We refer to these objects as the secondary constraint functions. In order for any vector field solution $X_{H}$, defined over the points of $M_{1}$, to be physically meaningful, it must remain tangent to $M_{1}$. If this were not the case, the solution would tend to evolve off the constraint surface, and thus cease to obey the restrictions placed upon the system. Tangency of the solution is not guaranteed, and so must be imposed as an additional constraint, leading us to consider only those points of $M_{1}$ at which the solution may be written as $X_{H}=\jmath_{1*}Z$ for some $Z\in TM_{1}$; generally, $\jmath_{k}:M_{k}\hookrightarrow M_{k-1}$ is used to refer to the inclusion of $M_{k}$. The condition $X_{H}=\jmath_{1*}Z$ is not readily imposed locally, and so we shall use the following alternative criterion to characterise $M_{2}$
$M_{2}:=\{x\in M_{1}\;|\;\langle(dH_{0})_{x},T_{x}M_{1}^{\perp}\rangle=0\,\}$ (12)
That this provides an equivalent definition to that given by imposing $X_{H}=\jmath_{1*}Z$ follows from a series of simple calculations, outlined in [11]. Vector field solutions on $M_{2}$ are now tangent to $M_{1}$, but, in general, not to $M_{2}$, requiring this condition to be imposed as a further restriction. It is then clear from (9) and (12) how the algorithm must proceed: at the $k^{\textrm{th}}$ step, the submanifold $M_{k}$ is composed of all points of $M_{k-1}$ at which the solution is tangent to $M_{k-1}$, but not necessarily to $M_{k}$, leading us to restrict $M_{k}$ to the smaller set $M_{k+1}$. This condition may be imposed requiring that $X_{H}=\jmath_{k*}Z$ for some $Z\in TM_{k}$, or equivalently, by demanding that $dH_{0}$ belong to the annihilator $(TM_{k}^{\perp})^{o}$, so that
$M_{k+1}=\{x\in M_{k}\;|\;\langle(dH_{0})_{x},T_{x}M_{k}^{\perp}\rangle=0\,\}$ (13)
The constraint procedure may terminate in one of three ways. Firstly, there may be some $N\in\mathbb{N}$, for which $M_{N}=\emptyset$; in this case, the equation of motion (8) possesses no solution and the system $(M_{0},\omega_{0},H_{0})$ does not describe any physical configuration. Alternatively, the algorithm may yield a submanifold $M_{N}$, which is non-empty, but of dimension zero; such a space may be identified as a collection of isolated points, which forces the vector field solution to be identically zero. Evidently, neither of these eventualities is particularly interesting from a physical perspective. Consequently, we shall only consider those systems for which there exists some $N\in\mathbb{N}$, such that $M_{N+1}=M_{N}$, with $\textrm{dim}\,M_{N}\neq 0$. Upon deducing the value of $N$ for which the algorithm terminates, we shall declare $M_{f}:=M_{N}$ to be the final constraint manifold.
III.2 Geometry of the Final Constraint Submanifold
The algorithm of the previous section provides a systematic way to restrict the set of points of the cotangent bundle at which the equations of motion of a singular system admit everywhere-tangent solutions. At no point was it necessary to specify the nature of the constraints defining each subspace; once the final submanifold has been found, however, the dynamical evolution of observable quantities is deduced locally, which does require the distinction between first and second class constraints to be made.
Suppose that the constraint surface $M_{f}$ is defined as the zero-set of the functions $\phi^{\alpha}$, with $\alpha=1,\,\cdots,D$. We introduce the matrix $J$, whose entries are the Poisson brackets of the constraints i.e $J^{\alpha\beta}:=\{\phi^{\alpha},\phi^{\beta}\}$. If $\textrm{rank}\,J=D-F$, it follows that there exist $F$ $C^{\infty}$-linear combinations of the $\phi^{\alpha}$ which generate unobservable (gauge) transformations. [^1] Here, and throughout all subsequent analyses, we take the Dirac conjecture - that all first class constraints generate gauge transformations - to be true [44, 45, 46]. The goal is then to identify such combinations, corresponding to first class functions, and to modify the original set of constraints, so as to extract a maximal subfamily of second class functions. Note that, since $J$ is calculated from the Poisson bracket, it follows that $\textrm{rank}\,\omega_{f}=D-F$, in which $\omega_{f}:=\iota^{*}\omega$ is the closed 2-form naturally inherited by $M_{f}$, and $\imath:=\jmath_{f}\circ\cdots\circ\jmath_{0}$ defines the embedding of $M_{f}$ into $T^{*}Q$.
In order to extract the first class functions, we introduce a ‘vector’ $v_{\alpha}=(v_{1},\,\cdots,v_{D})$, and demand that
$\left(v_{\alpha}J^{\alpha\beta}\right)|_{M_{f}}=0\quad\textrm{for all}\;\beta=1,\,\cdots,D$ (14)
It is sufficient to seek these null vectors on the constraint manifold, since a function $f$ is only required to have weakly-vanishing Poisson bracket with the $\phi^{\alpha}$, to be characterised as first class. Having identified all linearly independent combinations satisfying (14), indexing the resulting vectors via $v^{(a)}$, the $a^{\textrm{th}}$ first class constraint is simply $\Omega^{a}:=v^{(a)}_{\alpha}\phi^{\alpha}$. In order to extract the maximal subfamily of second class constraints, we seek DFD-F linearly independent vectors $w^{(j)}$, which are not in the kernel of $J$, so that
$\underbrace{\Omega^{a}=v^{(a)}_{\alpha}\phi^{\alpha}}_{1^{\textrm{st}}\;\textrm{class}}\quad\quad\quad\quad\underbrace{\chi^{j}=w^{(j)}_{\alpha}\phi^{\alpha}}_{2^{\textrm{nd}}\;\textrm{class}}$
The constraint surface $M_{f}$ may now be characterised as
$M_{f}=\{x\in T^{*}Q\;|\;\Omega^{a}(x)=0\quad\textrm{for}\;a=1\,,\cdots,F\quad\textrm{and}\quad\chi^{j}(x)=0\quad\textrm{for}\;j=1,\,\cdots,D-F\;\}$ (15)
In the presence of second class functions, it is necessary to replace the usual Poisson bracket with the Dirac structure $\{\,\cdot\,,\cdot\,\}_{D}$ [44, 47]; we introduce the matrix $C^{ij}:=\{\chi^{i},\chi^{j}\}$, which, by construction, is of maximal rank DFD-F, and thus non-singular. If $C_{ij}$ are the elements of the inverse $C^{-1}$, the Dirac bracket of two phase space functions $f,g\in C^{\infty}(T^{*}Q)$ is given by
$\{f,g\}_{D}:=\{f,g\}-\{f,\chi^{i}\}\,C_{ij}\{\chi^{j},g\}$ (16)
The Dirac bracket has a number of noteworthy properties:
  • $\{f,gh\}_{D}=\{f,g\}_{D}\,h+g\{f,h\}_{D}$ for all $f,g,h\in C^{\infty}(T^{*}Q)$
  • $\{f,\chi^{j}\}_{D}=0$ for any $f\in C^{\infty}(T^{*}Q)$
  • $\{f,g\}_{D}\approx\{f,g\}$ for $f$ first class and $g$ arbitrary
We now impose the second class constraints $\chi^{j}=0$ as strong equalities, which defines a hypersurface
$M_{\chi}:=\{x\in T^{*}Q\;|\;\chi^{j}(x)=0\quad\textrm{for}\;j=1,\,\cdots,D-F\,\}\;\subset\;T^{*}Q$ (17)
with inclusion $\jmath_{\chi}:M_{\chi}\hookrightarrow T^{*}Q$, which inherits a closed 2-form $\omega_{\chi}:=\jmath^{*}_{\chi}\omega$ which is symplectic. The Dirac bracket is precisely the algebraic structure required to ‘transfer’ this symplectic structure to functions defined on the whole of $T^{*}Q$. In particular, the dynamical evolution of any phase space function $f\in C^{\infty}(T^{*}Q)$ may be deduced from $\dot{f}=\{f,H_{T}\}_{D}$, in which $H_{T}:=H+\lambda_{a}\Omega^{a}$ is the total Hamiltonian; $\lambda^{a}$ are arbitrary Lagrange multipliers, and $H$ is any extension of $H_{0}$ to the full phase space.
Within $M_{\chi}$, we consider the first class surface
$M_{\scriptscriptstyle\Omega}:=\{x\in M_{\chi}\;|\;\Omega^{a}(x)=0\quad\textrm{for}\;a=1,\,\cdots,F\,\}$ (18)
$M_{\scriptscriptstyle\Omega}$ is an embedded submanifold of $M_{\chi}$, with inclusion $\jmath_{\scriptscriptstyle\Omega}:M_{\scriptscriptstyle\Omega}\hookrightarrow M_{\chi}$; the 2-form $\omega_{\scriptscriptstyle\Omega}:=\jmath^{*}_{\scriptscriptstyle\Omega}\,\omega_{\chi}$ is maximally degenerate, since $M_{\scriptscriptstyle\Omega}$ is first class. It is straightforward to show that the Hamiltonian vector fields $X_{a}$ corresponding to $\Omega^{a}$ via $\iota_{X_{a}}\omega_{\chi}=d\Omega^{a}$ are tangent to $M_{\scriptscriptstyle\Omega}$, since
$\mathcal{L}_{X_{a}}\Omega^{b}=d\Omega^{b}[X_{a}]=\omega_{\chi}(X_{b},X_{a})=\{\Omega^{b},\Omega^{a}\}\approx 0$
Here, the weak equality is taken to signify ‘when restricted to $M_{\scriptscriptstyle\Omega}$.’ In general, a vector field $\xi\in\mathfrak{X}^{\infty}(T^{*}Q)$ may only be tangent to $M_{\scriptscriptstyle\Omega}$ if $\mathcal{L}_{\xi}\Omega^{a}\approx 0$ for all $\Omega^{a}$: that is if $\xi\in\textrm{ker}\,\omega_{\scriptscriptstyle\Omega}$; since the first class constraints are assumed to be independent (by which we mean that their differentials are linearly independent) it follows that $\textrm{ker}\,\omega_{\scriptscriptstyle\Omega}$ is spanned by the Hamiltonian vector fields $X_{a}$.
More formally, at each point $x\in M_{\scriptscriptstyle\Omega}$, the symplectic orthogonal
$T_{x}M_{\scriptscriptstyle\Omega}^{\perp}:=\{v\in T_{x}M_{\chi}\;|\;\omega_{\chi}(v,w)=0\quad\textrm{for all}\;w\in T_{x}M_{\scriptscriptstyle\Omega}\,\}$
satisfies $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}\subseteq T_{x}M_{\scriptscriptstyle\Omega}$, and so $M_{\scriptscriptstyle\Omega}$ constitutes a coisotropic submanifold of $M_{\chi}$ [48]; for such a class of space, the distribution $x\mapsto T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ is involutive, since, for any two $X,Y\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, we take some $Z\in T_{x}M_{\scriptscriptstyle\Omega}$, and compute
$d\omega_{\chi}(X,Y,Z)=X\omega_{\chi}(Y,Z)-Y\omega_{\chi}(Z,X)+Z\omega_{\chi}(X,Y)-\omega_{\chi}([X,Y],Z)+\omega_{\chi}([X,Z],Y)-\omega_{\chi}([Y,Z],X)$
Since $\omega_{\chi}$ is closed, $d\omega_{\chi}(X,Y,Z)=0$; moreover, $X$ and $Y$ were taken to be in the symplectic orthogonal $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, so that the first three terms on the right-hand side vanish. Similar considerations lead us to conclude that
$d\omega_{\chi}(X,Y,Z)=-\,\omega_{\chi}([X,Y],Z)=0$
so that if $X,Y\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, then $X,Y]\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$. Hence, $x\mapsto T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ is an involutive distribution, which, locally, we may write as the span of the $X_{a}$ [49]: that is to say
$T_{x}M_{\scriptscriptstyle\Omega}^{\perp}=\textrm{span}(X_{1}|_{x}\,,\cdots,X_{F}|_{x})\,\subseteq\,T_{x}M_{\scriptscriptstyle\Omega}$
Since the Hamiltonian vector fields of the first class constraints define an involutive distribution, we know, from the Frobenius theorem, that there exists a foliation $\mathcal{F}$ of $M_{\scriptscriptstyle\Omega}$ into $F$-dimensional submanifolds, which, in this context, we refer to as gauge leaves [50, 51]. The gauge orbits correspond to integral curves of the $X_{a}$, and they map points of $M_{\scriptscriptstyle\Omega}$ into other points, which produce dynamically indistinguishable configurations. Note that these results are valid only on $M_{\scriptscriptstyle\Omega}$: elsewhere, the vector fields $X_{a}$ do not, in general, close under the Lie bracket.
As a result of the degeneracy of $\omega_{\scriptscriptstyle\Omega}$, the constraint surface does not inherit a well-defined bracket structure. It is therefore necessary to construct a reduced space, quotienting out by the gauge orbits; this reduced phase space identifies all points related by gauge transformations, and thus contains only configurations which are physically inequivalent. Upon constructing this space, there arise two distinct possibilities: in the most favourable of cases, the algebra formed by the first class constraints is a genuine Lie algebra; this occurs precisely when the structure functions $f^{ab}_{c}$, appearing in $\{\Omega^{a},\Omega^{b}\}=f^{ab}_{c}\Omega^{c}$, are constant over the whole of $M_{\scriptscriptstyle\Omega}$. In these conditions, the finite gauge transformations are generated by the elements of the corresponding Lie group $G$ that are path-connected to the identity, [^2] This lifting of the elements of the Lie algebra to the group is, of course, provided by the exponential map. and they give rise to a smooth action of $G$ on $M_{\scriptscriptstyle\Omega}$. It is then possible to perform a reduction procedure, using results similar to those of the Marsden-Weinstein theorem [52, 53], to obtain a phase space $M_{\scriptscriptstyle\Omega}/G$, which is symplectic [54, 55].
In most cases, such conditions will not be realised; it is often found that the Dirac bracket between two first class constraints is strongly equal to some combination of the phase space variables which vanishes on $M_{\scriptscriptstyle\Omega}$. The structure functions are then of the form $f^{ab}_{c}(q^{i},p_{i})$, and they define a Lie algebroid on $M_{\scriptscriptstyle\Omega}$ [56]. Formally, it is again possible to quotient out by the gauge orbits, obtaining a physical phase space that is symplectic; however it is almost invariably more straightforward to introduce a choice of gauge-fixing. Since we are only concerned with the dynamical evolution of gauge-invariant observable quantities, the precise choice of gauge-fixing is somewhat arbitrary, provided that the gauge is obtainable via a finite number of infinitesimal transformations, and that it fixes all degrees of freedom, leaving no remnant gauge symmetry [57, 58].
Upon carrying out the pre-symplectic reduction process, we obtain a physical state space $\mathcal{P}$, together with a projection map $\pi:M_{\scriptscriptstyle\Omega}\rightarrow\mathcal{P}$, sending each point on the constraint surface to its equivalence class under gauge orbits. The reduced phase space $\mathcal{P}$ is symplectic, with associated 2-form $\pi^{*}\omega_{\scriptscriptstyle\mathcal{P}}=\jmath^{*}\omega$, where $\jmath:=\jmath_{\chi}\circ\jmath_{\scriptscriptstyle\Omega}:M_{\Omega}\rightarrow T^{*}Q$ defines the embedding of $M_{\scriptscriptstyle\Omega}$ into $T^{*}Q$.
IV Contact Geometry
In order to examine the commutative relationship between contact reduction by scaling symmetries, and the restriction of the phase space of mechanical systems described by singular Lagrangians, it will be necessary to understand how the geometric algorithm we have presented for pre-symplectic manifolds is to be adapted to the pre-contact spaces, which naturally arise upon eliminating a scaling degree of freedom [14, 15, 10]. In anticipation of this, we shall briefly introduce the fundamental ideas of contact and pre-contact geometry [59, 60], together with their implementation in the description of Lagrangian and Hamiltonian systems [61, 62, 63].
IV.1 Contact Manifolds
A contact structure on a smooth manifold $C$, of dimension $2n+1$, is a maximally-non-integrable distribution $\xi\subset TC$; locally, this distribution may be described as the kernel of some $\eta\in\Omega^{1}(U\subset C)$, referred to as a contact form. If the quotient line bundle $TC/\xi\rightarrow C$ is trivial, then $\xi$ is coorientable, and $\eta$ may be extended to the whole of $C$ [64]. The non-integrability condition of $\xi$ may be expressed as the requirement that
$\eta\wedge d\eta^{n}\neq 0$ (19)
over the whole of $C$. The coorientability of $\xi$ implies that $C$ itself is orientable; in this case, $\eta\wedge d\eta^{n}$ provides a volume form on $C$, and since, in general, $\xi$ is of rank $2n$, at each point $p\in C$, we have the decomposition
$T_{p}C=\textrm{ker}\,\eta_{p}\oplus\textrm{ker}\,d\eta_{p}$
and $d\eta_{p}$ defines a skew-symmetric, non-degenerate, bilinear form of maximal rank on the $2n$-dimensional subspace $\xi_{p}$. We henceforth assume that $\xi$ is coorientable, and work with the globally-defined form $\eta$; every contact manifold $(C,\eta)$ admits a distinguished Reeb vector field $\mathcal{R}\in\mathfrak{X}^{\infty}(C)$, defined via
$\iota_{\mathcal{R}}d\eta=0\quad\quad\quad\quad\iota_{\mathcal{R}}\eta=1$ (20)
As for symplectic manifolds, around each point $p\in C$, we may always find a local chart of Darboux coordinates $(x^{1},\cdots,x^{n},y_{1},\cdots,y_{n},z)$, in which $\eta$ takes the form $\eta=dz-y_{i}\,dx^{i}$.
IV.2 Lagrangian Mechanics
Consider the $n$-dimensional configuration space $Q$, with local coordinates $(q^{i})$; the extended tangent bundle $TQ\times\mathbb{R}$ is a manifold of dimension $2n+1$, with induced bundle coordinates $(q^{i},v^{i},z)$. The canonical structures of $TQ$ are easily extended to $TQ\times\mathbb{R}$ [62], and have local coordinate expression (1). A contact Lagrangian is a function $L:TQ\times\mathbb{R}\rightarrow\mathbb{R}$, which allows us to define the Cartan forms $\theta_{L}$ and $\omega_{L}$, together with the energy function $E_{L}$. The space $TQ\times\mathbb{R}$ is made into a contact manifold upon introducing the 1-form $\eta_{L}:=dz-\theta_{L}$, and we refer to the triple $(TQ\times\mathbb{R},\eta_{L},E_{L})$ as a contact Lagrangian system.
IV.3 The Legendre Map
Whilst contact Lagrangian systems are introduced on the space $TQ\times\mathbb{R}$, their Hamiltonian counterparts are studied on the extended cotangent bundle $T^{*}Q\times\mathbb{R}$, which admits the following projections
$\begin{split}\pi_{0}:T^{*}Q\times\mathbb{R}\rightarrow Q\times\mathbb{R}&\quad\quad\quad\quad\pi_{1}:T^{*}Q\times\mathbb{R}\rightarrow T^{*}Q\\ \rho:\,&T^{*}Q\times\mathbb{R}\rightarrow\mathbb{R}\end{split}$ (21)
The Legendre map $FL:TQ\times\mathbb{R}\rightarrow T^{*}Q\times\mathbb{R}$ acts on the local bundle coordinates $(q^{i},p_{i},z)$ of $T^{*}Q\times\mathbb{R}$ according to
$FL^{*}q^{i}=q^{i}\quad\quad FL^{*}p_{i}=\frac{\partial L}{\partial v^{i}}\quad\quad FL^{*}z=z$ (22)
The extended cotangent bundle inherits a contact structure, defined in terms of the canonical structures of $T^{*}Q$; in particular, if $\vartheta\in\Omega^{1}(T^{*}Q)$ and $\Omega=-\,d\vartheta\in\Omega^{2}(T^{*}Q)$ denote the Liouville 1- and 2-forms on the cotangent bundle, then we have
$\theta=\pi_{1}^{*}\vartheta\quad\quad\textrm{and}\quad\quad\omega=-\,d\theta=\pi_{1}^{*}\Omega$ (23)
The Legendre map relates these forms with $\theta_{L}$ and $\omega_{L}$ in the obvious manner
$FL^{*}\theta=\theta_{L}\quad\quad\textrm{and}\quad\quad FL^{*}\omega=\omega_{L}$ (24)
The contact 1-form $\eta$ is then $\eta=dz-\theta$, and the Reeb vector field $\mathcal{R}$ follows from (20); in local coordinates, we have
$\theta=p_{i}\,dq^{i}\quad\quad\omega=dq^{i}\wedge dp_{i}\quad\quad\eta=dz-p_{i}\,dq^{i}\quad\quad\mathcal{R}=\frac{\partial}{\partial z}$ (25)
In complete analogy to the symplectic case, the Legendre map provides a set of criteria, according to which we may classify a given contact Lagrangian system; in particular, we shall restrict ourselves to those singular systems which are categorised as ‘almost regular.’
IV.4 Hamiltonian Formalism
If $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a hyperregular contact Lagrangian system, we introduce the Hamiltonian as the unique function $H:T^{*}Q\times\mathbb{R}\rightarrow\mathbb{R}$, such that $FL^{*}H=E_{L}$. The triple $(T^{*}Q\times\mathbb{R},\eta,H)$ is then said to constitute a contact Hamiltonian system, and there exists a bundle morphism
$\begin{split}\bar{\flat}\;:\;T(T^{*}Q\times\mathbb{R})&\longrightarrow T^{*}(T^{*}Q\times\mathbb{R})\\ v\;&\longmapsto\iota_{v}\,d\eta+(\iota_{v}\eta)\eta\end{split}$ (26)
The equations of motion are deduced seeking a vector field $X_{H}\in\mathfrak{X}^{\infty}(T^{*}Q\times\mathbb{R})$, which satisfies
$\bar{\flat}(X_{H})=dH-\left(\mathcal{R}(H)+H\right)\eta$ (27)
Suppose that we decompose the vector field $X_{H}$ as
$X_{H}=A^{i}\frac{\partial}{\partial q^{i}}+B_{i}\frac{\partial}{\partial p_{i}}+C\frac{\partial}{\partial z}$ (28)
then the coefficient functions $A^{i}$, $B_{i}$, and $C$ satisfy
$A^{i}=\frac{\partial H}{\partial p_{i}}\quad\quad B_{i}=-\,\biggr(\frac{\partial H}{\partial q^{i}}+p_{i}\,\frac{\partial H}{\partial z}\biggr)\quad\quad C=p_{i}\,\frac{\partial H}{\partial p_{i}}-H$ (29)
V Pre-contact Systems
All results presented thus far are applicable to regular and hyperregular systems; however, the main objective of this work is to present a mathematical framework within which to analyse scaling symmetries in singular theories. Consequently, we shall dedicate the following sections to developing a number of tools required to treat such systems.
V.1 Jacobi Structures
Both symplectic and contact manifolds belong to a broader class of object known as Jacobi manifolds [65, 66, 67]; in general, a Jacobi manifold is a triple $(M,\Lambda,E)$, in which $\Lambda\in\mathfrak{X}^{2}(M)$ is a multivector field of degree two, and $E\in\mathfrak{X}^{\infty}(M)$ is a vector field; together, these objects satisfy [60]
$\Lambda,\Lambda]=2E\wedge\Lambda\quad\quad\quad\mathcal{L}_{E}\Lambda=[E,\Lambda]=0$ (30)
in which $\cdot\,$$\cdot\,$] denotes the Schouten-Nijenhuis bracket [68, 69, 70, 71, 72]. Given a Jacobi manifold $(M,\Lambda,E)$, the bivector field $\Lambda$ allows us to introduce a bundle morphism [^3] It should be noted that our conventions differ to those of [60] and [73]; this is primarily owing to our desire to recover a Poisson bracket that is in-line with conventions used throughout the physics literature.
$\begin{split}\sharp_{\Lambda}\;:\;T^{*}M&\longrightarrow TM\\ \xi&\longmapsto\Lambda(\;\cdot\;,\,\xi)\end{split}$ (31)
To every function $f\in C^{\infty}(M)$, there corresponds a Hamiltonian vector field $X_{f}\in\mathfrak{X}^{\infty}(M)$, given by
$X_{f}:=\sharp_{\Lambda}(df)-fE$ (32)
Finally, we define the Jacobi bracket between two functions $f,g\in C^{\infty}(M)$ as
$\begin{split}\{\,\cdot\;,\,\cdot\,\}_{J}\;:\;C^{\infty}(M)\times C^{\infty}(M)&\longrightarrow C^{\infty}(M)\\ (f,g)&\longmapsto\{f,g\}_{J}:=\Lambda(df,dg)+fE(g)-gE(f)\end{split}$ (33)
The pair $\left(C^{\infty}(M),\{\,\cdot\;,\,\cdot\,\}_{J}\right)$ then constitutes a Lie algebra, satisfying a weaker form of the Leibnitz rule
$\{f,gh\}_{J}=h\{f,g\}_{J}+g\{f,h\}_{J}+ghE(f)$ (34)
with $f,g,h\in C^{\infty}(M)$. The evolution of an observable $f\in C^{\infty}(M)$ along the Hamiltonian flow is given by
$\dot{f}=X_{H}(f)=df[X_{H}]=\{f,H\}_{J}-fE(H)$ (35)
For a contact manifold $(M,\eta)$, the vector field $E$ is simply $\mathcal{R}$, and $\Lambda$ is given locally by
$\Lambda=\frac{\partial}{\partial q^{i}}\wedge\frac{\partial}{\partial p_{i}}+p_{i}\,\frac{\partial}{\partial z}\wedge\frac{\partial}{\partial p_{i}}$ (36)
We also have $\dot{f}=X_{H}(f)=\{f,H\}_{J}-f\mathcal{R}(H)$, with the following expression for the Jacobi bracket
$\{f,g\}_{J}=\frac{\partial f}{\partial q^{i}}\frac{\partial g}{\partial p_{i}}-\frac{\partial f}{\partial p_{i}}\frac{\partial g}{\partial q^{i}}+\left(p_{i}\frac{\partial g}{\partial p_{i}}-g\right)\frac{\partial f}{\partial z}-\left(p_{i}\frac{\partial f}{\partial p_{i}}-f\right)\frac{\partial g}{\partial z}$ (37)
V.2 Pre-contact Manifolds
The final piece of mathematical background required to introduce the constraint algorithm is that of pre-contact manifolds [62, 74]; suppose that $(M,\eta)$ is a $(2n+1)$-dimensional manifold, equipped with a 1-form $\eta$, such that the distribution
$\mathcal{C}:=\textrm{ker}\,\eta\,\cap\,\textrm{ker}\,d\eta\;\subseteq\;TM$ (38)
is of rank 2n+1p2n+1-p; in this case, we refer to $\mathcal{C}$ as the characteristic distribution, and say that $\eta$ is of class $p$. In general, if $\eta$ is of class $2k+1$, then we have
$\eta\wedge d\eta^{k}\neq 0\quad\quad\textrm{but}\quad\quad\eta\wedge d\eta^{k+1}=0$ (39)
In addition, around each point $p\in M$, there exist local coordinates $(x_{1},\cdots,x_{k},y^{1},\cdots,y^{k},z,w_{1},\cdots,w_{s})$, with s=2(nk)s=2(n-k), such that $\eta$ is given by $\eta=dz-y_{i}\,dx^{i}$, and the characteristic distribution is spanned by
$\mathcal{C}\,=\,\,\biggr\langle\frac{\partial}{\partial w_{1}},\,\cdots,\frac{\partial}{\partial w_{s}}\biggr\rangle$
Upon introducing a function $H:M\rightarrow\mathbb{R}$, we refer to the triple $(M,\eta,H)$ as a pre-contact Hamiltonian system; as in the case of regular contact systems, we introduce the vector bundle morphism $\bar{\flat}:TM\rightarrow T^{*}M$, with precisely the same action as in (26).
Every pre-contact manifold admits a Reeb vector field $\mathcal{R}$, defined according to the conditions $\iota_{\mathcal{R}}\,d\eta=0$ and $\iota_{\mathcal{R}}\eta=1$; however, the novelty is that this field is no longer unique. In particular, any $X\in\mathfrak{X}^{\infty}(M)$ is a Reeb vector field if $\bar{\flat}(X)=\eta$ [60].
Given a distribution $\mathcal{D}\subset TM$, we define the following notion of orthogonality
$\mathcal{D}^{\perp}:=\{v\in TM\;|\;\bar{\flat}(w)(v)=0\quad\textrm{for all}\;w\in\mathcal{D}\,\}$ (40)
Note that since $\bar{\flat}(v)(w)$ is neither symmetric or antisymmetric under the exchange of $v$ and $w$, the above definition is strictly the right complement of $\mathcal{D}$; it is of course possible to define the left complement; however, $\mathcal{D}^{\perp}$ is the space we shall use within the constraint algorithm.
VI Pre-contact Constraint Algorithm
We begin by considering a pre-contact Lagrangian system $(TQ\times\mathbb{R},\eta_{L},E_{L})$, of class $2k+1$, which occurs precisely when the rank of the Hessian matrix is equal to $k$ over the whole of $TQ\times\mathbb{R}$. The extended cotangent bundle $T^{*}Q\times\mathbb{R}$ has a canonical contact form $\eta$, whose construction was given in (23); if we introduce the inclusion $\kappa_{0}:P_{0}:=FL(TQ\times\mathbb{R})\hookrightarrow T^{*}Q\times\mathbb{R}$, then $P_{0}$ inherits the structure of a pre-contact manifold, with corresponding 1-form $\eta_{0}=\kappa_{0}^{*}\eta=FL_{0}^{*}\,\eta_{L}$, where $FL_{0}$ refers to the restriction of the Legendre map to $P_{0}$. This pre-contact manifold is then made into a Hamiltonian system upon introducing the unique function $H_{0}:P_{0}\rightarrow\mathbb{R}$, which satisfies $FL_{0}^{*}\,H_{0}=E_{L}$. In the interest of reducing cumbersome notation in what follows, we introduce
$\alpha_{0}:=dH_{0}-\left(\mathcal{R}(H_{0})+H_{0}\right)\eta_{0}$
and write the restriction of the bundle morphism $\bar{\flat}$ to $P_{0}$ as $\bar{\flat}_{0}$; thus, the dynamical problem we seek to solve is that of finding a vector field $X_{H}\in\mathfrak{X}^{\infty}(P_{0})$, such that
$\bar{\flat}_{0}(X_{H})=\alpha_{0}$ (41)
We first restrict $P_{0}$ to the subset of points at which $\alpha_{0}$ is in the range of $\bar{\flat}_{0}$: that is to say
$P_{1}:=\{x\in P_{0}\;|\;(\alpha_{0})_{x}\in\bar{\flat}_{0}(T_{x}P_{0})\,\}$ (42)
As was the case for the pre-symplectic algorithm, it is convenient to make use of the relationship between the annihilator of the space $TP_{0}^{\perp}$ and the image of $TP_{0}$ under $\bar{\flat}_{0}$ to write $P_{1}$ as
$P_{1}=\{x\in P_{0}\;|\;\langle\,(\alpha_{0})_{x},T_{x}P_{0}^{\perp}\,\rangle=0\,\}$ (43)
Recall that for the pre-symplectic manifold $M_{0}$, we observed that $TM_{0}^{\perp}=\textrm{ker}\,\omega_{0}$, which led to the recharacterisation of $M_{1}$ given in (10); in the present context, we find that $TP_{0}^{\perp}=\mathcal{C}$, which provides the following, more practical condition for deducing $P_{1}$
$P_{1}=\{x\in P_{0}\;|\;\langle(\alpha_{0})_{x},\mathcal{C}_{x}\rangle=0\,\}$ (44)
Upon finding a basis for the characteristic distribution, $P_{1}$ may be described as the zero-set of a collection of secondary constraint functions. We must now impose that any vector field solution on $P_{1}$ be tangent to this space, which generally requires us to restrict to the smaller set
$P_{2}:=\{x\in P_{1}\;|\;\langle(\alpha_{0})_{x},T_{x}P_{1}^{\perp}\rangle=0\,\}$ (45)
The tangency condition is then imposed recursively, producing a series of embedded submanifolds $\kappa_{i+1}:P_{i+1}\hookrightarrow P_{i}$, until we reach the final constraint manifold $P_{f}$, where further iteration of the algorithm produces no new conditions. In this case, the geometric equation $\bar{\flat}_{0}(X_{H})=\alpha_{0}$ has a well-defined, albeit non-unique solution which is everywhere-tangent to $P_{f}$. The non-uniqueness follows from the observation that we may add to $X_{H}$ any element of $\mathcal{C}\cap TP_{f}$, without altering the physical dynamics or leaving the constraint surface.
In order to deduce local expressions for the dynamical evolution of extended phase space functions, we shall suppose that $P_{f}$ may be written as the zero-set of the $N$ functions $\{\Phi^{\alpha}\}$. Introducing the matrix of brackets $K^{\alpha\beta}:=\{\Phi^{\alpha},\Phi^{\beta}\}_{J}$, with $\textrm{rank}\,K=N-F$, we extract $F$ first class functions $\Omega^{a}$, and NFN-F second class functions $\chi^{j}$, following precisely the same procedure as in (14).
The Jacobi bracket of the second class constraints again forms an invertible matrix $C^{ij}:=\{\chi^{i},\chi^{j}\}_{J}$ of size $(N-F)\times(N-F)$, whose inverse we denote $C_{ij}$. The extended cotangent bundle admits a second Jacobi structure $(\Lambda_{DJ},\mathcal{R}_{DJ})$, whose corresponding bracket is the Dirac-Jacobi bracket $\{\,\cdot\;,\,\cdot\,\}_{DJ}$. For two functions $f,g\in C^{\infty}(T^{*}Q\times\mathbb{R})$, we have
$\{f,g\}_{DJ}:=\{f,g\}_{J}-\{f,\chi^{i}\}_{J}\,C_{ij}\{\chi^{j},g\}_{J}$ (46)
The Reeb field $\mathcal{R}_{DJ}$ is deduced from the observation that, for any function $f$, we have $\mathcal{R}_{DJ}(f)=\{1,f\}_{DJ}$; we find that
$\mathcal{R}_{DJ}:=\mathcal{R}+C_{ij}\mathcal{R}(\chi^{j})\biggr[\chi^{i}\mathcal{R}-\sharp_{\Lambda}(d\chi^{i})\biggr]$ (47)
in which $\sharp_{\Lambda}$ is defined in (31). In addition to satisfying the weak Leibnitz rule, the Dirac-Jacobi bracket vanishes when one or more of its arguments is second class. With this bracket at our disposal, it is now straightforward to deduce the dynamical evolution of any extended phase space function. The total Hamiltonian is $H_{T}:=H+u_{a}\Omega^{a}$, in which $H$ is an arbitrary extension of $H_{0}$ to $T^{*}Q\times\mathbb{R}$, and $u_{a}$ are undetermined Lagrange multipliers. The time evolution of a function $f\in C^{\infty}(T^{*}Q\times\mathbb{R})$ is then given by
$\dot{f}=X_{H_{T}}(f)=\{f,H_{T}\}_{DJ}-f\mathcal{R}_{DJ}(H_{T})$ (48)
Upon expanding this expression, making use of the weak Leibnitz rule, we find that
$\dot{f}=\left(X_{H}+u_{a}X_{\Omega^{a}}\right)(f)+\Omega^{a}\Lambda_{DJ}(df,du_{a})$ (49)
When evaluated on the constraint surface $P_{f}$, the term proportional to $\Omega^{a}$ vanishes identically, leaving
$\dot{f}\approx\left(X_{H}+u_{a}X_{\Omega^{a}}\right)(f)$ (50)
VII Contact Reduction of Singular Systems
Over the course of previous sections, we have constructed a framework in which the dynamics of constrained Hamiltonian systems may be analysed in a geometrical way. With these ideas in-mind, we shall now illustrate how contact reduction by scaling symmetries generalises to singular theories. In order to facilitate our analysis, we begin with an overview of the reduction process for regular systems; a more detailed discussion of these ideas, as well as explicit examples, may be found in [15, 59, 75, 76].
VII.1 Scaling Symmetries of Regular Hamiltonian Systems
Given a symplectic Hamiltonian system, $(M,\omega,H)$, we declare a vector field $D\in\mathfrak{X}^{\infty}(M)$ to constitute a scaling symmetry of degree $\Lambda$ if the following two conditions hold
$\mathfrak{L}_{D}\hskip 0.56905pt\omega=\omega\quad\quad\quad\quad\mathfrak{L}_{D}H=\Lambda H$ (51)
Here, and throughout what follows, we use the symbol $\mathfrak{L}$ to denote the Lie derivative, reserving the calligraphic $\mathcal{L}$ for the Lagrangian density, encountered in our treatment of multisymplectic field theory.
As a consequence of (51), we see that if $X_{H}\in\mathfrak{X}^{\infty}(M)$ is the unique Hamiltonian vector field corresponding to $H$ via the relationship $\iota_{X_{H}}\omega=dH$, then
$\iota_{\scriptscriptstyle[D,X_{H}]}\hskip 0.85358pt\omega=[\mathfrak{L}_{D},\iota_{X_{H}}]\hskip 1.13809pt\omega=(\Lambda-1)\,dH$ (52)
Since $\omega$ is, by assumption, non-degenerate, it follows that $D,X_{H}]=(\Lambda-1)X_{H}$, so that the effect of $D$ is to rescale all trajectories by the same non-zero factor, and may thus be considered a generator of changes in global scale. The deduction of physical law is, fundamentally, an empirical process; a measurement of length, for instance, is only meaningful as compared to some fiducial length, taken to represent a ‘metre stick’ [77]. Under the effect of a scaling symmetry, not only is the length to be measured scaled, but so too is the instrument used to perform the measurement. Consequently, a dynamical similarity maps between configurations whose constituents are rescaled in such a way so as to produce no observable effect. Such a transformation is the mathematical realisation of Poincaré’s thought experiment, where one imagines waking up to a world in which all distances have suffered a doubling in scale [78]: any attempt to discern the rescaled world from the original configuration will be futile, since observers have access only to ratios of dimensionful quantities. We shall develop these ideas somewhat further in section (XII), in the context of low-energy effective gauge theories, examining arguments presented against Poincaré’s stance that a rescaled world would be indistinguishable from our own.
From a minimalistic viewpoint, a mathematical description whose ontology contains redundant degrees of freedom, pertaining to unobservable global scales, is highly undesirable [79, 80]. We argue that, where possible, any theory of the natural world should be formulated exclusively in terms of empirically accessible quantities, together with those (possibly unobservable) parameters, whose presence contributes to the closure of the algebra of dynamical observables. The identification of a scaling symmetry indicates that our mathematical framework contains just such a redundancy, and the role of ‘contact reduction’ is to make precise the heuristic notion of eliminating this superfluous structure.
If the vector field $D$ is such that its flow acts freely and properly on $M$, then the quotient space $M/\hskip-2.84526pt\sim$ formed by identifying points connected by $D$-orbits has the structure of a smooth manifold of dimension $\textrm{dim}\,M-1$; moreover, the map $\pi:M\rightarrow M/\hskip-2.84526pt\sim$, sending each point of $M$ to its equivalence class, is a submersion [59, 81, 82]. In addition to being a smooth manifold, the quotient space also inherits a contact structure, defined according to $\xi:=\pi_{*}\,\textrm{ker}\left(\iota_{D}\hskip 0.56905pt\omega\right)$ [14]. We define a scaling function to be any $\rho:M\rightarrow\mathbb{R}_{+}$, such that $\mathfrak{L}_{D}\rho=\rho$; the existence of a global scaling function allows the contact distribution $\xi$ to be expressed as the kernel of a well-defined 1-form $\eta$ on $M/\hskip-2.84526pt\sim\,$
$\pi^{*}\eta:=\frac{\iota_{D}\hskip 0.56905pt\omega}{\rho}$ (53)
Assuming the existence of a global scaling function, we henceforth adopt the notation $(C,\eta)$ to refer to the contact manifold $M/\hskip-2.84526pt\sim$. Orbits of the symplectic Hamiltonian system project to well-defined curves on $C$; more precisely, the Hamiltonian vector field $X_{H}\in\mathfrak{X}^{\infty}(M)$ generates a line field $\textrm{span}\left(X_{H}\right)$ on $M$, which is projected to a line field $\pi_{*}\,\textrm{span}\left(X_{H}\right)$ on $C$. Additionally, there exists a contact Hamiltonian function $H^{c}:C\rightarrow\mathbb{R}$, which, on the level-set $\pi(H=0)$, is calculated according to
$\pi^{*}H^{c}:=\frac{H}{\rho^{\Lambda}}$ (54)
Elsewhere, i.e on $C\hskip 1.70717pt\backslash\pi(H=0)$, the appropriate function is $|H^{c}|^{1/\Lambda}$ [14]. If $X_{H^{c}}$ denotes the Hamiltonian vector field of the contact Hamiltonian (54), and $\mathcal{R}$ is the Reeb field of $\eta$, then the vector field
$X:=X_{H^{c}}+(\Lambda-1)H^{c}\,\mathcal{R}$ (55)
generates $\pi_{*}\,\textrm{span}(X_{H})$. In general, a contact vector field is one whose flow preserves the contact distribution; consequently, unless the scaling symmetry of interest is of degree one (which, throughout subsequent discussion, will often be the case) or we restrict ourselves to the zero-set of $H^{c}$, the vector field $X$ does not generally preserve $\xi=\textrm{ker}\,\eta$. Additionally, it is clear that this construction is dependent upon the relation $\mathfrak{L}_{D}X_{H}=(\Lambda-1)X_{H}$, which itself derives from the symplectic Hamiltonian equation of motion $\iota_{X_{H}}\omega=dH$; as a result, scaling symmetries are strictly dynamical in nature: that is to say, the contact-reduced system faithfully reproduces the original symplectic theory only when the equations of motion are satisfied.
In general, the presence of the scaling function $\rho$ introduces a reparameterisation of the temporal coordinate on $C$; in particular, for a scaling symmetry of degree $\Lambda$, trajectories on the symplectic phase space, parameterised by the coordinate $t$, are projected to curves, governed by changes in $\tau$, with $d\tau=\rho^{\Lambda-1}dt$. Accordingly, if $(q^{i}(\tau),p_{i}(\tau),S(\tau))$ describes an integral curve of (55), the equations of motion read
$\frac{dq^{i}}{d\tau}=\frac{\partial H^{c}}{\partial p_{i}}\quad\quad\frac{dp_{i}}{d\tau}=-\left(\frac{\partial H^{c}}{\partial q^{i}}+p_{i}\frac{\partial H^{c}}{\partial S}\right)\quad\quad\frac{dS}{d\tau}=p_{i}\frac{\partial H^{c}}{\partial p_{i}}-\Lambda H^{c}$ (56)
This concludes our introductory treatment of contact reduction, and we now apply these ideas to constrained Hamiltonian systems. Before commencing, we remark that there are two conceivable ways of proceeding: on the one hand, given a Hamiltonian $H_{0}$, corresponding to a singular Lagrangian, we may apply the techniques of section (III.1), deduce the final constraint manifold, eliminating any gauge degrees of freedom, and then seek a scaling symmetry, to make the contact reduction. Alternatively, we might first identify a scaling degree of freedom within the unconstrained Hamiltonian system, excise this from our ontology, and then use the pre-contact constraint algorithm, to deduce the subset of the reduced space, upon which solutions to the contact equations of motion are well-defined. More succinctly, in what follows, we would like to examine the commutativity of ‘restriction &\& reduction’ and ‘reduction &\& restriction.’
VII.2 Reduction &\& Restriction
Let us suppose that $(TQ,\eta_{L},E_{L})$ is an almost-regular Lagrangian system; we begin by deducing the canonical Hamiltonian $H_{0}:M_{0}\rightarrow\mathbb{R}$, defined on the primary constraint manifold $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$. This is then extended to a function $H:T^{*}Q\rightarrow\mathbb{R}$ on the full phase space, such that $H|_{M_{0}}=H_{0}$. [^4] In practice, $H$ and $H_{0}$ have identical coordinate expressions; however, formally, they belong to different spaces. The cotangent bundle $T^{*}Q$ is a symplectic manifold, whose corresponding 2-form is expressed in local Darboux coordinates as $\omega=dq^{i}\wedge dp_{i}$; supposing that $M_{0}$ is described as the zero-set of the primary constraint functions $\phi^{\alpha}$, we seek a vector field $D\in\mathfrak{X}^{\infty}(T^{*}Q)$ which satisfies
$\mathfrak{L}_{D}H=\Lambda H\quad\quad\quad\mathfrak{L}_{D}\hskip 0.56905pt\omega=\omega\quad\quad\quad\mathfrak{L}_{D}\phi^{\alpha}=C^{\alpha}_{\,\beta}\phi^{\beta}$ (57)
for a set of functions $C^{\alpha}_{\,\beta}\in C^{\infty}(T^{*}Q)$, with $\textrm{det}\,C^{\alpha}_{\;\beta}\neq 0$. The first two conditions simply identify $D$ as a scaling symmetry, while the third ensures that the flow of $D$ maps $M_{0}$ onto itself. This is essential, as $M_{0}$ describes the maximal subset of points at which solutions to the equations of motion could exist (evidently, implementation of the constraint algorithm is likely to restrict $M_{0}$ further, but any such restriction will necessarily be a subset of $M_{0}$); the flow of the scaling field must not, therefore, map points of $M_{0}$, which are (potentially) dynamically admissible, to points of $T^{*}Q\hskip 2.27621pt\backslash M_{0}$, which are not.
In principle, we should now verify that the flow of $D$ does in fact act freely and properly on $T^{*}Q$; supposing that this is the case, the space $C:=T^{*}Q/\hskip-2.84526pt\sim$ is a contact manifold with submersion $\beta:T^{*}Q\rightarrow C$, and contact distribution $\xi:=\beta_{*}\,\textrm{ker}\left(\iota_{D}\hskip 0.56905pt\omega\right)$. We have a contact form $\eta$ and Hamiltonian $H^{c}$ defined precisely as in (53) and (54) respectively, with the result that $(C,\eta,H^{c})$ constitutes a contact Hamiltonian system.
We highlight that this construction does not directly give rise to a pre-contact manifold; instead, we have symmetry reduced the full phase space, into which the pre-symplectic system is embedded. Had we sought a vector field on $M_{0}$ as our scaling symmetry, it is not clear that we should always obtain a well-defined pre-contact form on the quotient space. By working on the full phase space, we have the map $\beta:T^{*}Q\rightarrow C$, which allows us to construct functions $\gamma^{\alpha}\in C^{\infty}(C)$, whose zero-set defines a submanifold $\kappa_{0}:C_{0}\hookrightarrow C$, which does inherit a pre-contact structure; in particular, we have
$C_{0}:=\{y\in C\;|\;\gamma^{\alpha}(y)=0\,\}\quad\quad\textrm{with}\quad\beta^{*}\gamma^{\alpha}:=\frac{\phi^{\alpha}}{\rho^{\Lambda}}$ (58)
The contact Hamiltonian $H^{c}$ is restricted to a function $H^{c}_{0}:=H^{c}|_{C_{0}}$, and similarly, $\eta_{0}:=\eta|_{C_{0}}$ defines a pre-contact form on $C_{0}$, with the result that the triple $(C_{0},\eta_{0},H^{c}_{0})$ is a well-defined pre-contact manifold, to which the constraint algorithm of section (VI) may freely be applied. Supposing that the algorithm stabilises on the final constraint surface $C_{f}$, we eliminate any gauge degrees of freedom, denoting physical state space $C_{\scriptscriptstyle\mathcal{P}}$, with $\kappa_{\scriptscriptstyle\mathcal{P}}:C_{\scriptscriptstyle\mathcal{P}}\hookrightarrow C_{f}$. On $C_{\scriptscriptstyle\mathcal{P}}$, we have a well-defined contact form $\eta_{\scriptscriptstyle\mathcal{P}}$, and Hamiltonian $H^{c}_{\scriptscriptstyle\mathcal{P}}$.
VII.3 Restriction &\& Reduction
Having shown how we obtain the contact Hamiltonian equations of motion by first contact reducing the pre-symplectic system, and then applying the appropriate constraint algorithm, we shall now consider the reverse process, in which we first carry out the restriction at the symplectic level, before seeking a scaling symmetry with which to contact reduce the constrained theory.
We begin, once again, from the pre-symplectic system $(M_{0},\omega_{0},H_{0})$, and carry out the constraint procedure presented in section (III.1), deducing the final submanifold $M_{f}\hookrightarrow T^{*}Q$. In order to be able to carry out a contact reduction via scaling symmetry, we must either introduce a particular choice of gauge-fixing, or formally quotient out by the action of the gauge transformations; since we are concerned with the dynamics of observable ($i.e$ gauge-invariant) quantities, this choice is somewhat inconsequential, and we denote the physical phase space $\mathcal{P}$, with $\jmath_{\scriptscriptstyle\mathcal{P}}:\mathcal{P}\hookrightarrow M_{f}$.
From above, we know that $\mathcal{P}$ is symplectic, with non-degenerate 2-form $\omega_{\scriptscriptstyle\mathcal{P}}$ and Hamiltonian $H_{\scriptscriptstyle\mathcal{P}}$. As such, we now seek a vector field $Z\in\mathfrak{X}^{\infty}(\mathcal{P})$, such that
$\mathfrak{L}_{Z}\omega_{\scriptscriptstyle\mathcal{P}}=\omega_{\scriptscriptstyle\mathcal{P}}\quad\quad\quad\mathfrak{L}_{Z}H_{\scriptscriptstyle\mathcal{P}}=\Lambda H_{\scriptscriptstyle\mathcal{P}}$ (59)
Note that, by construction, the constraint algorithm produces a surface to which all dynamics must remain tangent; thus, we need impose no conditions on the preservation of the constraints. The dynamical evolution of some $f\in C^{\infty}(\mathcal{P})$ is determined via the bracket induced by $\omega_{\scriptscriptstyle\mathcal{P}}$, that is
$\dot{f}=\{f,H_{\scriptscriptstyle\mathcal{P}}\}_{\scriptscriptstyle\mathcal{P}}:=\omega_{\scriptscriptstyle\mathcal{P}}(X_{f},X_{H_{\scriptscriptstyle\mathcal{P}}})$
in which $X_{f}$ refers to the Hamiltonian vector field associated to $f$ via $\iota_{X_{f}}\omega_{\scriptscriptstyle\mathcal{P}}=df$, and similarly for $X_{H_{\scriptscriptstyle\mathcal{P}}}$.
Assuming that the flow of $Z$ acts freely and properly on $\mathcal{P}$, we have a well-defined quotient space $\mathcal{P}/\hskip-2.84526pt\sim$, with submersion $\sigma:\mathcal{P}\rightarrow\mathcal{P}/\hskip-2.84526pt\sim\,$; if $\rho:\mathcal{P}\rightarrow\mathbb{R}_{+}$ is a global scaling function on $\mathcal{P}$, the contact form and Hamiltonian are given by
$\sigma^{*}\eta_{\scriptscriptstyle\mathcal{P}}=\frac{\iota_{Z}\omega_{\scriptscriptstyle\mathcal{P}}}{\rho}\quad\quad\quad\sigma^{*}H_{\scriptscriptstyle\mathcal{P}}^{c}=\frac{H_{\scriptscriptstyle\mathcal{P}}}{\rho^{\Lambda}}$ (60)
Supposing that $(q^{i},\Pi_{i},S)$ constitutes a set of local Darboux coordinates on $\mathcal{P}/\hskip-2.84526pt\sim$, so that $\eta_{\scriptscriptstyle\mathcal{P}}=dS-\Pi_{i}dq^{i}$, the symmetry-reduced equations of motion are simply the contact Hamiltonian equations (56).
From the constructions given above, it should be clear that the pre-symplectic and pre-contact constraint algorithms run in parallel. At the $i^{\textrm{th}}$ stage of the pre-symplectic algorithm, the submanifold $M_{i}$, with inclusion $\jmath_{i}:M_{i}\hookrightarrow M_{i-1}$, may be described as the zero-set of the functions $\phi^{\alpha}_{(i)}$. The projection of these $\phi^{\alpha}_{(i)}$ defines a second set of functions $\gamma^{\alpha}_{(i)}$ according to
$\beta^{*}\gamma^{\alpha}_{(i)}:=\frac{\phi^{\alpha}_{(i)}}{\rho^{\Lambda}}$ (61)
Up to multiplicative factors, these $\gamma^{\alpha}_{(i)}$ coincide with the constraint functions obtained in $i^{\textrm{th}}$ iteration of the pre-contact algorithm. These ideas are summarised in figure 1.
$(T^{*}Q,\omega,H)$$(C,\eta,H^{c})$$(M_{0},\omega_{0},H_{0})$$(C_{0},\eta_{0},H^{c}_{0})$$M_{1}$$\vdots$$\vdots$$\vdots$$M_{f}$$C_{f}$$(\mathcal{P},\omega_{\scriptscriptstyle\mathcal{P}},H_{\scriptscriptstyle\mathcal{P}})$$(C_{\scriptscriptstyle\mathcal{P}},\eta_{\scriptscriptstyle\mathcal{P}},H^{c}_{\scriptscriptstyle\mathcal{P}})$$\scriptstyle{\beta}$$\scriptstyle{\jmath_{0}}$$\scriptstyle{\kappa_{0}}$$\scriptstyle{\jmath_{1}}$$\scriptstyle{\kappa_{1}}$$\scriptstyle{\jmath_{2}}$$\scriptstyle{\jmath_{f}}$$\scriptstyle{\kappa_{f}}$$\scriptstyle{\jmath_{\scriptscriptstyle\mathcal{P}}}$$\scriptstyle{\sigma}$$\scriptstyle{\kappa_{\scriptscriptstyle\mathcal{P}}}$
Figure 1: Commutative diagram showing how the two constraint algorithms run in parallel. Starting from the cotangent bundle, with its canonical symplectic form, the right-hand path shows how we may restrict to the primary constraint manifold $M_{0}$, and carry out the pre-symplectic algorithm, to reach the final space $\mathcal{P}$. This manifold is symplectic, and thus allows a contact reduction to be carried out. Alternatively, the scaling degree of freedom may be excised from the full symplectic phase space, defining the contact manifold $C$. Upon projecting the primary constraint functions that define $M_{0}$, $C$ is restricted to the pre-contact space $C_{0}$, and the left-hand path illustrates the implementation of the pre-contact algorithm. Both paths lead to the same final reduced space $C_{\scriptscriptstyle\mathcal{P}}$.
When applying a restriction of the phase space (prior to contact reduction) we arrive at a symplectic manifold $\mathcal{P}$, which, by construction, contains all points at which the equations of motion admit vector field solutions. We know that the reduced theory faithfully reproduces the original symplectic dynamics only when the equations of motion are satisfied. Since the constraint procedure systematically eliminates all points which are not dynamically-admissible, we expect that the contact-reduced dynamics on $C_{\scriptscriptstyle\mathcal{P}}$ coincide with that of the constrained symplectic system at all points.
VIII An Example
We shall now apply the mathematical formalism we have developed throughout to a simple example; for completeness, we obtain the reduced space dynamics via both paths of 1, thereby providing an explicit illustration of the commutativity of the reduction and constraint processes.
VIII.1 Restriction &\& Reduction
For this particular example, we have a configuration manifold $Q=\mathbb{R}_{+}^{4}$, with local coordinates $(x,y,u,\phi)$, and Lagrangian function
$L=2\dot{u}^{2}+\frac{1}{2}u^{2}\biggr[\dot{\phi}^{2}+(\dot{x}+\dot{y}-\dot{\phi})^{2}-(x^{2}+y^{2}-2\phi^{2})\biggr]$ (62)
Our motivation for choosing this particular Lagrangian is purely pedagogical; we do not consider this to represent any realistic physical system. Clearly, the Lagrangian is singular, as the Hessian matrix with respect to the velocities $(\dot{x},\dot{y},\dot{u},\dot{\phi})$ is
$W=u^{2}\begin{pmatrix}1&1&0&-1\\ 1&1&0&-1\\ 0&0&4u^{-2}&0\\ -1&-1&0&2\end{pmatrix}$
which is of constant rank 3, provided $u\neq 0$, which we have assumed. This Lagrangian admits a fairly simple scaling symmetry of the form
$\Sigma=\frac{1}{2}\left(u\frac{\partial}{\partial u}+\dot{u}\frac{\partial}{\partial\dot{u}}\right)$ (63)
Making the change of variable $u\mapsto e^{\rho/2}$, the scaling symmetry becomes $\Sigma=\partial_{\rho}$, while the Lagrangian adopts the form
$L=e^{\rho}\biggr[\frac{1}{2}\left(\dot{\rho}^{2}+\dot{\phi}^{2}\right)+\frac{1}{2}\left(\dot{x}+\dot{y}-\dot{\phi}\right)^{2}-\frac{1}{2}\left(x^{2}+y^{2}\right)+\phi^{2}\biggr]$ (64)
This is the most convenient parameterisation of $L$, and the one with which we shall henceforth work; the momenta conjugate to $(\dot{x},\dot{y},\dot{\rho},\dot{\phi})$ are given by
$p_{\rho}$ $=e^{\rho}\dot{\rho}$ $p_{x}$ $=e^{\rho}\left(\dot{x}+\dot{y}-\dot{\phi}\right)$
$p_{\phi}$ $=e^{\rho}\left(2\dot{\phi}-(\dot{x}+\dot{y})\right)$ $p_{y}$ $=e^{\rho}\left(\dot{x}+\dot{y}-\dot{\phi}\right)$
From this, we see that there is a single primary constraint, defining the submanifold $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$
$M_{0}:=\{p\in T^{*}Q\;|\;\psi^{1}(p):=p_{x}-p_{y}=0\,\}$ (65)
The canonical Hamiltonian $H_{0}:M_{0}\rightarrow\mathbb{R}$ is readily calculated to be
$H_{0}=\frac{e^{-\rho}}{2}\left[p_{\rho}^{2}+\left(p_{x}+p_{\phi}\right)^{2}+p_{x}^{2}\right]+e^{\rho}\left[\frac{1}{2}(x^{2}+y^{2})-\phi^{2}\right]$ (66)
Pulling back the canonical symplectic form $\omega$ on $T^{*}Q$ to $M_{0}$, we have $\omega_{0}:=\jmath^{*}\omega\in\Omega^{2}(M_{0})$, which, in local coordinates, is given by
$\omega_{0}=d\rho\wedge dp_{\rho}+d\phi\wedge dp_{\phi}+(dx+dy)\wedge dp_{x}$ (67)
In accordance with the pre-symplectic constraint algorithm developed in section (III.1), we know that the symplectic orthogonal $TM_{0}^{\perp}$ coincides precisely with $\textrm{ker}\,\omega_{0}$. In order for the vector field
$X=A^{x}\frac{\partial}{\partial x}+A^{y}\frac{\partial}{\partial y}+A^{\rho}\frac{\partial}{\partial\rho}+A^{\phi}\frac{\partial}{\partial\phi}+B^{x}\frac{\partial}{\partial p_{x}}+B^{\rho}\frac{\partial}{\partial p_{\rho}}+B^{\phi}\frac{\partial}{\partial p_{\phi}}$
to belong to $\textrm{ker}\,\omega_{0}$, we require $A^{x}+A^{y}=0$ and that the remaining coefficients be zero; consequently, we have
$\textrm{ker}\,\omega_{0}=TM_{0}^{\perp}\,=\;\biggr\langle\frac{\partial}{\partial x}-\frac{\partial}{\partial y}\biggr\rangle$ (68)
The secondary constraint manifold $M_{1}$ is composed of those points $p\in M_{0}$, at which $\langle(dH_{0})_{p},T_{p}M_{0}^{\perp}\rangle=0$; from (66), we see that this gives rise to a single constraint [^5] Our inclusion of the exponential factor in $\psi^{2}$ may seem somewhat pedantic, since clearly $e^{\rho}\neq 0$; however, its presence is highly relevant when considering the reduction process.
$M_{1}=\{p\in M_{0}\;|\;\psi^{2}(p):=e^{\rho}\left(x-y\right)=0\,\}$ (69)
Any vector field solution $X_{H}$ to the geometrical equation $\iota_{X_{H}}\omega_{0}-dH_{0}=0$, restricted to $M_{1}$, must remain tangent to this surface, in order to be considered physically meaningful. As we have seen, this requirement is entirely equivalent to imposing that $\langle(dH_{0})_{p},T_{p}M_{1}^{\perp}\rangle=0$ for all points $p\in M_{1}$. It is straightforward to verify that
$TM_{1}^{\perp}\,=\hskip 4.2679pt\biggr\langle\frac{\partial}{\partial x}-\frac{\partial}{\partial y}\biggr\rangle$
Consequently, the algorithm stabilises on $M_{1}$, with the two constraints $\psi^{1}$ and $\psi^{2}$. In accordance with our general procedure, we introduce the matrix of Poisson brackets $J^{\alpha\beta}:=\{\psi^{\alpha},\psi^{\beta}\}$, which is given by
$J=e^{\rho}\begin{pmatrix}0&-2\\ 2&0\end{pmatrix}$
Clearly, this is invertible, which indicates that $\psi^{1}$ and $\psi^{2}$ constitute a pair of second class constraints; it is therefore unnecessary to eliminate gauge degrees of freedom, and we henceforth refer to $M_{1}$ as $\mathcal{P}$. The inverse $J^{-1}$ is given by
$J^{-1}=\frac{e^{-\rho}}{2}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$
and so the Dirac bracket of two functions $f,g\in C^{\infty}(T^{*}Q)$ is
$\{f,g\}_{D}=\{f,g\}-\frac{e^{-\rho}}{2}\{f,\psi^{1}\}\{\psi^{2},g\}+\frac{e^{-\rho}}{2}\{f,\psi^{2}\}\{\psi^{1},g\}$ (70)
In order to deduce the dynamical evolution of phase space functions, we should construct the total Hamiltonian; however, having identified $\psi^{1}$ and $\psi^{2}$ as second class, it follows that the total Hamiltonian coincides with $H_{0}$. It is then straightforward to show that the equations of motion for our phase space variables, which follow from $\dot{f}=\{f,H_{0}\}_{D}|_{\scriptscriptstyle\mathcal{P}}$ are
$\dot{x}$ $=\frac{e^{-\rho}}{2}\left(2p_{x}+p_{\phi}\right)=\dot{y}$ $\dot{p}_{x}$ $=-e^{\rho}x=\dot{p}_{y}$
$\dot{\phi}$ $=e^{-\rho}\left(p_{x}+p_{\phi}\right)$ $\dot{p}_{\phi}$ $=2e^{\rho}\phi$ (71)
$\dot{\rho}$ $=e^{-\rho}p_{\rho}$ $\dot{p}_{\rho}$ $=\frac{e^{-\rho}}{2}\left(p_{\rho}^{2}+\left(p_{x}+p_{\phi}\right)^{2}+p_{x}^{2}\right)-e^{\rho}\left(x^{2}-\phi^{2}\right)$
These dynamical equations must be reproduced by the symmetry-reduced model, else our claim that the excision of a scaling variable is inconsequential for the evolution of observable quatities is unjustified. Turning now to the task of reducing this system, we note that the symplectic form on $\mathcal{P}$ is
$\omega_{\scriptscriptstyle\mathcal{P}}:=\omega_{0}|_{\scriptscriptstyle\mathcal{P}}=d\rho\wedge dp_{\rho}+d\phi\wedge dp_{\phi}+2\,dx\wedge dp_{x}$
We also have the following restricted Hamiltonian
$H_{\scriptscriptstyle\mathcal{P}}:=H_{0}|_{\scriptscriptstyle\mathcal{P}}=\frac{e^{-\rho}}{2}\left[p_{\rho}^{2}+\left(p_{x}+p_{\phi}\right)^{2}+p_{x}^{2}\right]+e^{\rho}\left[x^{2}-\phi^{2}\right]$
In accordance with the results of section (VII.3), we now seek a vector field $Z\in\mathfrak{X}^{\infty}(\mathcal{P})$ that satisfies $\mathfrak{L}_{Z}\hskip 0.85358pt\omega_{\scriptscriptstyle\mathcal{P}}=\omega_{\scriptscriptstyle\mathcal{P}}$, and $\mathfrak{L}_{Z}\hskip 0.56905ptH_{\scriptscriptstyle\mathcal{P}}=\Lambda H_{\scriptscriptstyle\mathcal{P}}$; a series of short calculations confirms that
$Z=\frac{\partial}{\partial\rho}+p_{\rho}\frac{\partial}{\partial p_{\rho}}+p_{\phi}\frac{\partial}{\partial p_{\phi}}+p_{x}\frac{\partial}{\partial p_{x}}$ (72)
does indeed satisfy both of these requirements, and is a scaling symmetry of degree one. An obvious choice of scaling function is $e^{\rho}$; thus, we introduce the map $\sigma:\mathcal{P}\rightarrow\mathcal{P}/\hskip-2.84526pt\sim$, and construct the 1-form and contact Hamiltonian according to
$\sigma^{*}\eta_{\scriptscriptstyle\mathcal{P}}=\frac{\iota_{Z}\hskip 0.56905pt\omega_{\scriptscriptstyle\mathcal{P}}}{e^{\rho}}\quad\quad\sigma^{*}H_{\scriptscriptstyle\mathcal{P}}^{c}=\frac{H_{\scriptscriptstyle\mathcal{P}}}{e^{\rho}}$
We take coordinates on $\mathcal{P}/\hskip-2.84526pt\sim$ to be $(x,\Pi_{x},\phi,\Pi_{\phi},S)$, where
$\sigma^{*}\Pi_{x}=\frac{p_{x}}{e^{\rho}}\quad\quad\sigma^{*}\Pi_{\phi}=\frac{p_{\phi}}{e^{\rho}}\quad\quad\sigma^{*}S=\frac{p_{\rho}}{e^{\rho}}$ (73)
It then follows that $\eta_{\scriptscriptstyle\mathcal{P}}$ and $H_{\scriptscriptstyle\mathcal{P}}^{c}$ have the following local coordinate expressions
$\eta_{\scriptscriptstyle\mathcal{P}}=dS-\Pi_{\phi}d\phi-2\hskip 0.56905pt\Pi_{x}dx\quad\quad\quad H_{\scriptscriptstyle\mathcal{P}}^{c}=\frac{1}{2}\biggr[S^{2}+\left(\Pi_{x}+\Pi_{\phi}\right)^{2}+\Pi_{x}^{2}\biggr]+x^{2}-\phi^{2}$ (74)
The contact Hamiltonian equations of motion are then deduced in the standard fashion; however, it should be noted that the multiplicative factor of two in our expression for $\eta_{\scriptscriptstyle\mathcal{P}}$ shows that $(x,2\Pi_{x},\phi,\Pi_{\phi},S)$ provides the correct set of local Darboux coordinates on $\mathcal{P}/\hskip-2.84526pt\sim$. Consequently, we have
$\dot{x}$ $=\frac{1}{2}\left(2\Pi_{x}+\Pi_{\phi}\right)$ $\dot{\Pi}_{x}$ $=-\left(2x+\Pi_{x}S\right)$
$\dot{\phi}$ $=\Pi_{x}+\Pi_{\phi}$ $\dot{\Pi}_{\phi}$ $=2\phi-\Pi_{\phi}S$ (75)
$\dot{S}$ $=\frac{1}{2}\biggr[-S^{2}+\left(\Pi_{x}+\Pi_{\phi}\right)^{2}\biggr]-x^{2}+\phi^{2}$
It is clear that the equations of motion for the coordinates $x$ and $\phi$ coincide with those given in (VIII.1); after some work, it can be shown that the additional action-dependent terms in the momentum equations correctly reproduce the original dynamics.
VIII.2 Reduction + Restriction
In order to obtain the contact Hamiltonian equations (VIII.1) via the alternative method presented in section (VII.2), we return to the canonical Hamiltonian $H_{0}$ on the primary constraint manifold; in accordance with our general procedure, we extend $H_{0}$ to a function $H$ on the full phase space $T^{*}Q$
$H=\frac{e^{-\rho}}{2}\left[p_{\rho}^{2}+\left(p_{x}+p_{\phi}\right)^{2}+p_{x}^{2}\right]+e^{\rho}\left[\frac{1}{2}(x^{2}+y^{2})-\phi^{2}\right]$ (76)
The vector field
$D=\frac{\partial}{\partial\rho}+p_{\rho}\frac{\partial}{\partial p_{\rho}}+p_{\phi}\frac{\partial}{\partial p_{\phi}}+p_{x}\frac{\partial}{\partial p_{x}}+p_{y}\frac{\partial}{\partial p_{y}}$ (77)
is a scaling symmetry of degree one, and, crucially, preserves the primary constraint $\psi^{1}$, since $\mathfrak{L}_{D}\psi^{1}=\psi^{1}$. Thus, we know that the quotient space $C:=T^{*}Q/\hskip-2.84526pt\sim$ is a contact manifold, and we introduce the surjective map $\beta:T^{*}Q\rightarrow C$, and take coordinates on $C$ to be $(x,\pi_{x},y,\pi_{y},\phi,\pi_{\phi},S)$, where
$\beta^{*}\pi_{x}=\frac{p_{x}}{e^{\rho}}\quad\quad\quad\beta^{*}\pi_{y}=\frac{p_{y}}{e^{\rho}}\quad\quad\quad\beta^{*}\pi_{\phi}=\frac{p_{\phi}}{e^{\rho}}\quad\quad\quad\beta^{*}S=\frac{p_{\rho}}{e^{\rho}}$ (78)
so that the contact form and Hamiltonian on $C$ become
$\eta=dS-\pi_{x}dx-\pi_{y}dy-\pi_{\phi}d\phi\quad\quad\quad H^{c}=\frac{1}{2}\biggr[S^{2}+\left(\pi_{x}+\pi_{\phi}\right)^{2}+\pi_{x}^{2}\biggr]+\frac{1}{2}\left(x^{2}+y^{2}\right)-\phi^{2}$
We emphasise that $(C,\eta,H^{c})$ is a contact Hamiltonian system, and that the symmetry-reduced analogue of the primary constraint manifold $M_{0}$ is found by projecting the function $\psi^{1}$, to obtain
$\beta^{*}\gamma^{1}=\frac{\psi^{1}}{e^{\rho}}\quad\implies\quad\gamma^{1}=\pi_{x}+\pi_{y}$
Consequently, the primary constraint submanifold $C_{0}$ is precisely the zero-set of $\gamma^{1}$, and the pre-contact system $(C_{0},\eta_{0},H^{c}_{0})$ is obtained, restricting $\eta$ and $H^{c}$ to this space
$\eta_{0}:=\eta|_{C_{0}}=dS-\pi_{x}\left(dx+dy\right)-\pi_{\phi}d\phi\quad\quad\quad H^{c}_{0}:=H^{c}|_{C_{0}}=\frac{1}{2}\biggr[S^{2}+\left(\pi_{x}+\pi_{\phi}\right)^{2}+\pi_{x}^{2}\biggr]+\frac{1}{2}\left(x^{2}+y^{2}\right)-\phi^{2}$
We now apply the pre-contact constraint algorithm developed in section (VI), which begins with the deduction of the characteristic distribution $\mathcal{C}$. From above, we see that $d\eta_{0}=(dx+dy)\wedge d\pi_{x}+d\phi\wedge d\pi_{\phi}$, whence it is straightforward to show that
$\mathcal{C}:=\textrm{ker}\,\eta_{0}\;\cap\;\textrm{ker}\,d\eta_{0}\;=\;\,\biggr\langle\frac{\partial}{\partial x}-\frac{\partial}{\partial y}\biggr\rangle$
The first iteration of the constraint algorithm requires us to restrict $C_{0}$ to the subset of points at which the covector $\alpha_{0}:=dH^{c}_{0}-\left(\mathcal{R}(H^{c}_{0})+H^{c}_{0}\right)\eta_{0}$ is in the annihilator of $TC_{0}^{\perp}$, which, we observed, was equivalent to imposing that $\langle\alpha_{0},\mathcal{C}\rangle=0$. We take, as our Reeb field $\mathcal{R}=\partial/\partial S$, and find that the condition $\langle\alpha_{0},\mathcal{C}\rangle=0$ gives rise to a single secondary constraint
$C_{1}=\{p\in C_{0}\;|\;\gamma^{2}(p):=x-y=0\,\}$
As expected, comparing this space to $M_{1}$ found in (69), we see that the constraint $\gamma^{2}$ satisfies
$\beta^{*}\gamma^{2}=\frac{\psi^{2}}{e^{\rho}}$
Upon calculating $TC_{1}^{\perp}$, we find that imposing $\langle\alpha_{0},TC_{1}^{\perp}\rangle=0$ produces no additional constraints, and so the algorithm stabilises on $C_{1}$. Thus, we compute the matrix of Jacobi brackets $K^{\alpha\beta}:=\{\gamma^{\alpha},\gamma^{\beta}\}_{J}$, and referring to (37), we find that
$K=\begin{pmatrix}0&-2\\ 2&0\end{pmatrix}\quad\quad\implies\quad\quad K^{-1}=\frac{1}{2}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$
The invertibility of this matrix confirms that, as expected, $\gamma^{1}$ and $\gamma^{2}$ are both second class constraints; thus, in order to calculate the equations of motion, we impose such conditions as strong equalities, and introduce the Dirac-Jacobi bracket, which reads
$\{f,g\}_{DJ}=\{f,g\}_{J}-\frac{1}{2}\{f,\gamma^{1}\}_{J}\,\{\gamma^{2},g\}_{J}+\frac{1}{2}\{f,\gamma^{2}\}_{J}\,\{\gamma^{1},g\}_{J}$ (79)
The evolution of a function $f$ is then computed from $\{f,H^{c}_{0}\}_{DJ}\,|_{\scriptscriptstyle C_{1}}$, and we find that
$\dot{x}$ $=\frac{1}{2}\left(2\pi_{x}+\pi_{\phi}\right)$ $\dot{\pi}_{x}$ $=-\left(2x+\pi_{x}S\right)$
$\dot{\phi}$ $=\pi_{x}+\pi_{\phi}$ $\dot{\pi}_{\phi}$ $=2\phi-\pi_{\phi}S$ (80)
$\dot{S}$ $=\frac{1}{2}\biggr[-S^{2}+\left(\pi_{x}+\pi_{\phi}\right)^{2}\biggr]-x^{2}+\phi^{2}$
which coincides precisely with the corresponding set of expressions (VIII.1), deduced by first restricting the pre-symplectic system, and then making a symmetry reduction. Additionally, since $\mathcal{P}$ was, by construction, the maximal set of points at which the solutions (VIII.1) are well-defined, we are assured that both (VIII.1) and (VIII.2) faithfully reproduce the original dynamics at all points of the reduced space.
In this example, we have provided an explicit illustration of how a constrained Hamiltonian system may be reduced to a simpler, dynamically-equivalent theory, by identifying and excising an unphysical scaling degree of freedom. Whilst relatively simple in nature, the system considered has allowed us to present a fully-worked example of the commutativity of contact reduction and phase space restriction, and thus concludes our treatment of particle dynamics. In the following sections, we introduce the geometrical framework required to describe singular field theories, which will then allow us to provide a field-theoretic discussion, which parallels the ideas presented thus far.
IX Multisymplectic Field Theory
The use of multisymplectic geometry to describe classical field theories is an area of active research interest, as the fibred manifolds it employs provide an arena in which a manifestly covariant formalism may be developed, in a finite-dimensional setting [83, 84, 85]. Scaling symmetries of classical field theories have been studied in [34], and a mathematical framework for the contact reduction of multisymplectic theories developed in [29]. The multisymplectic Hamiltonian formalism of singular field theories is still not entirely understood; for example, it is known that there exist certain classes of Lagrangian, for which the construction of the corresponding Hamiltonian is either ambiguous, or simply ill-defined [16]. Additionally, multisymplectic manifolds are not, in general, equipped with a well-defined bracket structure [86], and the local constraint algorithm employs geometrical principles, that render the use of such structures redundant [21, 22]. As a result, in contrast to our study of systems of particles, our analysis of singular field theories will favour the Lagrangian formalism, and in what follows, we shall provide a heavily abridged summary of the most pertinent ideas, referring to [29] for more details.
IX.1 Lagrangian Field Theory
In general, an $m$-dimensional manifold $\mathcal{M}$ is said to be multisymplectic if it admits a closed, 1-non-degenerate $k$-form $\Omega\in\Omega^{k}(\mathcal{M})$ (with $1<k\leqslant m$) [16]; the 1-non-degeneracy condition may be expressed locally as the requirement that for every $p\in\mathcal{M}$ and $X_{p}\in T_{p}\mathcal{M}$
$\iota_{X_{p}}\Omega_{p}=0\quad\quad\implies\quad\quad X_{p}=0$
If $\Omega$ is closed but 1-degenerate, we refer to the pair $(\mathcal{M},\Omega)$ as a pre-multisymplectic manifold; such objects will be of particular relevance in our study of singular theories.
In order to formulate the field equations of a dynamical system in a geometrical fashion, we require a class of objects known as multivector fields. In general, a multivector field of degree $r$ on $\mathcal{M}$ is a section $\boldsymbol{X}\in\Gamma(\wedge^{r}T\mathcal{M})$ of the $r^{\textrm{th}}$ exterior power of the tangent bundle. We denote the space of all such multivector fields $\mathfrak{X}^{r}(\mathcal{M}):=\Gamma(\wedge^{r}\,T\mathcal{M})$, and declare some $\boldsymbol{X}\in\mathfrak{X}^{r}(\mathcal{M})$ to be locally decomposable if, for every point $p\in\mathcal{M}$, there exists an open neighbourhood $\mathcal{U}_{p}\subset\mathcal{M}$, and vector fields $X_{1},\,\cdots\,,X_{r}\in\mathfrak{X}^{\infty}(\mathcal{U}_{p})$, such that
$\boldsymbol{X}|_{\mathcal{U}_{p}}=X_{1}\wedge\,\cdots\,\wedge X_{r}$
We say that an $m$-dimensional distribution $\mathcal{D}\subset T\mathcal{M}$ is locally associated to a non-zero $\boldsymbol{X}\in\mathfrak{X}^{m}(\mathcal{M})$, if there exists some connected open set $\mathcal{V}\subset\mathcal{M}$, such that $\boldsymbol{X}|_{\mathcal{V}}$ is a section of $\wedge^{m}\mathcal{D}|_{\mathcal{V}}$; further, $\boldsymbol{X}$ is said to be integrable if its locally associated distribution is integrable.
Consider a fibre bundle $\pi:E\rightarrow M$ over the $d$-dimensional spacetime manifold $M$; we shall suppose that $M$ is orientable, with volume form $\omega$. Local coordinates on $M$ are written as $(x^{\mu})$, with $\mu=0,\,\cdots,d-1$, so that $\omega=dx^{0}\wedge\,\cdots\wedge dx^{d-1}:=d^{d}x$. The $(n+d)$-dimensional manifold $E$ is referred to as the covariant configuration space, and the first jet bundle $\kappa:J^{1}E\rightarrow E$ of sections of $\pi$ is the natural space upon which to introduce a Lagrangian density [87, 88]; local coordinates on $J^{1}E$, adapted to the bundle structure of $E$, are given by $(x^{\mu},y^{a},y^{a}_{\mu})$, with $a=1,\,\cdots,n$. Introducing the bundle projection $\widehat{\pi}:=\pi\circ\kappa:J^{1}E\rightarrow M$, the Lagrangian density $\mathcal{L}$ may be expressed as a $\widehat{\pi}-$semibasic $d$-form on $J^{1}E$
$\mathcal{L}(x^{\mu},y^{a},y^{a}_{\mu})=L(x^{\mu},y^{a},y^{a}_{\mu})\,\widehat{\pi}^{*}\omega$ (81)
in which $L:J^{1}E\rightarrow\mathbb{R}$ is referred to as the Lagrangian function, and $\widehat{\pi}^{*}\omega$ is the volume form on $M$, pulled back to $J^{1}E$ [29, 89]. The Lagrangian function is then used to define the Cartan forms $\Theta_{L}\in\Omega^{d}(J^{1}E)$ and $\Omega_{L}\in\Omega^{d+1}(J^{1}E)$; in local bundle coordinates $(x^{\mu},y^{a},y^{a}_{\mu})$, these are expressed as
$\Theta_{L}=\frac{\partial L}{\partial y^{a}_{\mu}}\,dy^{a}\wedge d^{d-1}x_{\mu}-\left(\frac{\partial L}{\partial y^{a}_{\mu}}\,y^{a}_{\mu}-L\right)\,d^{d}x\quad\quad\quad\quad\Omega_{L}:=-\,d\Theta_{L}$ (82)
where $d^{d-1}x_{\mu}:=\iota_{\scriptscriptstyle\partial_{\mu}}d^{d}x$. The pair $(J^{1}E,\Omega_{L})$ defines a Lagrangian system, which is said to be regular if $\Omega_{L}$ is multisymplectic, and singular if it is pre-multisymplectic [22]. The subset of those singular systems which are categorised as almost-regular satisfy the same conditions with respect to the Legendre map as those given in section (II) for systems of particles.
Given a regular Lagrangian system $(J^{1}E,\Omega_{L})$, the dynamical equations are derived from a variational principle [16], which defines critical sections $\phi\in\Gamma(M,E)$; the canonical lifting $j^{1}\phi$ of these objects to $J^{1}E$ are then integral sections of an equivalence class of locally decomposable, $\widehat{\pi}-$transverse holonomic multivector fields $\{\boldsymbol{X}_{L}\}$, each of which satisfies
$\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$ (83)
The most general expression for a locally decomposable field $\boldsymbol{X}_{L}$ is
$\boldsymbol{X}_{L}=\bigwedge\limits_{\mu=0}^{d-1}f\left(\frac{\partial}{\partial x^{\mu}}+F^{a}_{\mu}\frac{\partial}{\partial y^{a}}+G^{a}_{\mu\nu}\frac{\partial}{\partial y^{a}_{\nu}}\right)$ (84)
for some non-zero $f\in C^{\infty}(J^{1}E)$. The $\widehat{\pi}-$transversality condition is most readily enforced by setting $\iota_{\boldsymbol{X}_{L}}(\widehat{\pi}^{*}\omega)=1$, which fixes the multiplicative function $f$ to unity. When $\boldsymbol{X}_{L}$ is holonomic, it is integrable, and the coefficient functions $F^{a}_{\mu}$ are simply $y^{a}_{\mu}$; if $\boldsymbol{X}_{L}$ has local coordinate expression (84), with $F^{a}_{\mu}=y^{a}_{\mu}$, but is not integrable, it is referred to as semi-holonomic.
The critical sections $\phi$ of the variational problem are such that their canonical lifting $j^{1}\phi$ satisfy the Euler-Lagrange field equations
$\frac{\partial}{\partial x^{\mu}}\left(\frac{\partial L}{\partial y^{a}_{\mu}}\circ j^{1}\phi\right)-\frac{\partial L}{\partial y^{a}}\circ j^{1}\phi=0$ (85)
X Pre-multicontact Systems
In our treatment of scaling symmetries of pre-symplectic systems, we observed that the elimination of a scaling degree of freedom and phase space restriction were commutative processes; indeed, we dedicated significant effort to the construction of the final symmetry-reduced theory via both paths of the commutative diagram 1. When considering field theories, we shall not attempt to retain this same level of generality, and so will restrict our attention to the implementation of a constraint algorithm only after the scaling degree of freedom has been eliminated.
The symmetry reduction of multisymplectic field theories has been studied in detail in [29]; there, we presented a brief account of multicontact geometry, as pertained to the reduction of regular systems. When the theories of interest are singular, far more care must be taken, for it is not guaranteed that the symmetry-reduced form derived from the pre-multisymplectic $\Theta_{L}$ defines a pre-multicontact structure. Consequently, as an additional step, we must ensure that the reduced theory possesses a well-defined Reeb distribution; such considerations will require the introduction of additional mathematical structures, that were not required in [29]. Our discussion of multicontact geometry shall be limited to those concepts which are fundamental to the construction of classical field theories; more details may be found in [26, 25, 27], for example.
X.1 Multicontact Lagrangian Field Theory
Dissipative (or frictional) field theories are the subject of notably less research interest than their conservative counterparts; this is perhaps due to a tendency to wish to embed any such non-conservative configuration within a larger system, often denominated ‘the environment’. The frictional behaviour of the original system is then palliated by the observation that any apparent losses may simply be attributed to an exchange of energy between the system and environment. The philosophical stance which motivates our study of scaling symmetries is diametrically opposed to the notion of creating larger, redundant mathematical structures, with the sole objective of streamlining practical calculations [77, 80, 90]. We posit that the use of the smallest amount of dynamically-irrelevant structure, even at the expense of superficial mathematical simplicity, should be prioritised, wherever possible [91, 92]. In excising superfluous scaling degrees of freedom, seeking a minimal ontology, we are forced to work with theories which are inherently frictional in nature; consequently, we shall dedicate this section to developing the geometrical tools required to describe such systems.
The fibred manifolds $\pi:E\rightarrow M$, and $\kappa:J^{1}E\rightarrow E$ (with $\textrm{dim}\,M=d$ and $\textrm{dim}\,E=n+d$) of the multisymplectic formalism continue to assume a prominent role in the description of frictional field theories; in particular, the Lagrangian density is now a $d$-form on the manifold
$\mathcal{S}:=J^{1}E\times_{M}\wedge^{d-1}T^{*}M\cong J^{1}E\times\mathbb{R}^{d}$ (86)
which is a bundle over both $E$, with projection $\tau:\mathcal{S}\rightarrow E$, and $M$, with $\beta=\pi\circ\tau:\mathcal{S}\rightarrow M$. Local coordinates on $\mathcal{S}$ are denoted $(x^{\mu},y^{a},y^{a}_{\mu},s^{\mu})$, in which the quantities $s^{\mu}$ correspond to an action density. As in (81), we write the Lagrangian density in terms of a local function
$\mathcal{L}(x^{\mu},y^{a},y^{a}_{\mu},s^{\mu})=L(x^{\mu},y^{a},y^{a}_{\mu},s^{\mu})\,\beta^{*}\omega$ (87)
In contrast to the multisymplectic case, when the Lagrangian is singular, the $d$-form $\Theta_{L}$
$\Theta_{L}=\left(ds^{\mu}-\frac{\partial L}{\partial y^{a}_{\mu}}dy^{a}\right)\,\wedge\,d^{d-1}x_{\mu}+\left(\frac{\partial L}{\partial y^{a}_{\mu}}\,y^{a}_{\mu}-L\right)\,d^{d}x$ (88)
does not necessarily define a pre-multicontact structure. It is therefore necessary to introduce a number of additional concepts, which allow us to determine when a given Lagrangian $d$-form does in fact give rise to such a structure.
We begin by considering a generalisation of the annihilator of a given subspace, that we first encountered in section (II); for the purpose of defining the pre-multicontact structures of this section, in addition to the Lagrangian $d$-form $\Theta_{L}$, we also introduce the notation $\Xi:=\beta^{*}\omega\in\Omega^{d}(\mathcal{S})$; this volume form serves as a reference object, while $\Theta_{L}$ endows the space $\mathcal{S}$ with the formal geometrical structure of a pre-multicontact manifold. Given a regular distribution $\mathcal{D}\subset T\mathcal{S}$, we introduce the space of $r$-forms which annihilate all sections of $\mathcal{D}$
$\textrm{Ann}^{r}(\mathcal{D}):=\{\xi\in\Omega^{r}(\mathcal{S})\;|\;\iota_{X}\xi=0\quad\textrm{for all}\;\,X\in\Gamma(\mathcal{D})\,\}$ (89)
The Reeb distribution associated with $(\mathcal{S},\Theta_{L},\Xi)$ is then defined pointwise as
$\mathcal{D}^{R}_{p}:=\{X\in\textrm{ker}\,\Xi_{p}\;|\;\iota_{X}d\Theta_{L}\in\textrm{Ann}^{d}_{p}(\textrm{ker}\,\Xi)\,\}$ (90)
Sections of $\mathcal{D}^{R}$ are the Reeb vector fields; we denote this space $\mathfrak{R}:=\Gamma(\mathcal{D}^{R})$. In practice, we shall always work with coordinates on $\mathcal{S}$ adapted to the bundle structure $E\rightarrow M$, and so $\Xi=dx^{0}\wedge\,\cdots\,\wedge dx^{d-1}$ is a closed form, from which it follows that $\textrm{ker}\,\Xi$ consists of those vector fields whose vertical components $V^{\mu}\partial_{x^{\mu}}$ are vanishing. The space $\textrm{Ann}^{d}(\textrm{ker}\,\Xi)$ then comprises all $d$-forms $\xi\in\Omega^{d}(\mathcal{S})$, for which $\iota_{X}\xi=0$ for any choice of vector field of the form
$X=A^{a}\frac{\partial}{\partial y^{a}}+B^{a}_{\mu}\frac{\partial}{\partial y^{a}_{\mu}}+C^{\mu}\frac{\partial}{\partial s^{\mu}}$
In addition to the Reeb distribution, we also have the characteristic distribution $\mathcal{C}$, defined as the intersection $\mathcal{C}:=\textrm{ker}\,\Xi\cap\textrm{ker}\,\Theta_{L}\cap\textrm{ker}\,d\Theta_{L}$.
With these geometrical constructions in-hand, the conditions under which the triple $(\mathcal{S},\Theta_{L},\Xi)$ constitutes a pre-multicontact manifold are that, for some $0<k\leqslant n(1+d)$
  • $\star$
    $\textrm{rank\;ker}\,\Xi=d+n+nd$
  • $\star$
    $\textrm{rank}\,\mathcal{D}^{R}=d+k$
  • $\star$
    $\textrm{rank}\,\mathcal{C}=k$
  • $\star$
    $\textrm{Ann}^{d-1}(\textrm{ker}\,\Xi)=\{\iota_{R}\Theta_{L}\;|\;R\in\mathfrak{R}\,\}$
Note that we have excluded the possibility that $k=0$, for this is precisely the condition required for $(\mathcal{S},\Theta_{L},\Xi)$ to be multicontact, and not pre-multicontact. In practice, when working in adapted coordinates, the local Reeb fields are deduced from $\iota_{R_{\mu}}\Theta_{L}=d^{d-1}x_{\mu}$; however, observe that this does not determine the $R_{\mu}$ uniquely, for it is always possible to add to $R_{\mu}$ an element of the characteristic distribution, that is to say $\mathfrak{R}=\textrm{span}\hskip 1.13809pt(R_{\mu})+\mathcal{C}$.
Given an $r$-form $\xi\in\Omega^{r}(\mathcal{S})$, we introduce the following operator
$\begin{split}\bar{d}\;:\;\Omega^{r}(\mathcal{S})\;&\longrightarrow\;\Omega^{r+1}(\mathcal{S})\\ \xi\;&\longmapsto\;\bar{d}\xi:=d\xi+\sigma_{\Theta_{L}}\wedge\,\xi\end{split}$
where $\sigma_{\Theta_{L}}\in\Omega^{1}(\mathcal{S})$ is known as the dissipation form, and is expressed in local coordinates as
$\sigma_{\Theta_{L}}=-\,\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}$ (91)
The $(d+1)$-form $\Omega_{L}$ is then defined as
$\Omega_{L}=\bar{d}\Theta_{L}=d\Theta_{L}-\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}\wedge\,\Theta_{L}$ (92)
Given a (pre-)multicontact Lagrangian system $(\mathcal{S},\Theta_{L},\Xi)$, the equations of motion for sections $\Psi\in\Gamma(M,\mathcal{S})$ are given by
$\Psi^{*}\Theta_{L}=0\quad\quad\quad\textrm{and}\quad\quad\quad\Psi^{*}\iota_{Z}\Omega_{L}=0\quad\textrm{for all}\;Z\in\mathfrak{X}^{\infty}(\mathcal{S})$ (93)
We also have a similar pair equations for locally decomposable $\beta$-transverse multivector fields $\boldsymbol{X}_{L}\in\mathfrak{X}^{d}(\mathcal{S})$, which take the form
$\iota_{\boldsymbol{X}_{L}}\Theta_{L}=0\quad\quad\quad\quad\quad\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$ (94)
When $(\mathcal{S},\Theta_{L},\Xi)$ is multicontact, the multivector field solutions are integrable, with holonomic integral sections; in local coordinates, such a section may be expressed as
$\Psi(x)=\left(x^{\mu},y^{a}(x),\frac{\partial y^{a}}{\partial x^{\mu}}\biggr|_{x},s^{\mu}(x)\right)$
with which the coordinate-free equations (93) become
$\begin{split}\frac{\partial}{\partial x^{\mu}}\left(\frac{\partial L}{\partial y^{a}_{\mu}}\circ\Psi\right)=\left(\frac{\partial L}{\partial y^{a}}+\frac{\partial L}{\partial y^{a}_{\mu}}\frac{\partial L}{\partial s^{\mu}}\right)\circ\Psi\\ \frac{\partial s^{\mu}}{\partial x^{\mu}}=L\circ\Psi\end{split}$ (95)
We refer to these expressions as the Herglotz-Lagrange field equations [93]. When $(\mathcal{S},\Theta_{L},\Xi)$ is pre-multicontact, if multivector field solutions of (94) do exist, which is not guaranteed, they are generally not integrable. Consequently, the goal of the constraint algorithm, is to deduce the maximal submanifold $\mathcal{S}_{f}\hookrightarrow\mathcal{S}$ upon which holonomic multivector field solutions exist, and crucially, are tangent to $\mathcal{S}_{f}$.
X.2 The Pre-multicontact Constraint Algorithm
The geometrical constraint algorithm for singular field theories may be formulated in an intrinsic manner [21]; however, in what follows, we shall provide a somewhat less formal exposition, favouring a local, more practical approach. Recall that the dynamical problem to be solved is the following: given an almost-regular Lagrangian system $(\mathcal{S},\Theta_{L},\Xi)$, we seek the maximal submanifold $S_{f}\hookrightarrow\mathcal{S}$, upon which there exist locally decomposable, $\beta$-transverse, holonomic multivector field solutions to (94).
We begin by introducing the space of $d$-multivector fields which are solutions to the equations of motion, but are not necessarily $\beta$-transverse or integrable
$\textrm{ker}^{d}\,(\Theta_{L},\Omega_{L}):=\{\boldsymbol{X}_{L}\in\mathfrak{X}^{d}(\mathcal{S})\;|\;\iota_{\boldsymbol{X}_{L}}\Theta_{L}=0\quad\textrm{and}\quad\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0\,\}$
Within this space, we then seek the subset of those multivector fields which are locally decomposable and $\beta$-transverse, denoting the resulting subspace $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})\subset\textrm{ker}^{d}\,(\Theta_{L},\Omega_{L})$. In practice, as a first step, we shall often simply assume that a solution $\boldsymbol{X}_{L}$ has the local coordinate decomposition
$\boldsymbol{X}_{L}=\bigwedge\limits_{\mu=0}^{d-1}\left(\frac{\partial}{\partial x^{\mu}}+F^{a}_{\mu}\frac{\partial}{\partial y^{a}}+G^{a}_{\mu\nu}\frac{\partial}{\partial y^{a}_{\nu}}+K_{\mu}^{\nu}\frac{\partial}{\partial s^{\nu}}\right)$ (96)
which is clearly an element of $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$. Upon substituting this into the dynamical equations $\iota_{\boldsymbol{X}_{L}}\Theta_{L}=0$ and $\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$, we must ensure that the resulting expressions do not contain any inconsistencies. Supposing that we do in fact obtain a compatible set of equations, we now restrict $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$ to the subset of semi-holonomic solutions, which amounts to setting $F_{\mu}^{a}=y^{a}_{\mu}$ in the decomposition (96); in general, this step gives rise to further consistency conditions, and possibly to new constraints.
As was the case for the pre-symplectic algorithm, we must, at every stage, ensure that the dynamics of the system remain confined to the constraint submanifold; thus, for each local constraint function $\Phi$, if $\boldsymbol{X}_{L}$ is expressed as the product $\boldsymbol{X}_{L}=X_{0}\wedge\cdots\wedge X_{d-1}$, we must demand that $\mathfrak{L}_{X_{\mu}}\Phi=0$ for $\mu=0,\,\cdots,d-1$. If these tangency conditions themselves give rise to additional constraints, such functions must again have vanishing Lie derivative along each of the $X_{\mu}$.
The final step of the algorithm requires us to examine the integrability of our semi-holonomic solutions. Such an analysis is conducted by considering the constraints which arise as a result of imposing that $X_{\mu},X_{\nu}]=0$ for $\mu,\nu=0,\,\cdots,d-1$. This leads to the final subspace $\textrm{ker}^{d}_{\scriptscriptstyle H}\,(\Theta_{L},\Omega_{L})$, consisting of locally decomposable, $\beta$-transverse, holonomic multivector fields, which, in general, will only exist on the submanifold $S_{f}\hookrightarrow\mathcal{S}$.
XI Contact Reduction of Singular Field Theories
We shall now present the field-theoretic construction, which generalises the concepts developed in section (VII); as stated previously, we shall favour the Lagrangian formalism, as opposed to the Hamiltonian framework, as the latter suffers from significantly more pathological features, when considering singular field theories. Our exposition follows closely that of [29], and so we shall focus on features of the reduction process that are unique to singular theories.
Referring to the notation developed throughout section (IX.1), we suppose that $(J^{1}E,\Omega_{L})$ is a pre-multisymplectic Lagrangian system, with corresponding Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$. From the Cartan $d$-form $\Theta_{L}$, we define
$\theta_{L}^{\mu}:=-\,\iota_{\partial_{d-1}}\cdots\;\iota_{\partial_{0}}\left(\Theta_{L}\wedge dx^{\mu}\right)=\frac{\partial L}{\partial y^{a}_{\mu}}dy^{a}$ (97)
The vector field $\Sigma\in\mathfrak{X}^{\infty}(J^{1}E)$ then constitutes a scaling symmetry of degree $\Lambda$ if
$\mathfrak{L}_{\Sigma}L=\Lambda L\quad\quad\quad\textrm{and}\quad\quad\quad\mathfrak{L}_{\Sigma}\theta_{L}^{\mu}=\theta_{L}^{\mu}\quad\textrm{for all}\;\mu=0,\,\cdots,d-1$
Consider a Herglotz (i.e multicontact) Lagrangian $L^{H}$, embedded within a multisymplectic manifold of one dimension higher through the expression
$L=e^{\rho}(L^{H}+\rho_{\mu}s^{\mu})$ (98)
We shall suppose that $L^{H}$ depends upon the scalar fields $\phi^{a}$ ($a=1,\,\cdots,k$), and their first derivatives $\phi^{a}_{\mu}$, together with the action density $s^{\mu}$. We define a field variable $\rho$ in such a way that
$\rho_{\mu}=-\,\frac{\partial L^{H}}{\partial s^{\mu}}$ (99)
It is relatively straightforward to show that the equations of motion of $L$ directly imply the Herglotz-Lagrange field equations for $L^{H}$, when the former is restricted to the subspace upon which $L^{H}$ is defined. In light of this, suppose that, having identified a scaling symmetry $\Sigma$ of our pre-multisymplectic system, we adopt coordinates on $J^{1}E$ in such a way so as to render this vector field of the form
$\Sigma=\xi\frac{\partial}{\partial\xi}+\xi_{\mu}\frac{\partial}{\partial\xi_{\mu}}$ (100)
In these coordinates, the Lagrangian function depends upon both $\xi$ and $\xi_{\mu}$, together with a set of unscaled field variables $\psi^{a}$, and their corresponding velocities $\psi^{a}_{\mu}$. Finally, we make the redefinition $\xi=e^{\rho/\Lambda}$, so that the scaling symmetry vector field is simply $\Sigma=\Lambda\hskip 0.56905pt\partial_{\rho}$, and the Lagrangian takes the form
$L=e^{\rho}f(\rho_{\mu},\psi^{a},\psi^{a}_{\mu})$ (101)
The Euler-Lagrange field equation for $\rho$ then implies that the function $f$ may be written as
$f=\rho_{\mu}\frac{\partial f}{\partial\rho_{\mu}}+\frac{\partial}{\partial x^{\mu}}\frac{\partial f}{\partial\rho_{\mu}}$
which, when compared to (98), and recalling that the Herglotz-Lagrange equations require that $\partial_{\mu}s^{\mu}=L^{H}$, suggests we should identify
$s^{\mu}=\frac{\partial f}{\partial\rho_{\mu}}\quad\quad\quad\quad L^{H}=f-\rho_{\mu}s^{\mu}$ (102)
All steps outlined above are applicable to regular Lagrangian systems; however, the novelty that arises in the singular case, is that the Lagrangian $d$-form $\Theta_{L^{H}}$ calculated from the symmetry-reduced function $L^{H}$ is not guaranteed to be pre-multicontact. Consequently, as a crucial additional step, we must compute the characteristic and Reeb distributions, after making the ‘naïve’ contact reduction; if these objects do not satisfy the conditions given in section (X), the symmetry reduction procedure fails. Of course, we shall only consider those physically interesting cases, in which the reduced space does inherit a pre-multicontact structure; nevertheless, it is a non-trivial condition that must be verified.
Having calculated $\Theta_{L^{H}}$, and confirmed that $\textrm{rank}\,\mathcal{D}^{R}=d+k$ and $\textrm{rank}\,\mathcal{C}=k$, for some $0<k\leqslant n(1+d)$, we have a pre-multicontact manifold $(\mathcal{S}_{0},\Theta_{L^{H}})$, in which $\mathcal{S}_{0}$ denotes the reduced space, obtained as a result of excising the scaling variable $\rho$. Geometrically speaking, if the system of interest is such that we may globally choose a single covariant configuration space variable as a scaling function: that is to say, if the quantity $\xi$ in (100) corresponds to a single field variable on $E$, then, upon writing the Lagrangian in the form (101), $E$ separates into two connected pieces $E_{\pm}$, where each of $E_{\pm}$ is the product of a trivial bundle, and a codimension-1 subspace: $E_{\pm}\cong\widetilde{E}\times_{M}(M\times\mathbb{R}_{\pm})$. The two components $E_{\pm}$ arise as a result of the change of variable $\xi\rightarrow e^{\rho/\Lambda}$, which selects only the positive component of $\xi$. In order to cover the full range of $\xi$, we must also consider a change of variables in which $\xi\rightarrow-\,e^{\rho/\Lambda}$. Within each of $E_{\pm}$, it is precisely the trivial bundle $M\times\mathbb{R}_{\pm}\rightarrow M$ that is eliminated by the contact reduction. Thus, provided we consider both components, so as not to lose any dynamical information, the reduced space is isomorphic to $J^{1}\widetilde{E}\times\mathbb{R}^{d}$, in which the codimension-1 subspace $\widetilde{E}$ may intuitively be identified with a covariant configuration space comprised of all original field variables, except $\rho$. The reduced space does not simply inherit the structure of the jet bundle $J^{1}\widetilde{E}$, since the field variable $\rho$ has been excised, but the velocities $\rho_{\mu}$ have not; indeed these $d$ independent parameters span a copy of $\mathbb{R}^{d}$, and the Cartan form $\Theta_{L^{H}}$ determines whether this defines a Reeb distribution.
Supposing that $(\mathcal{S}_{0},\Theta_{L^{H}})$ does constitute a pre-multicontact system, with characteristic distribution $\mathcal{C}$, we calculate the $(d+1)$-form $\Omega_{L^{H}}:=\bar{d}\Theta_{L^{H}}$, and introduce a multivector field $\boldsymbol{X}_{L^{H}}\in\mathfrak{X}^{d}(\mathcal{S}_{0})$, with local decomposition
$\boldsymbol{X}_{L^{H}}=\bigwedge\limits_{\mu=0}^{d-1}\left(\frac{\partial}{\partial x^{\mu}}+F^{a}_{\mu}\frac{\partial}{\partial\psi^{a}}+G^{a}_{\mu\nu}\frac{\partial}{\partial\psi^{a}_{\nu}}+K^{\nu}_{\mu}\frac{\partial}{\partial s^{\nu}}\right)$
in which we employ notation consistent with the decomposition (101), denoting the unscaled fields $\psi^{a}$. Following this, we compute the constraints that arise as a result of setting $\iota_{\boldsymbol{X}_{L^{H}}}\Theta_{L^{H}}=0$ and $\iota_{\boldsymbol{X}_{L^{H}}}\Omega_{L^{H}}=0$, following the algorithmic procedure detailed in section (X.2). The final constraint submanifold $\mathcal{S}_{f}\hookrightarrow\mathcal{S}_{0}$ is the maximal space upon which holonomic multivector field solutions exist, and further, on this space, the observable dynamics described by these solutions coincides with that which we would have obtained by constraining the original pre-multisymplectic system.
XII Effective Non-Abelian Gauge Theory
The appearance of scalar fields that directly couple to gauge curvature terms is a phenomenon which arises most notably in string-inspired models [94, 95]. Indeed, the imprint of dilaton fields on low-energy effective actions has speculatively been regarded as a means to offer novel insight into problems such as colour confinement [96, 97]. In this section, we present an example of a non-Abelian gauge theory coupled to a dilaton-like scalar field; our goal is to illustrate how the concepts of previous sections are applied in practice, as well as to discuss the physical interpretation of our results.
XII.1 Geometrical Setting
Let $\pi:P\rightarrow M$ be a principle $G$-bundle over the $d$-dimensional spacetime manifold $M$; on $P$, there exists a unique, Lie algebra valued connection $1$-form $\omega\in\Omega^{1}(P,\mathcal{L}(G))$, which descends to a local form $A:=s^{*}\omega\in\Omega^{1}(U,\mathcal{L}(G))$, in which $s:U\rightarrow P$ is a local trivialising section of $P$, defined on the open subset $U\subset M$. We shall suppose that the spacetime manifold $M$ is equipped with a Lorentzian metric $g$ of signature $(+,-,-,\,\cdots\,)$; this metric is considered to be parametric, and non-variational [19], so that, physically, our theory is defined on a curved background, which we may change, but is not coupled to gravity. The appropriate covariant configuration space for this theory is
$\mathcal{E}=C(P)\times_{M}\left(M\times\mathbb{R}\right)\times_{M}\textrm{Sym}^{1,d-1}_{2}(M)$ (103)
where $C(P)\rightarrow M$ is the bundle of connections on $P$, with $C(P)\cong J^{1}P/G$ [98]. The space $\textrm{Sym}^{1,d-1}_{2}(M)$ refers to the set of symmetric covariant tensors of rank 2 and Lorentzian signature (1,d1)(1,d-1) on $M$, and thus parameterises our choice of metric. Local coordinates on $E$ are denoted $(x^{\mu},A^{a}_{\mu},\hskip 1.13809pt\phi\hskip 1.13809pt;g_{\mu\nu})$, in which the semicolon separates the parametric and variational degrees of freedom, and the index $a$ runs from $a=1,\,\cdots,\textrm{dim}\,\mathcal{L}(G):=n$. The Lagrangian density is a semibasic $d$-form on the space
$\mathcal{Q}:=J^{1}\left(C(P)\times_{M}\left(M\times\mathbb{R}\right)\right)\times_{M}\textrm{Sym}^{1,d-1}_{2}(M)$ (104)
upon which we take local coordinates to be $(x^{\mu},A^{a}_{\mu},\hskip 1.13809pt\phi,A^{a}_{\mu,\nu},\hskip 1.13809pt\phi_{\mu}\hskip 1.13809pt;g_{\mu\nu})$. The corresponding Lagrangian function $L:\mathcal{Q}\rightarrow\mathbb{R}$ is
$L=\textrm{Tr}\left[-\frac{\phi^{2}}{2g^{2}}F_{\mu\nu}F^{\mu\nu}+2J_{\mu}A^{\mu}\right]+\frac{1}{2}g^{\mu\nu}\phi_{\mu}\phi_{\nu}-V(\phi)$ (105)
where the trace is taken over the indices of $\mathcal{L}(G)$, and $F$ refers to the Lie algebra valued curvature 2-form $F:=dA+\frac{1}{2}[A,A]$.
In what follows, we shall adopt group-theoretic conventions more prevalent in the physics literature, taking the generators $T_{a}$ of the Lie algebra to be Hermitian, with $T_{a},T_{b}]=if_{ab}^{\;\;\;c}\,T_{c}$. Supposing that $G$ is a semi-simple compact Lie group, the Hermiticity of the $T_{a}$ implies that the Killing form is both non-degenerate, and positive semi-definite; in the fundamental representation, we adopt the following normalisation with respect to the trace
$\textrm{Tr}\left(T_{a}T_{b}\right)=\frac{1}{2}\delta_{ab}$ (106)
and we raise and lower indices with $\delta^{ab}$ and $\delta_{ab}$, writing $f_{abc}=\delta_{cd}\,f_{ab}^{\;\;\;d}$, for example. Returning to the Lagrangian (105), $J_{\mu}^{a}$ is a Lie algebra valued 1-form that couples to the gauge field, acting as a source term, and $V(\phi)$ is the non-perturbative dilaton potential. For simplicity of exposition, we shall take $V(\phi)$ to be a single mass term $\frac{1}{2}m^{2}\phi^{2}$, and temporarily ignore the source; having carried out the contact reduction, we will briefly comment on the implications of reinstating $J_{\mu}^{a}$.
We remark that, with the above conventions, we have implicitly assumed that the dilaton transforms trivially under $G$; were $\phi$ to transform in some non-trivial representation $\rho:G\rightarrow GL(V)$, we would instead have to consider the associated bundle $P\times_{\rho}V$, with the result that our covariant configuration space $\mathcal{E}$ would instead be
$\mathcal{E}=C(P)\times_{M}\left(P\times_{\rho}V\right)\times_{M}\textrm{Sym}_{2}^{1,d-1}(M)$
where the trivial bundle $M\times\mathbb{R}$ has been replaced with the vector bundle $P\times_{\rho}V$ associated to $P$ via the representation $\rho$. In the interest of limiting the number of additional complications, we shall content ourselves with a dilaton that transforms trivially. With these considerations, we rewrite (105) as
$L=-\,\frac{\phi^{2}}{4g^{2}}F^{a}_{\mu\nu}F_{a}^{\mu\nu}+\frac{1}{2}g^{\mu\nu}\phi_{\mu}\phi_{\nu}-\frac{1}{2}m^{2}\phi^{2}$ (107)
in which the curvature $F^{a}_{\mu\nu}$ has the following local coordinate expression
$F_{\mu\nu}^{a}=A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f_{bc}^{\;\;\;a}A^{b}_{\mu}A^{c}_{\nu}$
The degeneracy of this Lagrangian arises as a result of the gauge symmetry; indeed, the Hessian matrix with respect to $A^{a}_{\mu,\nu}$ and $\phi_{\mu}$ has the following entries
$\frac{\partial^{2}L}{\partial A^{a}_{\nu,\mu}\,\partial A^{b}_{\lambda,\kappa}}=-\,\frac{\phi^{2}}{g^{2}}\delta_{ab}\left(g^{\nu\lambda}g^{\mu\kappa}-g^{\nu\kappa}g^{\mu\lambda}\right)\quad\quad\frac{\partial^{2}L}{\partial\phi_{\mu}\,\partial\phi_{\nu}}=g^{\mu\nu}\quad\quad\frac{\partial^{2}L}{\partial A^{a}_{\nu,\mu}\,\partial\phi_{\lambda}}=0$
which, from the antisymmetric combinations of $g^{\mu\nu}$, is non-invertible. The multisymplectic form $\Theta_{L}$ is of degree $d$, and from (82), we find that
$\begin{split}\Theta_{L}=\left(-\,\frac{\phi^{2}}{g^{2}}F^{\mu\nu}_{a}\,dA^{a}_{\nu}+g^{\mu\nu}\phi_{\nu}\,d\phi\right)\wedge\,d^{d-1}x_{\mu}-\left(-\,\frac{\phi^{2}}{g^{2}}F^{\mu\nu}_{a}A^{a}_{\nu,\mu}\right.\\ \left.+\frac{\phi^{2}}{4g^{2}}F^{a}_{\mu\nu}F_{a}^{\mu\nu}+\frac{1}{2}g^{\mu\nu}\phi_{\mu}\phi_{\nu}+\frac{1}{2}m^{2}\phi^{2}\right)d^{d}x\end{split}$ (108)
As described in section (XI), from $\Theta_{L}$, we extract the following 1-forms
$\theta^{\mu}_{L}:=-\,\iota_{\partial_{d-1}}\,\cdots\;\,\iota_{\partial_{0}}\left(\Theta_{L}\wedge dx^{\mu}\right)=-\,\frac{\phi^{2}}{g^{2}}F^{\mu\nu}_{a}\,dA^{a}_{\nu}+g^{\mu\nu}\phi_{\nu}\,d\phi$ (109)
From the structure of $\theta^{\mu}_{L}$, and the Lagrangian (107), it is relatively clear that the vector field
$\Sigma=\frac{1}{2}\left(\phi\frac{\partial}{\partial\phi}+\phi_{\mu}\frac{\partial}{\partial\phi_{\mu}}\right)$
satisfies $\mathfrak{L}_{\Sigma}\theta_{L}^{\mu}=\theta^{\mu}_{L}$ and $\mathfrak{L}_{\Sigma}L=L$, with which we conclude that $\Sigma$ is a scaling symmetry of degree one.
XII.2 Contact Reduction
Upon making the change of variables $\phi\mapsto e^{\rho/2}$, the scaling symmetry vector field $\Sigma$ is now simply $\partial_{\rho}$, and the Lagrangian, with the curvature term expanded in full, reads
$\begin{split}L=-\,\frac{e^{\rho}}{4g^{2}}\biggr[g^{\mu\lambda}g^{\nu\kappa}\left(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\right)\left(A^{a}_{\kappa,\lambda}-A^{a}_{\lambda,\kappa}+f_{de}^{\;\;\;a}A^{d}_{\lambda}A^{e}_{\kappa}\right)\biggr]\\ +e^{\rho}\left(\frac{1}{8}g^{\mu\nu}\rho_{\mu}\rho_{\nu}-\frac{1}{2}m^{2}\right)\end{split}$ (110)
Recall from section (XI) that when the Lagrangian is expressed in the form $L=e^{\rho}f(\rho_{\mu},A_{\mu},A_{\mu,\nu})$, the action density is found from $s^{\mu}=\partial f/\partial\rho_{\mu}$, whilst the Herglotz Lagrangian takes the form $L^{H}=f-\rho_{\mu}s^{\mu}$; for the former, we find
$s^{\mu}:=\frac{\partial f}{\partial\rho_{\mu}}=\frac{1}{4}g^{\mu\nu}\rho_{\nu}\quad\quad\implies\quad\quad\rho_{\mu}=4g_{\mu\nu}s^{\nu}$ (111)
In anticipation of the elimination of $\rho_{\mu}$ in favour of the action density, we have expressed the former in terms of the latter. The Herglotz Lagrangian for this system is then
$L^{H}=-\,\frac{1}{4g^{2}}\biggr[g^{\mu\lambda}g^{\nu\kappa}\left(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\right)\left(A^{a}_{\kappa,\lambda}-A^{a}_{\lambda,\kappa}+f^{\;\;\;a}_{de}A^{d}_{\lambda}A^{e}_{\kappa}\right)\biggr]-\frac{1}{2}m^{2}-2g_{\mu\nu}s^{\mu}s^{\nu}$ (112)
Due to the singular nature of the Lagrangian (107) from which $L^{H}$ is constructed, it is not a priori guaranteed that the $d$-form $\Theta_{L^{H}}$ endows the reduced space with a pre-multicontact structure. Consequently, prior to implementing the constraint algorithm, we must compute the characteristic and Reeb distributions, verifying that they do indeed satisfy the conditions required for $\Theta_{L^{H}}$ to be pre-multicontact.
Geometrically, referring to the decomposition (103), we see that the excision of the dilaton corresponds to the removal of the trivial bundle $M\times\mathbb{R}\rightarrow M$, defining a reduced covariant configuration space $\mathcal{E}_{\textrm{red}}$. More formally, since $\phi$ is a globally-defined scaling function, the symplectification $\widetilde{\mathcal{E}}_{\textrm{red}}$ is a trivial $\mathbb{R}^{\times}$-bundle over $M$, composed of two connected components $\widetilde{\mathcal{E}}_{\textrm{red}}^{\pm}$, which, with the change of variables $\phi\rightarrow e^{\rho/2}$, correspond to symplectification via $\pm e^{\rho/2}$. Consequently, with the understanding that care must be taken to ensure that both components are considered, we may somewhat informally speak of a reduced space $\mathcal{E}_{\textrm{red}}$ that is isomorphic to $C(P)\times_{M}\textrm{Sym}^{1,d-1}_{2}(M)$. The Herglotz Lagrangian is defined on the space
$\mathcal{Q}_{\textrm{red}}=\left(J^{1}C(P)\times_{M}\textrm{Sym}_{2}^{1,d-1}(M)\right)\times\mathbb{R}^{d}$ (113)
where the additional factor of $\mathbb{R}^{d}$ comes from the $d$ velocities $\rho_{\mu}$, which have assumed the role of the action density $s^{\mu}$. The task is then to deduce whether the $d$-form $\Theta_{L^{H}}$, given by
$\Theta_{L^{H}}=\left(ds^{\mu}+\frac{1}{g^{2}}F^{\mu\nu}_{a}\,dA^{a}_{\nu}\right)\wedge d^{d-1}x_{\mu}+\left(-\,\frac{1}{g^{2}}F^{\mu\nu}_{a}A_{\nu,\mu}^{a}+\frac{1}{4g^{2}}F^{\mu\nu}_{a}F^{a}_{\mu\nu}+\frac{1}{2}m^{2}+2g_{\mu\nu}s^{\mu}s^{\nu}\right)d^{d}x$ (114)
endows $\mathcal{Q}_{\textrm{red}}$ with a pre-multicontact structure. In order to calculate $\Omega_{L^{H}}$, we require the dissipation form $\sigma$
$\sigma=-\,\frac{\partial L^{H}}{\partial s^{\mu}}\,dx^{\mu}=4g_{\mu\nu}s^{\nu}\hskip 0.56905ptdx^{\mu}$
Thus we find that the $(d+1)$-form $\Omega_{L^{H}}$ is given by
$\begin{split}\Omega_{L^{H}}=\frac{1}{g^{2}}dF^{\mu\nu}_{a}\wedge\,dA^{a}_{\nu}\wedge d^{d-1}x_{\mu}+\left(\frac{1}{2g^{2}}\left(F_{a}^{\mu\nu}dF_{\mu\nu}^{a}-2F_{a}^{\mu\nu}\,dA^{a}_{\nu,\mu}-2A^{a}_{\nu,\mu}\,dF^{\mu\nu}_{a}\right)\right.\\ \left.-\frac{4}{g^{2}}g_{\mu\lambda}s^{\lambda}F^{\mu\nu}_{a}\,dA^{a}_{\nu}\right)\wedge d^{d}x\end{split}$ (115)
From the local expression (114), we see that $\iota_{\partial_{s^{\mu}}}\Theta_{L^{H}}=d^{d-1}x_{\mu}$, and that the velocity coordinates $A^{a}_{\nu,\mu}$ only appear in the antisymmetric combinations $A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}$. Additionally, it is straightforward to calculate that $\iota_{\partial_{s^{\mu}}}\,d\Theta_{L^{H}}=4g_{\mu\nu}s^{\nu}\,d^{d}x$, which is a $d$-form that is annihilated by any vector field in $\textrm{ker}\,\Xi$. From these observations, we deduce that the characteristic and Reeb distributions are given by
$\mathcal{C}=\;\biggr\langle\frac{\partial}{\partial A^{a}_{\nu,\mu}}+\frac{\partial}{\partial A_{\mu,\nu}^{a}}\biggr\rangle\quad\quad\quad\quad\mathcal{D}^{R}=\;\biggr\langle\frac{\partial}{\partial A^{a}_{\nu,\mu}}+\frac{\partial}{\partial A_{\mu,\nu}^{a}},\frac{\partial}{\partial s^{\mu}}\biggr\rangle$ (116)
for $\mu,\nu=0,\,\cdots,d-1$ and $a=1,\,\cdots,\textrm{dim}\,\mathcal{L}(G)=n$. Note that the ranks of these distributions are
$\textrm{rank}\,\mathcal{C}=n\,\frac{d(d+1)}{2}\quad\quad\quad\textrm{rank}\,\mathcal{D}^{R}=n\,\frac{d(d+1)}{2}+d$
which, referring to the discussion of section (X.1), are consistent with a pre-multicontact distribution with $k=d$. Consequently, we conclude that $\Theta_{L^{H}}$ does in fact define a pre-multicontact structure on $\mathcal{Q}_{\textrm{red}}$, and so may proceed with the constraint analysis; for the purpose of discussing the construction of multivector field solutions, we introduce the projection $\beta:\mathcal{Q}_{\textrm{red}}\rightarrow M$.
XII.3 The Constraint Algorithm
As described in section (X.2), the objective of the constraint procedure is to deduce the maximal submanifold $\mathcal{Q}_{f}\hookrightarrow\mathcal{Q}_{\textrm{red}}$ upon which there exist multivector field solutions $\boldsymbol{X}$ of the equations $\iota_{\boldsymbol{X}}\Theta_{L^{H}}=0$ and $\iota_{\boldsymbol{X}}\Omega_{L^{H}}=0$, that are locally decomposable, $\beta$-transverse, and holonomic. We beging by supposing that a locally decomposable, $\beta$-transverse solution $\boldsymbol{X}\in\textrm{ker}^{d}_{\beta}\,(\Theta_{L^{H}},\Omega_{L^{H}})$ exists, with
$\boldsymbol{X}=\bigwedge\limits_{\mu=0}^{d-1}\left(\frac{\partial}{\partial x^{\mu}}+C^{a}_{\mu\nu}\frac{\partial}{\partial A^{a}_{\nu}}+G^{a}_{\mu\nu\kappa}\frac{\partial}{\partial A^{a}_{\kappa,\nu}}+K_{\mu}^{\nu}\frac{\partial}{\partial s^{\nu}}\right):=\bigwedge\limits_{\mu=0}^{d-1}V_{\mu}$ (117)
The contraction of $\boldsymbol{X}$ with the pre-multicontact form $\Theta_{L^{H}}$ provides a single equation, as the degree of the multivector field coincides with that of the differential form. Imposing that $\iota_{\boldsymbol{X}}\Theta_{L^{H}}=0$, we find that
$K^{\mu}_{\mu}=\frac{1}{g^{2}}\left(A^{a}_{\nu,\mu}-C^{a}_{\mu\nu}\right)F^{\mu\nu}_{a}+\left(-\,\frac{1}{4g^{2}}F^{\mu\nu}_{a}F^{a}_{\mu\nu}-\frac{1}{2}m^{2}-2g_{\mu\nu}s^{\mu}s^{\nu}\right)$ (118)
Note that the final quantity on the right-hand side is precisely $L^{H}$, and the bracket $\left(A^{a}_{\nu,\mu}-C^{a}_{\mu\nu}\right)$ vanishes upon imposing semi-holonomy.
The contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$ is somewhat involved, and the lengthy algebraic manipulations required are not particularly illuminating; for completeness, we provide the details in appendix A. Performing the contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$, and setting the result to zero, we obtain the following pair of expressions
$0$ $=(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu})-(C^{a}_{\mu\nu}-C^{a}_{\nu\mu})$ (119)
$0$ $=g^{\mu\rho}g^{\nu\sigma}\left(G_{\mu\rho\sigma}^{a}-G^{a}_{\mu\sigma\rho}\right)+f_{bc}^{\;\;\;a}g^{\nu\rho}A^{b\mu}\left(C^{c}_{\mu\rho}-C^{c}_{\rho\mu}\right)+f_{bc}^{\;\;\;a}\left(g^{\nu\rho}A^{b\mu}-g^{\mu\rho}A^{b\nu}\right)C^{c}_{\mu\rho}$ (120)
$\quad+f_{bc}^{\;\;\;e}f_{de}^{\;\;\;a}A^{b\mu}A^{c\nu}A^{d}_{\mu}+4g_{\mu\lambda}s^{\lambda}F^{a\mu\nu}$
Having computed $\iota_{\boldsymbol{X}}\Theta_{L^{H}}$ and $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$ in local coordinates, the next stage of the algorithm requires that we examine the effects of imposing semi-holonomy; [^6] Recall that it is semi-holonomy that we must impose first, since we have not yet considered the integrability of $\boldsymbol{X}$. for the multivector field (117), this implies we should set $C^{a}_{\mu\nu}=A^{a}_{\nu,\mu}$. With this, (119) is rendered trivial, and (118) reduces to $K^{\mu}_{\mu}=L^{H}$, as expected. Following the procedure detailed in appendix A, combining the $G^{a}_{\mu\rho\sigma}$ terms with $f_{bc}^{\;\;\;a}\left(g^{\nu\rho}A^{b\mu}-g^{\mu\rho}A^{b\nu}\right)C^{c}_{\mu\rho}$, and denoting the result $g^{\mu\rho}g^{\nu\sigma}F^{a}_{\rho\sigma,\mu}$, we find that (120) becomes
$0=g^{\mu\rho}g^{\nu\sigma}F^{a}_{\rho\sigma,\mu}+f_{bc}^{\;\;\;a}g^{\nu\rho}A^{b\mu}F^{c}_{\mu\rho}+4g_{\mu\lambda}s^{\lambda}F^{a\mu\nu}$ (121)
It should be noted that, in the above, care has been taken to ensure that our notation for the velocity coordinates of the gauge field is merely reminiscent of a partial derivative with respect to the spacetime coordinates. In particular, until we have examined the integrability of our system, $A^{a}_{\nu,\mu}\neq\partial_{\mu}A^{a}_{\nu}$ and $G^{a}_{\mu\nu\sigma}\neq\partial_{\mu}\partial_{\nu}A^{a}_{\sigma}$. However, upon imposing the necessary conditions for $\textrm{span}\,(V_{\mu})$ to define an involutive distribution, we may affirm that the multivector field $\boldsymbol{X}$ possesses holonomic integral sections $\psi:M\rightarrow\mathcal{Q}_{\textrm{red}}$, whose local coordinate expression is
$\psi(x)=\left(x^{\mu},A^{a}_{\mu}(x),\frac{\partial A^{a}_{\mu}}{\partial x^{\nu}}(x),s^{\mu}(x)\right)$ (122)
Only after introducing these integral sections may we identify $A^{a}_{\nu,\mu}$ with the familiar $\partial_{\mu}A^{a}_{\nu}$. In light of this, we now turn to an analysis of the integrability of $\boldsymbol{X}$; the distribution formed by the $V_{\mu}$ will be involutive if, and only if it is closed under the Lie bracket; however, since partial derivatives with respect to the spacetime coordinates commute, involutivity is only guaranteed if $V_{\mu},V_{\nu}]=0$ for all $\mu,\nu=0,\,\cdots,d-1$. Thus, computing the Lie bracket, and demanding that the result be zero, we find that
$G^{a}_{\nu\mu\sigma}-G^{a}_{\mu\nu\sigma}=A^{a}_{\sigma,\mu\nu}-A^{a}_{\sigma,\nu\mu}$ (123)
This condition produces no inconsistencies for the coefficients of the multivector field, and so (118) and (121) may be reexpressed as
$\begin{split}\partial_{\mu}s^{\mu}=L^{H}\\ \partial_{\mu}F^{a\mu\nu}+f_{bc}^{\;\;\;a}A^{b}_{\mu}F^{c\mu\nu}+4g_{\mu\lambda}s^{\lambda}F^{a\mu\nu}=0\end{split}$ (124)
The first two terms of the second expression form the familiar covariant derivative $(\mathcal{D}_{\mu}F^{\mu\nu})^{a}$, whilst the final term is a common feature of contact-reduced theories, that arises as a result of eliminating a scaling degree of freedom from our ontology. More concretely, in the original Lagrangian (105), there was a coupling between the scalar field $\phi$ and the gauge-kinetic term $\textrm{Tr}\,(F_{\mu\nu}F^{\mu\nu})$; the removal of the dilaton, in favour of a minimal ontology, comes at the cost of introducing a dissipative characteristic into the theory, which is made manifest through the coupling of the gauge field strength to the action density.
Before concluding, there remains an important matter to be discussed, concerning the physical interpretation of the excision of the scalar field $\phi$. In particular, one may be inclined to object that, for theories in which parameter values are determined by expectation values of scalar fields, it is not possible to claim that such fields are ‘redundant’. However, this is not the case; consider, for instance, an experiment in which one attempts to deduce the expectation value $\langle\phi\rangle$, which, for the current discussion, we shall suppose to fix some coupling parameter $g_{s}$. The value of $g_{s}$ is not something directly accessible to an observer, but rather is deduced via comparison to a reference object; this is entirely analogous to how an object of length 1 m is only known to be such because its size is in a 1:1 ratio with that of a previously-standardised instrument. Consequently, under changes generated by $\phi$, the measuring apparatus is subjected to precisely the same rescaling as $g_{s}$, with the result that the overall scaling is empirically inaccessible. Thus, we are justified in asserting that this degree of freedom is redundant, from the perspective of the algebra of dynamical observables.
In fact, we make a somewhat bolder claim: if one is able to construct a theory whose empirically-accessible degrees of freedom are invariant under a particular transformation, we assert that it is a necessary condition, for any relational description, that the mathematical artifice generating such transformations be excisable.
To appreciate why this must be so, consider an arbitrary system $\mathcal{S}$, together with an intrinsic observer, who wishes to construct a relational description of $\mathcal{S}$. The observer has at their disposal only those quantities whose values are ascertainable through physical measurement. For instance, all lengths are deduced with respect to a standard metre stick, and similarly, all times are measured as multiples of the standardised ‘tick’ of a clock.
Suppose that we ourselves are extrinsic ($i.e$ external) observers to $\mathcal{S}$, and we note that, under a certain transformation $\mathcal{T}$, the set of variables we use to describe $\mathcal{S}$ suffer a rescaling. However, this rescaling is such that the dynamical equations we deduce remain invariant. As a trivial example, we consider a pair of variables $(x,t)$, with $\mathcal{T}:(x,t)\mapsto(\lambda x,\lambda t)$; the velocity $\dot{x}$ remains invariant, even though the variables used to deduce $\dot{x}$ do not. The generator of the transformation $\mathcal{T}$ must be a redundant component of our description; if it were redundant, but non-excisable, then it would necessarily appear in the ontology of the observer, else the intrinsic and extrinsic theories would not describe the same system. This contradicts our assumption that the observer’s description is relational. Similarly, if the generator of $\mathcal{T}$ were non-excisable because its presence was required to deduce dynamical evolution, it would be false to claim that the relational description is invariant under $\mathcal{T}$. Consequently, in order for it to be simultaneously true that the observer’s description is wholly relational and that there exists a transformation which leaves the dynamical observables invariant, it must be the case that the generator of this transformation is a feature of our mathematical description that can be removed.
Finally, before concluding, we briefly comment that, had we retained the source term $J_{\mu}^{a}$ in our Lagrangian, after introducing the change of variables $\phi\rightarrow e^{\rho/2}$, (110) would instead read
$L=e^{\rho}\left(-\frac{1}{4g^{2}}F_{\mu\nu}^{a}F^{\mu\nu}_{a}+\frac{1}{8}g^{\mu\nu}\rho_{\mu}\rho_{\nu}-\frac{1}{2}m^{2}\right)+J^{\mu}_{a}A^{a}_{\mu}$ (125)
Clearly the presence of this additional term destroys the scaling symmetry, and we are no longer able to make a contact reduction. For theories of particles (described within the Hamiltonian formalism), an expedient trick for treating such systems is to make the coupling dynamical, by enlarging the phase space, appending a copy of $\mathbb{R}\times\mathbb{R}_{+}$ to the cotangent bundle. With this, the coupling is promoted to a momentum variable, such that the Hamiltonian is independent of the corresponding conjugate position. In this way, Hamilton’s equations reveal that the new momentum is constant in time, which is precisely the desired outcome [14]. At present, more work is required to find a field-theoretic analogue - if such an extension exists - which allows coupling parameters to acquire their own dynamics, rendering Lagrangians of the type (125) reducible.
XIII Conclusions and Outlook
Many of the most mathematically-rich models in contemporary theoretical physics possess, as a central component, gauge symmetries; such theories are described by singular Lagrangians, and over the course of this article, we have provided a basic introduction to gauge theories, from the perspective of geometric mechanics. For systems of particles, the appropriate formalism was that of pre-symplectic geometry, while a field-theoretic discussion required the introduction of the fibred manifolds of pre-multisymplectic geometry. We have demonstrated that singular theories which exhibit invariance under global scaling transformations may be subjected to a reduction process similar to that which has been studied for regular systems, provided that degeneracies are controlled, through use of a suitable constraint algorithm. We observed that this constraint procedure could be implemented pre- or post-reduction; from a physical perspective, this is to be expected: if the processes of contact reduction and phase space restriction were non-commutative, this would imply that the constraint functions had, in some sense, information about the global scaling variable. Had the reduction of a constrained system led to a theory distinct from that obtained via the constraint of a reduced theory, then any system described by a degenerate Lagrangian would offer a means to detect changes in global scale, which is a highly unphysical outcome.
Having described the reduction of singular field theories, we then provided an example of an effective non-Abelian gauge theory coupled to a dilaton-like field; it was observed that this real scalar field was in fact redundant, and could therefore be excised. The elimination of this degree of freedom was, at first sight, relatively alarming, particularly because the coupling parameters of a theory are often fixed by the expectation value of such scalar fields. However, we argued that the strength of an interaction is something that must necessarily be deduced via empirical methods; as such, any hypothetical experiment, conducted with the aim of measuring such parameters will be insensitive to a rescaling of their values, since the apparatus itself will suffer an identical change in scale.
The results presented throughout constitute a framework within which to analyse singular theories; it is known that, upon making a conformal decomposition of the spacetime metric, the Einstein-Hilbert Lagrangian possesses a scaling symmetry, corresponding precisely to the conformal factor [34]. To our knowledge, this scaling symmetry has not been studied within the first-order Palatini formalism, for which the multisymplectic Lagrangian description has been developed in [99] and [100]. From the results of the present work, we know that it is possible to implement the geometrical constraint algorithm after having eliminated the scaling degree of freedom. As such, it will be of great interest to examine any simplifications that arise, as a result of working with a reduced ontology, in which we retain only those degrees of freedom required to close the algebra of dynamical observables.
In general, we distinguish between two principal approaches to the treatment of classical field theory. On the one hand, as is customary when attempting to construct a quantum analogue of the theory, one may consider the temporal evolution of the initial data, defined on a particular choice of Cauchy slice [101]. This ‘canonical’ formulation studies spatially-dependent fields, defined at fixed instants of time, and has the notably inconvenient features of possessing an infinite-dimensional phase space, and sacrificing manifest Lorentz covariance [102]. The ADM formalism of General Relativity is a constrained Hamiltonian theory that adopts just such an approach, considering the (globally-hyperbolic) spacetime manifold to be foliated by spacelike hypersurfaces [103, 104].
The Hamilton De-Donder Weyl formalism we have employed throughout makes no distinction between the spacetime coordinates, which immediately guarantees the Lorentz covariance that was absent in the canonical formulation. Having established a framework within which to treat constrained theories, it will be highly illustrative to contrast the constraint of a contact-reduced action for General Relativity with the more familiar canonical approach.
Appendix A Multivector Field Manipulations for Non-Abelian Gauge Theory
Here, we provide details of the algebraic manipulations required to compute $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$, that were omitted from the main text. For convenience, we reproduce the local form of $\Omega_{L^{H}}$ given in (115)
$\begin{split}\Omega_{L^{H}}=\frac{1}{g^{2}}dF^{\mu\nu}_{a}\wedge\,dA^{a}_{\nu}\wedge d^{d-1}x_{\mu}+\left(\frac{1}{2g^{2}}\left(F_{a}^{\mu\nu}dF_{\mu\nu}^{a}-2F_{a}^{\mu\nu}\,dA^{a}_{\nu,\mu}-2A^{a}_{\nu,\mu}\,dF^{\mu\nu}_{a}\right)\right.\\ \left.-\frac{4}{g^{2}}g_{\mu\lambda}s^{\lambda}F^{\mu\nu}_{a}\,dA^{a}_{\nu}\right)\wedge d^{d}x\end{split}$ (126)
We begin by simplifying this expression, noting that the term $(F_{\mu\nu}^{a}-2A_{\nu,\mu}^{a})\,dF^{\mu\nu}_{a}$ may be expanded as
$\Omega_{L^{H}}\;\supset\;(F_{\mu\nu}^{a}-2A_{\nu,\mu}^{a})\,dF^{\mu\nu}_{a}=-\left(A^{a}_{\nu,\mu}+A^{a}_{\mu,\nu}-f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\right)\,dF^{\mu\nu}_{a}=f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\,dF^{\mu\nu}_{a}$
where, in the final equality, we have used the fact that $(A^{a}_{\nu,\mu}+A^{a}_{\mu,\nu})$ is symmetric under the exchange of $\mu$ and $\nu$, whereas $F^{\mu\nu}_{a}$ is antisymmetric. The differential $dF^{\mu\nu}_{a}$ may itself be expanded, allowing the second bracket of (126) to be written as
$\frac{1}{2g^{2}}\left(f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\,dF^{\mu\nu}_{a}-2\,F^{\mu\nu}_{a}\,dA^{a}_{\nu,\mu}\right)$ $=\frac{1}{2g^{2}}\left(f_{abc}A^{b\mu}A^{c\nu}\,dF_{\mu\nu}^{a}-2\,F^{\mu\nu}_{a}\,dA^{a}_{\nu,\mu}\right)$
$=\frac{1}{2g^{2}}\left(\left(f_{abc}A^{b\mu}A^{c\nu}-F_{a}^{\mu\nu}\right)dA^{a}_{\nu,\mu}-\left(f_{abc}A^{b\mu}A^{c\nu}-F_{a}^{\mu\nu}\right)dA^{a}_{\mu,\nu}\right$.
$\quad+\left.f_{abc}f_{de}^{\;\;\;a}A^{b\mu}A^{c\nu}\left(A^{d}_{\mu}\,dA^{e}_{\nu}+A^{e}_{\nu}\,dA^{d}_{\mu}\right)\right)$
$=\frac{1}{g^{2}}\left(\left(A^{\mu,\nu}_{a}-A^{\nu,\mu}_{a}\right)dA^{a}_{\nu,\mu}+f_{abc}f_{de}^{\;\;\;a}A^{b\mu}A^{c\nu}A_{\mu}^{d}\,dA^{e}_{\nu}\right)$
In passing from the second to the third line, we have expanded each of the curvature terms, so that
$f_{abc}A^{b\mu}A^{c\nu}-F_{a}^{\mu\nu}=A_{a}^{\mu,\nu}-A_{a}^{\nu,\mu}$
We have then relabeled dummy indices, and used the antisymmetry of $F_{a}^{\mu\nu}$. With this, the final form of $\Omega_{L^{H}}$ we use to deduce the geometrical field equations is
$\begin{split}\Omega_{L^{H}}=\;&\frac{1}{g^{2}}\,dF^{\mu\nu}_{a}\wedge\,dA^{a}_{\nu}\wedge d^{d-1}x_{\mu}\\ &+\frac{1}{g^{2}}\biggr[\left(A^{\mu,\nu}_{a}-A^{\nu,\mu}_{a}\right)dA^{a}_{\nu,\mu}+\left(f_{abc}f_{de}^{\;\;\;a}A^{b\mu}A^{c\nu}A_{\mu}^{d}-4g_{\mu\lambda}s^{\lambda}F^{\mu\nu}_{e}\right)dA^{e}_{\nu}\biggr]\wedge d^{d}x\end{split}$ (127)
The multivector field $\boldsymbol{X}$ is of degree $d$, while $\Omega_{L^{H}}$ is a $(d+1)$-form; as such, the contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$ yields a 1-form. The pair of expressions (119) and (120) arise as a result of contracting the $d$ fields $V_{\mu}$ with $\Omega_{L^{H}}$ in such a way so as to leave a form proportional to $dA^{a}_{\nu,\mu}$ and $dA^{a}_{\nu}$ respectively. In the latter case, particular care must be taken, for $dF^{\mu\nu}_{a}$ contains $dA^{a}_{\nu}$ terms, which require us to swap the order of the triple wedge product, relabel dummy indices, and use the antisymmetry of $f_{abc}$ to obtain the result. The equations of motion, reproduced here for convenience, read
$0$ $=(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu})-(C^{a}_{\mu\nu}-C^{a}_{\nu\mu})$ (128)
$0$ $=g^{\mu\rho}g^{\nu\sigma}\left(G_{\mu\rho\sigma}^{a}-G^{a}_{\mu\sigma\rho}\right)+f_{bc}^{\;\;\;a}g^{\nu\rho}A^{b\mu}\left(C^{c}_{\mu\rho}-C^{c}_{\rho\mu}\right)+f_{bc}^{\;\;\;a}\left(g^{\nu\rho}A^{b\mu}-g^{\mu\rho}A^{b\nu}\right)C^{c}_{\mu\rho}$ (129)
$\quad+f_{bc}^{\;\;\;e}f_{de}^{\;\;\;a}A^{b\mu}A^{c\nu}A^{d}_{\mu}+4g_{\mu\lambda}s^{\lambda}F^{a\mu\nu}$
When semi-holonomy is imposed, the first term of (120) may be combined with $f_{bc}^{\;\;\;a}\left(g^{\nu\rho}A^{b\mu}-g^{\mu\rho}A^{b\nu}\right)C^{c}_{\mu\rho}$ to produce an object we shall denote $g^{\mu\rho}g^{\nu\sigma}F^{a}_{\rho\sigma,\mu}$; this is achieved by relabeling both spacetime and Lie algebra indices, and using the antisymmetric properties of $f_{bc}^{\;\;\;a}$. Finally, we write $(C^{c}_{\mu\rho}-C^{c}_{\rho\mu})$ as $F^{c}_{\mu\rho}-f_{de}^{\;\;\;c}A^{d}_{\mu}A^{e}_{\rho}$, and combine terms, using the Jacobi identity, as follows
$f_{bc}^{\;\;\;e}f_{de}^{\;\;\;a}A^{b\mu}A^{c\nu}A^{d}_{\mu}-f_{bc}^{\;\;\;a}f_{de}^{\;\;\;c}A^{e\nu}A^{d}_{\mu}A^{b\mu}=\left(f_{de}^{\;\;\;a}f_{bc}^{\;\;\;e}+f_{be}^{\;\;\;a}f_{cd}^{\;\;\;e}\right)A^{b\mu}A^{c\nu}A^{d}_{\mu}=-\,f_{ce}^{\;\;\;a}f_{db}^{\;\;\;e}A^{b\mu}A^{c\nu}A^{d}_{\mu}$
It is then clear that this term vanishes as a result of the antisymmetry of $f_{db}^{\;\;\;e}$ in $d$ and $b$. Combining these observations, the equation of motion
$0=g^{\mu\rho}g^{\nu\sigma}F^{a}_{\rho\sigma,\mu}+f_{bc}^{\;\;\;a}g^{\nu\rho}A^{b\mu}F^{c}_{\mu\rho}+4g_{\mu\lambda}s^{\lambda}F^{a\mu\nu}$ (130)
follows immediately.

### v2

@@
 In this article, we extend the formalism to physical models, of both particles and fields, described by singular Lagrangians.
-Our treatment of classical field theory is based on the manifestly covariant Hamilton De-Donder Weyl formalism, in which the Lagrangian density is introduced as a bundle morphism on the pre-multisymplectic velocity phase space $J^{1}E$.
+Our treatment of classical field theory is based on the manifestly covariant De-Donder Weyl formalism, in which the Lagrangian density is introduced as a bundle morphism on the pre-multisymplectic velocity phase space $J^{1}E$.
 The results obtained are subsequently applied to a number of physically-motivated examples, as well as a discussion presented on the implications of our work for classical General Relativity.
 I Introduction
-Singular field theories are, in many cases, amongst the most interesting class of models encountered in the description of natural phenomena [1, 2, 3].
+Singular field theories are often amongst the most interesting class of models encountered in the description of natural phenomena [1, 2, 3].
 Indeed, the most precisely tested mathematical theory to date, the Standard Model of Particle Physics, is a chiral quantum field theory, gauged under local transformations of the group $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ [4, 5, 6].
-Further, it is postulated that any UV-complete theory of quantum gravity may possess no global symmetries, and that all symmetries must therefore be either spontaneously broken, or gauged [7, 8, 9].
+Further, it is conjectured that any UV-complete theory of quantum gravity may possess no global symmetries, and that all symmetries must therefore be either spontaneously broken, or gauged [7, 8, 9].
 It is thus clear that singular theories are of central importance to the advancement of modern theoretical physics.
-Any attempt to rigorously develop a quantum theory requires a firm understanding of the classical counterpart upon which it is based.
-As such, in this article, we shall present a framework in which singular theories that possess scaling symmetries may be reduced to a lower-dimensional, dynamically-equivalent description, that is frictional in nature [10].
-For theories of particles, we take, as our starting point, the geometrical constraint algorithm for pre-symplectic systems, developed by Nester et.
-al.
-in [11, 12, 13].
-This is presented in section (III.1), following a brief review of elementary pre-symplectic geometry in section (II).
+Any attempt to rigorously develop a quantum theory requires a firm understanding of the underlying classical formalism.
+In this article, we present a framework in which singular theories that possess scaling symmetries may be reduced to a lower-dimensional, dynamically-equivalent description, that is frictional in nature [10].
+For theories of particles, we utilise the pre-symplectic geometrical constraint algorithm developed in [11, 12, 13].
+This is presented in section (III.1), following a review of elementary pre-symplectic geometry in section (II).
 Having introduced this preliminary contextual framework, we provide an overview of contact and pre-contact geometry, and how this is used to describe the dynamical evolution of non-conservative systems.
@@
 Suppose that $Q$ is an $n$-dimensional smooth manifold, corresponding to the configuration space of some mechanical system.
-Local coordinates on $Q$ are taken to be $q^{i}$, and we introduce a Lagrangian function $L:TQ\rightarrow\mathbb{R}$ on the tangent bundle [35, 36, 37].
-Fibre coordinates of $TQ$ are denoted $v^{i}$, and we have the bundle projection $\pi:TQ\rightarrow Q$.
-The tangent bundle is endowed with a number of canonical geometrical features, including the vertical endomorphism $J:T(TQ)\rightarrow T(TQ)$, satisfying $J^{2}=0$, and $\textrm{rank}\,J=n$, which constitutes an almost-tangent structure on $TQ$ [38, 39].
-Additionally, we have a distinguished Liouville vector field $\Delta\in\mathfrak{X}^{\infty}(TQ)$, defined such that its one-parameter family of transformations generates dilatations of the fibres of $TQ$.
-In local bundle coordinates $(q^{i},v^{i})$, we have
-$J=\frac{\partial}{\partial v^{i}}\otimes dq^{i}\quad\quad\quad\quad\Delta=v^{i}\,\frac{\partial}{\partial v^{i}}$ (1)
-Given a Lagrangian function $L$, the canonical structures $J$ and $\Delta$ are used to define a (pre-)symplectic Lagrangian system.
-In particular, the function $E_{L}:=\Delta(L)-L$ corresponds to the scalar energy of the mechanical system described by $L$, whilst $\theta_{L}:=\,^{t}\hskip-1.70717ptJ\circ dL$ is used to construct, the Poincaré-Cartan form $\omega_{L}:=-\,d\theta_{L}$.
-Clearly, $\omega_{L}$ is closed; non-degeneracy, on the other hand, is not assured, and throughout subsequent sections, we shall be interested in those systems for which $\omega_{L}$ is degenerate.
-In such cases, we refer to the triple $(TQ,\omega_{L},E_{L})$ as a pre-symplectic Lagrangian system [40], and from (1) it follows that
-$E_{L}=\frac{\partial L}{\partial v^{i}}v^{i}-L\quad\quad\quad\quad\omega_{L}=-\,d\left(\frac{\partial L}{\partial v^{i}}dq^{i}\right)$ (2)
+A Lagrangian function $L:TQ\rightarrow\mathbb{R}$ is introduced on the tangent bundle, upon which local coordinates are denoted $(q^{i},v^{i})$ [35, 36, 37].
+Throughout, the cases of greatest interest will be those in which the triple $(TQ,\omega_{L},E_{L})$ defines a pre-symplectic system [38].
+In local coordinates, the closed 2-form $\omega_{L}$ and energy function $E_{L}$ are expressed as
+$E_{L}=\frac{\partial L}{\partial v^{i}}v^{i}-L\quad\quad\quad\quad\omega_{L}=-\,d\theta_{L}=-\,d\left(\frac{\partial L}{\partial v^{i}}dq^{i}\right)$ (1)
 The degeneracy of the 2-form $\omega_{L}$ may be expressed as the requirement that the matrix of second derivatives
 $W_{ij}:=\left(\frac{\partial^{2}L}{\partial v^{i}\partial v^{j}}\right)$
-be of non-maximal rank, and thus singular. The equations of motion for a Lagrangian system $(TQ,\omega_{L},E_{L})$ are formulated geometrically via the following vector bundle morphism
-$\begin{split}\flat_{\scriptscriptstyle L}\;:\;T(TQ)\,&\longrightarrow\,T^{*}(TQ)\\ X\;&\longmapsto\;\iota_{X}\omega_{L}\end{split}$ (3)
-It should be noted that when $\omega_{L}$ is degenerate, $\flat_{\scriptscriptstyle L}$ is not an isomorphism.
-The dynamics are deduced seeking a vector field $X_{L}\in\mathfrak{X}^{\infty}(TQ)$, such that
-$\flat_{\scriptscriptstyle L}(X_{L})=\iota_{X_{L}}\omega_{L}=dE_{L}$ (4)
-When $(TQ,\omega_{L},E_{L})$ is a symplectic system, such vector fields are guaranteed to exist, and are unique.
-In addition, integral curves of $X_{L}$ project down to curves on the configuration space $Q$, and these latter curves correspond to physical trajectories of the mechanical system.
-Such vector fields are interchangeably referred to as holonomic or second-order differential equations (SODE) [41].
-When the system of interest possesses a singular Lagrangian, not only are vector field solutions of (4) not guaranteed to exist, but since $\textrm{ker}\,\omega_{L}$ is non-trivial, any solution that is found is no longer unique [40].
-In the most favourable of cases, there will exist a submanifold $P_{f}\hookrightarrow TQ$, upon which (4) admits solutions, and this space is deduced through use of a geometrical constraint algorithm [12, 13].
-For the purposes of studying the contact reduction of singular systems, we shall work almost exclusively within the Hamiltonian formalism.
-Our motivation for this is twofold: on the one hand, the Hamiltonian framework provides a well-defined bracket structure, which readily allows us to classify the constraint functions, and deduce the dynamical evolution of any phase space variable.
-Further, not only does the Lagrangian formalism lack such a bracket, but the geometrical constraint algorithm introduces a myriad of additional complications, which arise owing to a need to impose that, at each step, the vector field solution satisfy the second-order problem.
+be of non-maximal rank, and thus singular.
+When studying the contact reduction of singular systems, we shall work almost exclusively within the Hamiltonian formalism.
+Our motivation for this is twofold: on the one hand, the Hamiltonian framework provides a well-defined bracket structure, which readily allows us to classify the constraint functions, and compute the dynamical evolution of any phase space variable.
+Further, not only does the Lagrangian formalism lack such a bracket, but the geometrical constraint algorithm introduces additional complications, which arise owing to a need to impose that, at each step, the solution satisfy the second-order problem.
 Since the dynamical content of the two formalisms is identical, we shall evade the undesirable features of the Lagrangian description by using it only as a means to obtain the corresponding Hamiltonian function.
@@
 We shall refer to the triple $(M_{0},\omega_{0},H_{0})$ as a pre-symplectic Hamiltonian system, defined on the primary constraint manifold $M_{0}$.
-In order to formulate the dynamical problem, the vector bundle morphism $\flat:T(T^{*}Q)\rightarrow T^{*}(T^{*}Q)$ is restricted to $M_{0}$, and the resulting map denoted $\flat_{0}$.
-We then seek a vector field $X_{H}$, such that
+The dynamical problem is formulated via the introduction of a bundle morphism $\flat:T(T^{*}Q)\rightarrow T^{*}(T^{*}Q)$, with $\flat(X):=\iota_{X}\omega$; this is then restricted to $M_{0}$, with the resulting map being denoted $\flat_{0}$.
+We now seek a vector field $X_{H}$, such that
 $\flat_{0}(X_{H})=\iota_{X_{H}}\omega_{0}=dH_{0}$ (2)
-As in the Lagrangian case, solutions generally do not exist on the entirety of $M_{0}$, thereby necessitating the implementation of a constraint algorithm, which is described in detail in section (III.1).
-The final constraint manifold $M_{f}$ is then the maximal subspace of the cotangent bundle upon which the dynamical problem (5) possesses well-defined (albeit non-unique) solutions.
+Solutions to this equation generally do not exist on the entirety of $M_{0}$, thereby necessitating the implementation of a constraint algorithm.
+Such a procedure gives the maximal subspace of the cotangent bundle upon which the dynamical problem (2) possesses well-defined (albeit non-unique) solutions.
 III The Pre-symplectic Constraint Algorithm
@@
 III.1 The Hamiltonian Algorithm
-When analysing the dynamics of singular systems, we shall use the almost-regular Lagrangian $L:TQ\rightarrow\mathbb{R}$ to obtain the pre-symplectic Hamiltonian system $(M_{0},\omega_{0},H_{0})$, introducing the embedding $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$.
+Let $(M_{0},\omega_{0},H_{0})$ be the pre-symplectic Hamiltonian system, obtained from an almost regular Lagrangian $L$.
 It will often be convenient to describe $M_{0}$ as the zero-set of a collection of primary constraint functions $\{\phi^{a}\}$, with $a=1,\,\cdots,n-k$, in which $n=\textrm{dim}\,Q$, and $k=\textrm{rank}\,W_{ij}$ [40].
-We shall assume throughout that all subspaces generated are regular, closed submanifolds of $T^{*}Q$, whose natural injections are embeddings.
+We shall assume that all subspaces generated are regular, closed submanifolds of $T^{*}Q$, whose natural injections are embeddings, and are denoted $\jmath_{i}:M_{i}\hookrightarrow M_{i-1}$.
 We therefore begin by seeking a vector field $X_{H}$ which satisfies
 $\flat_{0}(X_{H})=\iota_{X_{H}}\omega_{0}=dH_{0}$ (5)
-Since $\omega_{0}$ is pre-symplectic, $\flat_{0}$ is not an isomorphism; consequently, an initial prerequisite for our geometrical equation to possess solutions is that, at every point, $dH_{0}$ must lie in the range of the map $\flat_{0}$.
-Given that such a condition is not guaranteed to be satisfied a priori, the first step of the constraint algorithm requires us to restrict $M_{0}$ to the following subset
-$M_{1}:=\{x\in M_{0}\;|\;(dH_{0})_{x}\in\,\flat_{0}(T_{x}M_{0})\,\}$ (9)
-From the discussion above, under the assumptions of reflexivity and topological closedness, we recall that the image $\flat_{0}(T_{x}M_{0})$ coincides with the annihilator of $T_{x}M_{0}^{\perp}$; this is particularly noteworthy, as $T_{x}M_{0}^{\perp}$ is precisely $\textrm{ker}\,(\omega_{0})_{x}$, which is readily deduced in local coordinates.
-Thus, from a pragmatic standpoint, it is preferable to characterise $M_{1}$ as follows
+Since $\omega_{0}$ is pre-symplectic, $\flat_{0}$ is not an isomorphism, and so we should only consider those points of $M_{0}$ at which $dH_{0}\in\textrm{Im}\,\flat_{0}$.
+From above, under the assumptions of reflexivity and topological closedness, we recall that $\flat_{0}(T_{x}M_{0})=T_{x}M_{0}^{\perp}$.
+However, $T_{x}M_{0}^{\perp}$ is precisely $\textrm{ker}\,(\omega_{0})_{x}$, and so the subset of $M_{0}$ of interest is
 $M_{1}=\{x\in M_{0}\;|\;\langle(dH_{0})_{x},\textrm{ker}\,(\omega_{0})_{x}\rangle=0\,\}$ (6)
@@
 If this were not the case, the solution would tend to evolve off the constraint surface, and thus cease to obey the restrictions placed upon the system.
-Tangency of the solution is not guaranteed, and so must be imposed as an additional constraint, leading us to consider only those points of $M_{1}$ at which the solution may be written as $X_{H}=\jmath_{1*}Z$ for some $Z\in TM_{1}$; generally, $\jmath_{k}:M_{k}\hookrightarrow M_{k-1}$ is used to refer to the inclusion of $M_{k}$.
-The condition $X_{H}=\jmath_{1*}Z$ is not readily imposed locally, and so we shall use the following alternative criterion to characterise $M_{2}$
+Tangency of the solution is not guaranteed, and so must be imposed as an additional constraint, leading us to consider only the following subset of points of $M_{1}$ [11]
 $M_{2}:=\{x\in M_{1}\;|\;\langle(dH_{0})_{x},T_{x}M_{1}^{\perp}\rangle=0\,\}$ (8)
-That this provides an equivalent definition to that given by imposing $X_{H}=\jmath_{1*}Z$ follows from a series of simple calculations, outlined in [11].
 Vector field solutions on $M_{2}$ are now tangent to $M_{1}$, but, in general, not to $M_{2}$, requiring this condition to be imposed as a further restriction.
-It is then clear from (9) and (12) how the algorithm must proceed: at the $k^{\textrm{th}}$ step, the submanifold $M_{k}$ is composed of all points of $M_{k-1}$ at which the solution is tangent to $M_{k-1}$, but not necessarily to $M_{k}$, leading us to restrict $M_{k}$ to the smaller set $M_{k+1}$.
-This condition may be imposed requiring that $X_{H}=\jmath_{k*}Z$ for some $Z\in TM_{k}$, or equivalently, by demanding that $dH_{0}$ belong to the annihilator $(TM_{k}^{\perp})^{o}$, so that
+It is then clear how the algorithm must proceed: at the $k^{\textrm{th}}$ step, the submanifold $M_{k}$ is composed of all points of $M_{k-1}$ at which the solution is tangent to $M_{k-1}$, but not necessarily to $M_{k}$, leading us to restrict $M_{k}$ to the smaller set $M_{k+1}$.
+This condition may be imposed requiring that $dH_{0}$ belong to the annihilator $(TM_{k}^{\perp})^{o}$, so that
 $M_{k+1}=\{x\in M_{k}\;|\;\langle(dH_{0})_{x},T_{x}M_{k}^{\perp}\rangle=0\,\}$ (9)
-The constraint procedure may terminate in one of three ways.
-Firstly, there may be some $N\in\mathbb{N}$, for which $M_{N}=\emptyset$; in this case, the equation of motion (8) possesses no solution and the system $(M_{0},\omega_{0},H_{0})$ does not describe any physical configuration.
-Alternatively, the algorithm may yield a submanifold $M_{N}$, which is non-empty, but of dimension zero; such a space may be identified as a collection of isolated points, which forces the vector field solution to be identically zero.
-Evidently, neither of these eventualities is particularly interesting from a physical perspective.
-Consequently, we shall only consider those systems for which there exists some $N\in\mathbb{N}$, such that $M_{N+1}=M_{N}$, with $\textrm{dim}\,M_{N}\neq 0$.
-Upon deducing the value of $N$ for which the algorithm terminates, we shall declare $M_{f}:=M_{N}$ to be the final constraint manifold.
+The constraint procedure may terminate in one of three ways [11]; however, the only dynamically-interesting eventuality is that there exists some $N\in\mathbb{N}$, such that $M_{N+1}=M_{N}$, with $\textrm{dim}\,M_{N}\neq 0$.
+Upon deducing the value of $N$ for which the algorithm terminates, we shall declare $M_{f}:=M_{N}$ to be the final constraint submanifold.
 III.2 Geometry of the Final Constraint Submanifold
-The algorithm of the previous section provides a systematic way to restrict the set of points of the cotangent bundle at which the equations of motion of a singular system admit everywhere-tangent solutions.
-At no point was it necessary to specify the nature of the constraints defining each subspace; once the final submanifold has been found, however, the dynamical evolution of observable quantities is deduced locally, which does require the distinction between first and second class constraints to be made.
-Suppose that the constraint surface $M_{f}$ is defined as the zero-set of the functions $\phi^{\alpha}$, with $\alpha=1,\,\cdots,D$.
-We introduce the matrix $J$, whose entries are the Poisson brackets of the constraints i.e $J^{\alpha\beta}:=\{\phi^{\alpha},\phi^{\beta}\}$.
+Suppose that the constraint surface $M_{f}$, deduced from the algorithm above, is defined as the zero-set of the functions $\phi^{\alpha}$, with $\alpha=1,\,\cdots,D$.
+We introduce the matrix of Poisson brackets $J^{\alpha\beta}:=\{\phi^{\alpha},\phi^{\beta}\}$.
 If $\textrm{rank}\,J=D-F$, it follows that there exist $F$ $C^{\infty}$-linear combinations of the $\phi^{\alpha}$ which generate unobservable (gauge) transformations.
@@
 The goal is then to identify such combinations, corresponding to first class functions, and to modify the original set of constraints, so as to extract a maximal subfamily of second class functions.
-Note that, since $J$ is calculated from the Poisson bracket, it follows that $\textrm{rank}\,\omega_{f}=D-F$, in which $\omega_{f}:=\iota^{*}\omega$ is the closed 2-form naturally inherited by $M_{f}$, and $\imath:=\jmath_{f}\circ\cdots\circ\jmath_{0}$ defines the embedding of $M_{f}$ into $T^{*}Q$.
-In order to extract the first class functions, we introduce a ‘vector’ $v_{\alpha}=(v_{1},\,\cdots,v_{D})$, and demand that
+To this end, we introduce a ‘vector’ $v_{\alpha}=(v_{1},\,\cdots,v_{D})$, and demand that
 $\left(v_{\alpha}J^{\alpha\beta}\right)|_{M_{f}}=0\quad\textrm{for all}\;\beta=1,\,\cdots,D$ (10)
-It is sufficient to seek these null vectors on the constraint manifold, since a function $f$ is only required to have weakly-vanishing Poisson bracket with the $\phi^{\alpha}$, to be characterised as first class.
-Having identified all linearly independent combinations satisfying (14), indexing the resulting vectors via $v^{(a)}$, the $a^{\textrm{th}}$ first class constraint is simply $\Omega^{a}:=v^{(a)}_{\alpha}\phi^{\alpha}$.
+Having identified all such linearly independent combinations, indexing the resulting vectors via $v^{(a)}$, the $a^{\textrm{th}}$ first class constraint is simply $\Omega^{a}:=v^{(a)}_{\alpha}\phi^{\alpha}$.
 In order to extract the maximal subfamily of second class constraints, we seek DFD-F linearly independent vectors $w^{(j)}$, which are not in the kernel of $J$, so that
@@
 $\{f,g\}_{D}:=\{f,g\}-\{f,\chi^{i}\}\,C_{ij}\{\chi^{j},g\}$ (12)
-The Dirac bracket has a number of noteworthy properties:
-  • $\{f,gh\}_{D}=\{f,g\}_{D}\,h+g\{f,h\}_{D}$ for all $f,g,h\in C^{\infty}(T^{*}Q)$
-  • $\{f,\chi^{j}\}_{D}=0$ for any $f\in C^{\infty}(T^{*}Q)$
-  • $\{f,g\}_{D}\approx\{f,g\}$ for $f$ first class and $g$ arbitrary
-We now impose the second class constraints $\chi^{j}=0$ as strong equalities, which defines a hypersurface
+Imposing the second class constraints $\chi^{j}=0$ as strong equalities, we obtain a hypersurface
 $M_{\chi}:=\{x\in T^{*}Q\;|\;\chi^{j}(x)=0\quad\textrm{for}\;j=1,\,\cdots,D-F\,\}\;\subset\;T^{*}Q$ (13)
-with inclusion $\jmath_{\chi}:M_{\chi}\hookrightarrow T^{*}Q$, which inherits a closed 2-form $\omega_{\chi}:=\jmath^{*}_{\chi}\omega$ which is symplectic.
+which inherits a closed 2-form $\omega_{\chi}$ which is symplectic.
 The Dirac bracket is precisely the algebraic structure required to ‘transfer’ this symplectic structure to functions defined on the whole of $T^{*}Q$.
@@
 $M_{\scriptscriptstyle\Omega}:=\{x\in M_{\chi}\;|\;\Omega^{a}(x)=0\quad\textrm{for}\;a=1,\,\cdots,F\,\}$ (14)
-$M_{\scriptscriptstyle\Omega}$ is an embedded submanifold of $M_{\chi}$, with inclusion $\jmath_{\scriptscriptstyle\Omega}:M_{\scriptscriptstyle\Omega}\hookrightarrow M_{\chi}$; the 2-form $\omega_{\scriptscriptstyle\Omega}:=\jmath^{*}_{\scriptscriptstyle\Omega}\,\omega_{\chi}$ is maximally degenerate, since $M_{\scriptscriptstyle\Omega}$ is first class.
-It is straightforward to show that the Hamiltonian vector fields $X_{a}$ corresponding to $\Omega^{a}$ via $\iota_{X_{a}}\omega_{\chi}=d\Omega^{a}$ are tangent to $M_{\scriptscriptstyle\Omega}$, since
-$\mathcal{L}_{X_{a}}\Omega^{b}=d\Omega^{b}[X_{a}]=\omega_{\chi}(X_{b},X_{a})=\{\Omega^{b},\Omega^{a}\}\approx 0$
-Here, the weak equality is taken to signify ‘when restricted to $M_{\scriptscriptstyle\Omega}$.’ In general, a vector field $\xi\in\mathfrak{X}^{\infty}(T^{*}Q)$ may only be tangent to $M_{\scriptscriptstyle\Omega}$ if $\mathcal{L}_{\xi}\Omega^{a}\approx 0$ for all $\Omega^{a}$: that is if $\xi\in\textrm{ker}\,\omega_{\scriptscriptstyle\Omega}$; since the first class constraints are assumed to be independent (by which we mean that their differentials are linearly independent) it follows that $\textrm{ker}\,\omega_{\scriptscriptstyle\Omega}$ is spanned by the Hamiltonian vector fields $X_{a}$.
-More formally, at each point $x\in M_{\scriptscriptstyle\Omega}$, the symplectic orthogonal
-$T_{x}M_{\scriptscriptstyle\Omega}^{\perp}:=\{v\in T_{x}M_{\chi}\;|\;\omega_{\chi}(v,w)=0\quad\textrm{for all}\;w\in T_{x}M_{\scriptscriptstyle\Omega}\,\}$
-satisfies $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}\subseteq T_{x}M_{\scriptscriptstyle\Omega}$, and so $M_{\scriptscriptstyle\Omega}$ constitutes a coisotropic submanifold of $M_{\chi}$ [48]; for such a class of space, the distribution $x\mapsto T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ is involutive, since, for any two $X,Y\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, we take some $Z\in T_{x}M_{\scriptscriptstyle\Omega}$, and compute
-$d\omega_{\chi}(X,Y,Z)=X\omega_{\chi}(Y,Z)-Y\omega_{\chi}(Z,X)+Z\omega_{\chi}(X,Y)-\omega_{\chi}([X,Y],Z)+\omega_{\chi}([X,Z],Y)-\omega_{\chi}([Y,Z],X)$
-Since $\omega_{\chi}$ is closed, $d\omega_{\chi}(X,Y,Z)=0$; moreover, $X$ and $Y$ were taken to be in the symplectic orthogonal $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, so that the first three terms on the right-hand side vanish.
-Similar considerations lead us to conclude that
-$d\omega_{\chi}(X,Y,Z)=-\,\omega_{\chi}([X,Y],Z)=0$
-so that if $X,Y\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$, then $X,Y]\in T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$.
-Hence, $x\mapsto T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ is an involutive distribution, which, locally, we may write as the span of the $X_{a}$ [49]: that is to say
+$M_{\scriptscriptstyle\Omega}$ is an embedded submanifold of $M_{\chi}$, whose 2-form $\omega_{\scriptscriptstyle\Omega}$ is maximally degenerate.
+Further, at each point $x\in M_{\scriptscriptstyle\Omega}$, the symplectic orthogonal $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ satisfies $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}\subseteq T_{x}M_{\scriptscriptstyle\Omega}$, and so $M_{\scriptscriptstyle\Omega}$ constitutes a coisotropic submanifold of $M_{\chi}$ [45].
+For such a class of space, the distribution $x\mapsto T_{x}M_{\scriptscriptstyle\Omega}^{\perp}$ is involutive, and may locally be written as the span of the Hamiltonian vector fields $X_{a}$ corresponding to $\Omega^{a}$ via $\iota_{X_{a}}\omega_{\chi}=d\Omega^{a}$ [46]: that is to say
 $T_{x}M_{\scriptscriptstyle\Omega}^{\perp}=\textrm{span}(X_{1}|_{x}\,,\cdots,X_{F}|_{x})\,\subseteq\,T_{x}M_{\scriptscriptstyle\Omega}$
-Since the Hamiltonian vector fields of the first class constraints define an involutive distribution, we know, from the Frobenius theorem, that there exists a foliation $\mathcal{F}$ of $M_{\scriptscriptstyle\Omega}$ into $F$-dimensional submanifolds, which, in this context, we refer to as gauge leaves [50, 51].
+Since the Hamiltonian vector fields of the first class constraints define an involutive distribution, we know that there exists a foliation of $M_{\scriptscriptstyle\Omega}$ into $F$-dimensional gauge leaves [47, 48].
 The gauge orbits correspond to integral curves of the $X_{a}$, and they map points of $M_{\scriptscriptstyle\Omega}$ into other points, which produce dynamically indistinguishable configurations.
-Note that these results are valid only on $M_{\scriptscriptstyle\Omega}$: elsewhere, the vector fields $X_{a}$ do not, in general, close under the Lie bracket.
-As a result of the degeneracy of $\omega_{\scriptscriptstyle\Omega}$, the constraint surface does not inherit a well-defined bracket structure.
-It is therefore necessary to construct a reduced space, quotienting out by the gauge orbits; this reduced phase space identifies all points related by gauge transformations, and thus contains only configurations which are physically inequivalent.
-Upon constructing this space, there arise two distinct possibilities: in the most favourable of cases, the algebra formed by the first class constraints is a genuine Lie algebra; this occurs precisely when the structure functions $f^{ab}_{c}$, appearing in $\{\Omega^{a},\Omega^{b}\}=f^{ab}_{c}\Omega^{c}$, are constant over the whole of $M_{\scriptscriptstyle\Omega}$.
-In these conditions, the finite gauge transformations are generated by the elements of the corresponding Lie group $G$ that are path-connected to the identity, [^2] This lifting of the elements of the Lie algebra to the group is, of course, provided by the exponential map.
-and they give rise to a smooth action of $G$ on $M_{\scriptscriptstyle\Omega}$.
-It is then possible to perform a reduction procedure, using results similar to those of the Marsden-Weinstein theorem [52, 53], to obtain a phase space $M_{\scriptscriptstyle\Omega}/G$, which is symplectic [54, 55].
-In most cases, such conditions will not be realised; it is often found that the Dirac bracket between two first class constraints is strongly equal to some combination of the phase space variables which vanishes on $M_{\scriptscriptstyle\Omega}$.
-The structure functions are then of the form $f^{ab}_{c}(q^{i},p_{i})$, and they define a Lie algebroid on $M_{\scriptscriptstyle\Omega}$ [56].
-Formally, it is again possible to quotient out by the gauge orbits, obtaining a physical phase space that is symplectic; however it is almost invariably more straightforward to introduce a choice of gauge-fixing.
-Since we are only concerned with the dynamical evolution of gauge-invariant observable quantities, the precise choice of gauge-fixing is somewhat arbitrary, provided that the gauge is obtainable via a finite number of infinitesimal transformations, and that it fixes all degrees of freedom, leaving no remnant gauge symmetry [57, 58].
-Upon carrying out the pre-symplectic reduction process, we obtain a physical state space $\mathcal{P}$, together with a projection map $\pi:M_{\scriptscriptstyle\Omega}\rightarrow\mathcal{P}$, sending each point on the constraint surface to its equivalence class under gauge orbits.
-The reduced phase space $\mathcal{P}$ is symplectic, with associated 2-form $\pi^{*}\omega_{\scriptscriptstyle\mathcal{P}}=\jmath^{*}\omega$, where $\jmath:=\jmath_{\chi}\circ\jmath_{\scriptscriptstyle\Omega}:M_{\Omega}\rightarrow T^{*}Q$ defines the embedding of $M_{\scriptscriptstyle\Omega}$ into $T^{*}Q$.
+Quotienting out by these gauge orbits, we obtain a physical phase space $\mathcal{P}$, which inherits a closed 2-form $\omega_{\scriptscriptstyle\mathcal{P}}$ that is symplectic.
 IV Contact Geometry
-In order to examine the commutative relationship between contact reduction by scaling symmetries, and the restriction of the phase space of mechanical systems described by singular Lagrangians, it will be necessary to understand how the geometric algorithm we have presented for pre-symplectic manifolds is to be adapted to the pre-contact spaces, which naturally arise upon eliminating a scaling degree of freedom [14, 15, 10].
-In anticipation of this, we shall briefly introduce the fundamental ideas of contact and pre-contact geometry [59, 60], together with their implementation in the description of Lagrangian and Hamiltonian systems [61, 62, 63].
-IV.1 Contact Manifolds
-A contact structure on a smooth manifold $C$, of dimension $2n+1$, is a maximally-non-integrable distribution $\xi\subset TC$; locally, this distribution may be described as the kernel of some $\eta\in\Omega^{1}(U\subset C)$, referred to as a contact form.
-If the quotient line bundle $TC/\xi\rightarrow C$ is trivial, then $\xi$ is coorientable, and $\eta$ may be extended to the whole of $C$ [64].
-The non-integrability condition of $\xi$ may be expressed as the requirement that
-$\eta\wedge d\eta^{n}\neq 0$ (19)
-over the whole of $C$.
-The coorientability of $\xi$ implies that $C$ itself is orientable; in this case, $\eta\wedge d\eta^{n}$ provides a volume form on $C$, and since, in general, $\xi$ is of rank $2n$, at each point $p\in C$, we have the decomposition
-$T_{p}C=\textrm{ker}\,\eta_{p}\oplus\textrm{ker}\,d\eta_{p}$
-and $d\eta_{p}$ defines a skew-symmetric, non-degenerate, bilinear form of maximal rank on the $2n$-dimensional subspace $\xi_{p}$.
-We henceforth assume that $\xi$ is coorientable, and work with the globally-defined form $\eta$; every contact manifold $(C,\eta)$ admits a distinguished Reeb vector field $\mathcal{R}\in\mathfrak{X}^{\infty}(C)$, defined via
+Contact structures arise naturally when scaling degrees of freedom are eliminated from the ontology of theories describing mechanical systems [14, 15, 10].
+In general, a contact structure on a smooth manifold $C$, of dimension $2n+1$, is a maximally-non-integrable distribution $\xi\subset TC$; locally, this distribution may be described as the kernel of some $\eta\in\Omega^{1}(U\subset C)$, referred to as a contact form [49, 50, 51, 52, 53].
+If the quotient line bundle $TC/\xi\rightarrow C$ is trivial, as we shall assume to be the case, then $\xi$ is coorientable, and $\eta$ may be extended to the whole of $C$ [54].
+Every contact manifold $(C,\eta)$ admits a distinguished Reeb vector field $\mathcal{R}\in\mathfrak{X}^{\infty}(C)$, defined via
 $\iota_{\mathcal{R}}d\eta=0\quad\quad\quad\quad\iota_{\mathcal{R}}\eta=1$ (15)
 As for symplectic manifolds, around each point $p\in C$, we may always find a local chart of Darboux coordinates $(x^{1},\cdots,x^{n},y_{1},\cdots,y_{n},z)$, in which $\eta$ takes the form $\eta=dz-y_{i}\,dx^{i}$.
-IV.2 Lagrangian Mechanics
-Consider the $n$-dimensional configuration space $Q$, with local coordinates $(q^{i})$; the extended tangent bundle $TQ\times\mathbb{R}$ is a manifold of dimension $2n+1$, with induced bundle coordinates $(q^{i},v^{i},z)$.
-The canonical structures of $TQ$ are easily extended to $TQ\times\mathbb{R}$ [62], and have local coordinate expression (1).
-A contact Lagrangian is a function $L:TQ\times\mathbb{R}\rightarrow\mathbb{R}$, which allows us to define the Cartan forms $\theta_{L}$ and $\omega_{L}$, together with the energy function $E_{L}$.
-The space $TQ\times\mathbb{R}$ is made into a contact manifold upon introducing the 1-form $\eta_{L}:=dz-\theta_{L}$, and we refer to the triple $(TQ\times\mathbb{R},\eta_{L},E_{L})$ as a contact Lagrangian system.
-IV.3 The Legendre Map
-Whilst contact Lagrangian systems are introduced on the space $TQ\times\mathbb{R}$, their Hamiltonian counterparts are studied on the extended cotangent bundle $T^{*}Q\times\mathbb{R}$, which admits the following projections
-$\begin{split}\pi_{0}:T^{*}Q\times\mathbb{R}\rightarrow Q\times\mathbb{R}&\quad\quad\quad\quad\pi_{1}:T^{*}Q\times\mathbb{R}\rightarrow T^{*}Q\\ \rho:\,&T^{*}Q\times\mathbb{R}\rightarrow\mathbb{R}\end{split}$ (21)
-The Legendre map $FL:TQ\times\mathbb{R}\rightarrow T^{*}Q\times\mathbb{R}$ acts on the local bundle coordinates $(q^{i},p_{i},z)$ of $T^{*}Q\times\mathbb{R}$ according to
+The extended tangent bundle $TQ\times\mathbb{R}$ of the $n$-dimensional configuration space $Q$ is a manifold of dimension $2n+1$, with local coordinates $(q^{i},v^{i},z)$.
+A contact Lagrangian is a function $L:TQ\times\mathbb{R}\rightarrow\mathbb{R}$, from which we define the Cartan forms $\theta_{L}$ and $\omega_{L}$, together with the energy function $E_{L}$, exactly as on $TQ$.
+The space $TQ\times\mathbb{R}$ is made into a (pre-)contact manifold, introducing the 1-form $\eta_{L}:=dz-\theta_{L}$, and we refer to the triple $(TQ\times\mathbb{R},\eta_{L},E_{L})$ as a (pre-)contact Lagrangian system.
+If the characteristic distribution
+$\mathcal{C}:=\textrm{ker}\,\eta_{L}\,\cap\,\textrm{ker}\,d\eta_{L}\;\subseteq\;T(TQ\times\mathbb{R})$ (16)
+is of rank 2(nk)2(n-k), we say that $\eta_{L}$ is of class $2k+1$ [52, 55]; locally, this implies
+$\eta\wedge d\eta^{k}\neq 0\quad\quad\textrm{but}\quad\quad\eta\wedge d\eta^{k+1}=0$ (17)
+Clearly, when $\textrm{rank}\,\mathcal{C}=2n$, $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a regular system; in such a case, the Legendre map $FL:TQ\times\mathbb{R}\rightarrow T^{*}Q\times\mathbb{R}$ is a diffeomorphism, and acts on local coordinates $(q^{i},p_{i},z)$ of $T^{*}Q\times\mathbb{R}$ according to
 $FL^{*}q^{i}=q^{i}\quad\quad FL^{*}p_{i}=\frac{\partial L}{\partial v^{i}}\quad\quad FL^{*}z=z$ (18)
-The extended cotangent bundle inherits a contact structure, defined in terms of the canonical structures of $T^{*}Q$; in particular, if $\vartheta\in\Omega^{1}(T^{*}Q)$ and $\Omega=-\,d\vartheta\in\Omega^{2}(T^{*}Q)$ denote the Liouville 1- and 2-forms on the cotangent bundle, then we have
-$\theta=\pi_{1}^{*}\vartheta\quad\quad\textrm{and}\quad\quad\omega=-\,d\theta=\pi_{1}^{*}\Omega$ (23)
-The Legendre map relates these forms with $\theta_{L}$ and $\omega_{L}$ in the obvious manner
-$FL^{*}\theta=\theta_{L}\quad\quad\textrm{and}\quad\quad FL^{*}\omega=\omega_{L}$ (24)
-The contact 1-form $\eta$ is then $\eta=dz-\theta$, and the Reeb vector field $\mathcal{R}$ follows from (20); in local coordinates, we have
-$\theta=p_{i}\,dq^{i}\quad\quad\omega=dq^{i}\wedge dp_{i}\quad\quad\eta=dz-p_{i}\,dq^{i}\quad\quad\mathcal{R}=\frac{\partial}{\partial z}$ (25)
-In complete analogy to the symplectic case, the Legendre map provides a set of criteria, according to which we may classify a given contact Lagrangian system; in particular, we shall restrict ourselves to those singular systems which are categorised as ‘almost regular.’
-IV.4 Hamiltonian Formalism
-If $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a hyperregular contact Lagrangian system, we introduce the Hamiltonian as the unique function $H:T^{*}Q\times\mathbb{R}\rightarrow\mathbb{R}$, such that $FL^{*}H=E_{L}$.
+The space $T^{*}Q\times\mathbb{R}$ has a canonical contact form $\eta=dz-p_{i}\,dq^{i}$, and the Reeb vector field $\mathcal{R}$ follows from (15).
+If $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a hyperregular Lagrangian system, we introduce the unique function $H:T^{*}Q\times\mathbb{R}\rightarrow\mathbb{R}$, such that $FL^{*}H=E_{L}$.
 The triple $(T^{*}Q\times\mathbb{R},\eta,H)$ is then said to constitute a contact Hamiltonian system, and there exists a bundle morphism
@@
 $A^{i}=\frac{\partial H}{\partial p_{i}}\quad\quad B_{i}=-\,\biggr(\frac{\partial H}{\partial q^{i}}+p_{i}\,\frac{\partial H}{\partial z}\biggr)\quad\quad C=p_{i}\,\frac{\partial H}{\partial p_{i}}-H$ (22)
-V Pre-contact Systems
-All results presented thus far are applicable to regular and hyperregular systems; however, the main objective of this work is to present a mathematical framework within which to analyse scaling symmetries in singular theories.
-Consequently, we shall dedicate the following sections to developing a number of tools required to treat such systems.
-V.1 Jacobi Structures
+V Jacobi Structures
 Both symplectic and contact manifolds belong to a broader class of object known as Jacobi manifolds [56, 57, 58]; in general, a Jacobi manifold is a triple $(M,\Lambda,E)$, in which $\Lambda\in\mathfrak{X}^{2}(M)$ is a multivector field of degree two, and $E\in\mathfrak{X}^{\infty}(M)$ is a vector field; together, these objects satisfy [50]
@@
 $\{f,g\}_{J}=\frac{\partial f}{\partial q^{i}}\frac{\partial g}{\partial p_{i}}-\frac{\partial f}{\partial p_{i}}\frac{\partial g}{\partial q^{i}}+\left(p_{i}\frac{\partial g}{\partial p_{i}}-g\right)\frac{\partial f}{\partial z}-\left(p_{i}\frac{\partial f}{\partial p_{i}}-f\right)\frac{\partial g}{\partial z}$ (30)
-V.2 Pre-contact Manifolds
-The final piece of mathematical background required to introduce the constraint algorithm is that of pre-contact manifolds [62, 74]; suppose that $(M,\eta)$ is a $(2n+1)$-dimensional manifold, equipped with a 1-form $\eta$, such that the distribution
-$\mathcal{C}:=\textrm{ker}\,\eta\,\cap\,\textrm{ker}\,d\eta\;\subseteq\;TM$ (38)
-is of rank 2n+1p2n+1-p; in this case, we refer to $\mathcal{C}$ as the characteristic distribution, and say that $\eta$ is of class $p$. In general, if $\eta$ is of class $2k+1$, then we have
-$\eta\wedge d\eta^{k}\neq 0\quad\quad\textrm{but}\quad\quad\eta\wedge d\eta^{k+1}=0$ (39)
-In addition, around each point $p\in M$, there exist local coordinates $(x_{1},\cdots,x_{k},y^{1},\cdots,y^{k},z,w_{1},\cdots,w_{s})$, with s=2(nk)s=2(n-k), such that $\eta$ is given by $\eta=dz-y_{i}\,dx^{i}$, and the characteristic distribution is spanned by
-$\mathcal{C}\,=\,\,\biggr\langle\frac{\partial}{\partial w_{1}},\,\cdots,\frac{\partial}{\partial w_{s}}\biggr\rangle$
-Upon introducing a function $H:M\rightarrow\mathbb{R}$, we refer to the triple $(M,\eta,H)$ as a pre-contact Hamiltonian system; as in the case of regular contact systems, we introduce the vector bundle morphism $\bar{\flat}:TM\rightarrow T^{*}M$, with precisely the same action as in (26).
-Every pre-contact manifold admits a Reeb vector field $\mathcal{R}$, defined according to the conditions $\iota_{\mathcal{R}}\,d\eta=0$ and $\iota_{\mathcal{R}}\eta=1$; however, the novelty is that this field is no longer unique.
-In particular, any $X\in\mathfrak{X}^{\infty}(M)$ is a Reeb vector field if $\bar{\flat}(X)=\eta$ [60].
-Given a distribution $\mathcal{D}\subset TM$, we define the following notion of orthogonality
-$\mathcal{D}^{\perp}:=\{v\in TM\;|\;\bar{\flat}(w)(v)=0\quad\textrm{for all}\;w\in\mathcal{D}\,\}$ (40)
-Note that since $\bar{\flat}(v)(w)$ is neither symmetric or antisymmetric under the exchange of $v$ and $w$, the above definition is strictly the right complement of $\mathcal{D}$; it is of course possible to define the left complement; however, $\mathcal{D}^{\perp}$ is the space we shall use within the constraint algorithm.
 VI Pre-contact Constraint Algorithm
-We begin by considering a pre-contact Lagrangian system $(TQ\times\mathbb{R},\eta_{L},E_{L})$, of class $2k+1$, which occurs precisely when the rank of the Hessian matrix is equal to $k$ over the whole of $TQ\times\mathbb{R}$.
-The extended cotangent bundle $T^{*}Q\times\mathbb{R}$ has a canonical contact form $\eta$, whose construction was given in (23); if we introduce the inclusion $\kappa_{0}:P_{0}:=FL(TQ\times\mathbb{R})\hookrightarrow T^{*}Q\times\mathbb{R}$, then $P_{0}$ inherits the structure of a pre-contact manifold, with corresponding 1-form $\eta_{0}=\kappa_{0}^{*}\eta=FL_{0}^{*}\,\eta_{L}$, where $FL_{0}$ refers to the restriction of the Legendre map to $P_{0}$.
-This pre-contact manifold is then made into a Hamiltonian system upon introducing the unique function $H_{0}:P_{0}\rightarrow\mathbb{R}$, which satisfies $FL_{0}^{*}\,H_{0}=E_{L}$.
+When $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a pre-contact system, the above construction requires only minor modification.
+The image $P_{0}:=FL(TQ\times\mathbb{R})$ is assumed to be a closed submanifold of $T^{*}Q\times\mathbb{R}$, with inclusion $\kappa_{0}:P_{0}\hookrightarrow T^{*}Q\times\mathbb{R}$; the canonical contact form on $T^{*}Q\times\mathbb{R}$ restricts to $\eta_{0}=\kappa_{0}^{*}\eta$, and endows $P_{0}$ with a pre-contact structure.
+This is then made into a Hamiltonian system upon introducing the unique function $H_{0}:P_{0}\rightarrow\mathbb{R}$, which satisfies $FL_{0}^{*}\,H_{0}=E_{L}$, where $FL_{0}$ denotes the restriction of $FL$ to its image.
+Denoting the restriction of the bundle morphism $\bar{\flat}$ to $P_{0}$ via $\bar{\flat}_{0}$, for a distribution $\mathcal{D}\subset TP_{0}$, we have the following notion of orthogonality
+$\mathcal{D}^{\perp}:=\{v\in TP_{0}\;|\;\bar{\flat}_{0}(w)(v)=0\quad\textrm{for all}\;w\in\mathcal{D}\,\}$ (31)
 In the interest of reducing cumbersome notation in what follows, we introduce
 $\alpha_{0}:=dH_{0}-\left(\mathcal{R}(H_{0})+H_{0}\right)\eta_{0}$
-and write the restriction of the bundle morphism $\bar{\flat}$ to $P_{0}$ as $\bar{\flat}_{0}$; thus, the dynamical problem we seek to solve is that of finding a vector field $X_{H}\in\mathfrak{X}^{\infty}(P_{0})$, such that
+so that the dynamical problem to be solved is that of finding a vector field $X_{H}\in\mathfrak{X}^{\infty}(P_{0})$, such that
 $\bar{\flat}_{0}(X_{H})=\alpha_{0}$ (32)
@@
 $P_{1}:=\{x\in P_{0}\;|\;(\alpha_{0})_{x}\in\bar{\flat}_{0}(T_{x}P_{0})\,\}$ (33)
-As was the case for the pre-symplectic algorithm, it is convenient to make use of the relationship between the annihilator of the space $TP_{0}^{\perp}$ and the image of $TP_{0}$ under $\bar{\flat}_{0}$ to write $P_{1}$ as
-$P_{1}=\{x\in P_{0}\;|\;\langle\,(\alpha_{0})_{x},T_{x}P_{0}^{\perp}\,\rangle=0\,\}$ (43)
-Recall that for the pre-symplectic manifold $M_{0}$, we observed that $TM_{0}^{\perp}=\textrm{ker}\,\omega_{0}$, which led to the recharacterisation of $M_{1}$ given in (10); in the present context, we find that $TP_{0}^{\perp}=\mathcal{C}$, which provides the following, more practical condition for deducing $P_{1}$
+As was the case for the pre-symplectic algorithm, it is convenient to make use of the relationship $(TP_{0}^{\perp})^{o}=\bar{\flat}_{0}(TP_{0})$; further, it is found that $TP_{0}^{\perp}=\mathcal{C}$, which provides the following, more practical condition for deducing $P_{1}$
 $P_{1}=\{x\in P_{0}\;|\;\langle(\alpha_{0})_{x},\mathcal{C}_{x}\rangle=0\,\}$ (34)
-Upon finding a basis for the characteristic distribution, $P_{1}$ may be described as the zero-set of a collection of secondary constraint functions.
 We must now impose that any vector field solution on $P_{1}$ be tangent to this space, which generally requires us to restrict to the smaller set
@@
 The non-uniqueness follows from the observation that we may add to $X_{H}$ any element of $\mathcal{C}\,\cap\,TP_{f}$, without altering the physical dynamics or leaving the constraint surface.
-In order to deduce local expressions for the dynamical evolution of extended phase space functions, we shall suppose that $P_{f}$ may be written as the zero-set of the $N$ functions $\{\Phi^{\alpha}\}$.
+In order to present the complementary local description, we shall suppose that $P_{f}$ may be written as the zero-set of the $N$ functions $\{\Phi^{\alpha}\}$.
 Introducing the matrix of brackets $K^{\alpha\beta}:=\{\Phi^{\alpha},\Phi^{\beta}\}_{J}$, with $\textrm{rank}\,K=N-F$, we extract $F$ first class functions $\Omega^{a}$, and NFN-F second class functions $\chi^{j}$, following precisely the same procedure as in (10).
@@
 $\dot{f}=X_{H_{T}}(f)=\{f,H_{T}\}_{DJ}-f\mathcal{R}_{DJ}(H_{T})$ (38)
-Upon expanding this expression, making use of the weak Leibnitz rule, we find that
-$\dot{f}=\left(X_{H}+u_{a}X_{\Omega^{a}}\right)(f)+\Omega^{a}\Lambda_{DJ}(df,du_{a})$ (49)
-When evaluated on the constraint surface $P_{f}$, the term proportional to $\Omega^{a}$ vanishes identically, leaving
+When evaluated on the constraint surface $P_{f}$, the constraints $\Omega^{a}$ vanish identically, and we find
 $\dot{f}\approx\left(X_{H}+u_{a}X_{\Omega^{a}}\right)(f)$ (39)
@@
 $\mathfrak{L}_{D}\hskip 0.56905pt\omega=\omega\quad\quad\quad\quad\mathfrak{L}_{D}H=\Lambda H$ (40)
-Here, and throughout what follows, we use the symbol $\mathfrak{L}$ to denote the Lie derivative, reserving the calligraphic $\mathcal{L}$ for the Lagrangian density, encountered in our treatment of multisymplectic field theory.
+Here, $\mathfrak{L}$ denotes the Lie derivative; we reserve the calligraphic $\mathcal{L}$ for the Lagrangian density.
 As a consequence of (40), we see that if $X_{H}\in\mathfrak{X}^{\infty}(M)$ is the unique Hamiltonian vector field corresponding to $H$ via the relationship $\iota_{X_{H}}\omega=dH$, then
@@
 Such a transformation is the mathematical realisation of Poincaré’s thought experiment, where one imagines waking up to a world in which all distances have suffered a doubling in scale [68]: any attempt to discern the rescaled world from the original configuration will be futile, since observers have access only to ratios of dimensionful quantities.
-We shall develop these ideas somewhat further in section (XII), in the context of low-energy effective gauge theories, examining arguments presented against Poincaré’s stance that a rescaled world would be indistinguishable from our own.
-From a minimalistic viewpoint, a mathematical description whose ontology contains redundant degrees of freedom, pertaining to unobservable global scales, is highly undesirable [79, 80].
+From a minimalistic viewpoint, a mathematical description whose ontology contains redundant degrees of freedom is highly undesirable [69, 70].
 We argue that, where possible, any theory of the natural world should be formulated exclusively in terms of empirically accessible quantities, together with those (possibly unobservable) parameters, whose presence contributes to the closure of the algebra of dynamical observables.
-The identification of a scaling symmetry indicates that our mathematical framework contains just such a redundancy, and the role of ‘contact reduction’ is to make precise the heuristic notion of eliminating this superfluous structure.
+The identification of a scaling symmetry indicates that our mathematical framework contains just such a redundancy, and the role of ‘contact reduction’ is to make precise the notion of eliminating this superfluous structure.
 If the vector field $D$ is such that its flow acts freely and properly on $M$, then the quotient space $M/\hskip-2.84526pt\sim$ formed by identifying points connected by $D$-orbits has the structure of a smooth manifold of dimension $\textrm{dim}\,M-1$; moreover, the map $\pi:M\rightarrow M/\hskip-2.84526pt\sim$, sending each point of $M$ to its equivalence class, is a submersion [49, 71, 72].
@@
 generates $\pi_{*}\,\textrm{span}(X_{H})$.
-In general, a contact vector field is one whose flow preserves the contact distribution; consequently, unless the scaling symmetry of interest is of degree one (which, throughout subsequent discussion, will often be the case) or we restrict ourselves to the zero-set of $H^{c}$, the vector field $X$ does not generally preserve $\xi=\textrm{ker}\,\eta$.
+A contact vector field is one whose flow preserves the contact distribution; consequently, unless the scaling symmetry of interest is of degree one (which will often be the case) or we restrict ourselves to the zero-set of $H^{c}$, the vector field $X$ does not generally preserve $\xi=\textrm{ker}\,\eta$.
 Additionally, it is clear that this construction is dependent upon the relation $\mathfrak{L}_{D}X_{H}=(\Lambda-1)X_{H}$, which itself derives from the symplectic Hamiltonian equation of motion $\iota_{X_{H}}\omega=dH$; as a result, scaling symmetries are strictly dynamical in nature: that is to say, the contact-reduced system faithfully reproduces the original symplectic theory only when the equations of motion are satisfied.
@@
 This concludes our introductory treatment of contact reduction, and we now apply these ideas to constrained Hamiltonian systems.
-Before commencing, we remark that there are two conceivable ways of proceeding: on the one hand, given a Hamiltonian $H_{0}$, corresponding to a singular Lagrangian, we may apply the techniques of section (III.1), deduce the final constraint manifold, eliminating any gauge degrees of freedom, and then seek a scaling symmetry, to make the contact reduction.
-Alternatively, we might first identify a scaling degree of freedom within the unconstrained Hamiltonian system, excise this from our ontology, and then use the pre-contact constraint algorithm, to deduce the subset of the reduced space, upon which solutions to the contact equations of motion are well-defined.
-More succinctly, in what follows, we would like to examine the commutativity of ‘restriction &\& reduction’ and ‘reduction &\& restriction.’
+Note that there are two conceivable ways of proceeding: on the one hand, given a Hamiltonian $H_{0}$, corresponding to a singular Lagrangian, we may apply the techniques of section (III.1), deduce the final constraint manifold, and then make the contact reduction.
+Alternatively, we might first identify a scaling degree of freedom within the unconstrained Hamiltonian system, excise this from our ontology, and then use the pre-contact constraint algorithm, to deduce the subset of the reduced space upon which solutions to the contact equations of motion are well-defined.
 VII.2 Reduction &\& Restriction
-Let us suppose that $(TQ,\eta_{L},E_{L})$ is an almost-regular Lagrangian system; we begin by deducing the canonical Hamiltonian $H_{0}:M_{0}\rightarrow\mathbb{R}$, defined on the primary constraint manifold $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$.
+Suppose that $(TQ,\eta_{L},E_{L})$ is an almost-regular Lagrangian system; we begin by deducing the canonical Hamiltonian $H_{0}:M_{0}\rightarrow\mathbb{R}$, defined on the primary constraint manifold $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$.
 This is then extended to a function $H:T^{*}Q\rightarrow\mathbb{R}$ on the full phase space, such that $H|_{M_{0}}=H_{0}$.
@@
 We highlight that this construction does not directly give rise to a pre-contact manifold; instead, we have symmetry reduced the full phase space, into which the pre-symplectic system is embedded.
-Had we sought a vector field on $M_{0}$ as our scaling symmetry, it is not clear that we should always obtain a well-defined pre-contact form on the quotient space.
-By working on the full phase space, we have the map $\beta:T^{*}Q\rightarrow C$, which allows us to construct functions $\gamma^{\alpha}\in C^{\infty}(C)$, whose zero-set defines a submanifold $\kappa_{0}:C_{0}\hookrightarrow C$, which does inherit a pre-contact structure; in particular, we have
+The map $\beta:T^{*}Q\rightarrow C$ allows us to construct functions $\gamma^{\alpha}\in C^{\infty}(C)$, whose zero-set defines a submanifold $\kappa_{0}:C_{0}\hookrightarrow C$, which does inherit a pre-contact structure; in particular, we have
 $C_{0}:=\{y\in C\;|\;\gamma^{\alpha}(y)=0\,\}\quad\quad\textrm{with}\quad\beta^{*}\gamma^{\alpha}:=\frac{\phi^{\alpha}}{\rho^{\Lambda}}$ (47)
@@
 VII.3 Restriction &\& Reduction
-Having shown how we obtain the contact Hamiltonian equations of motion by first contact reducing the pre-symplectic system, and then applying the appropriate constraint algorithm, we shall now consider the reverse process, in which we first carry out the restriction at the symplectic level, before seeking a scaling symmetry with which to contact reduce the constrained theory.
-We begin, once again, from the pre-symplectic system $(M_{0},\omega_{0},H_{0})$, and carry out the constraint procedure presented in section (III.1), deducing the final submanifold $M_{f}\hookrightarrow T^{*}Q$.
-In order to be able to carry out a contact reduction via scaling symmetry, we must either introduce a particular choice of gauge-fixing, or formally quotient out by the action of the gauge transformations; since we are concerned with the dynamics of observable ($i.e$ gauge-invariant) quantities, this choice is somewhat inconsequential, and we denote the physical phase space $\mathcal{P}$, with $\jmath_{\scriptscriptstyle\mathcal{P}}:\mathcal{P}\hookrightarrow M_{f}$.
+In order to carry out the reverse process, in which we seek a scaling symmetry within the constrained theory, we take the pre-symplectic system $(M_{0},\omega_{0},H_{0})$, and implement the procedure presented in section (III.1), deducing the final submanifold $M_{f}\hookrightarrow T^{*}Q$.
+In order to be able to carry out a contact reduction, we must either introduce a choice of gauge-fixing, or formally quotient out by the action of the gauge transformations; since we are concerned with the dynamics of observable, gauge-invariant quantities, this choice is somewhat inconsequential, and we denote the physical phase space $\mathcal{P}$, with $\jmath_{\scriptscriptstyle\mathcal{P}}:\mathcal{P}\hookrightarrow M_{f}$.
 From above, we know that $\mathcal{P}$ is symplectic, with non-degenerate 2-form $\omega_{\scriptscriptstyle\mathcal{P}}$ and Hamiltonian $H_{\scriptscriptstyle\mathcal{P}}$.
@@
 From the constructions given above, it should be clear that the pre-symplectic and pre-contact constraint algorithms run in parallel.
-At the $i^{\textrm{th}}$ stage of the pre-symplectic algorithm, the submanifold $M_{i}$, with inclusion $\jmath_{i}:M_{i}\hookrightarrow M_{i-1}$, may be described as the zero-set of the functions $\phi^{\alpha}_{(i)}$.
-The projection of these $\phi^{\alpha}_{(i)}$ defines a second set of functions $\gamma^{\alpha}_{(i)}$ according to
+At the $i^{\textrm{th}}$ stage of the former, the submanifold $M_{i}$, with inclusion $\jmath_{i}:M_{i}\hookrightarrow M_{i-1}$, may be described as the zero-set of the functions $\phi^{\alpha}_{(i)}$, whose projection defines a second set of functions $\gamma^{\alpha}_{(i)}$
 $\beta^{*}\gamma^{\alpha}_{(i)}:=\frac{\phi^{\alpha}_{(i)}}{\rho^{\Lambda}}$ (50)
@@
 Figure 1: Commutative diagram showing how the two constraint algorithms run in parallel.
-Starting from the cotangent bundle, with its canonical symplectic form, the right-hand path shows how we may restrict to the primary constraint manifold $M_{0}$, and carry out the pre-symplectic algorithm, to reach the final space $\mathcal{P}$.
+Starting from the cotangent bundle, the right-hand path shows implementation of the pre-symplectic algorithm, to deduce the final space $\mathcal{P}$.
 This manifold is symplectic, and thus allows a contact reduction to be carried out.
-Alternatively, the scaling degree of freedom may be excised from the full symplectic phase space, defining the contact manifold $C$.
-Upon projecting the primary constraint functions that define $M_{0}$, $C$ is restricted to the pre-contact space $C_{0}$, and the left-hand path illustrates the implementation of the pre-contact algorithm.
+Alternatively, the scaling variable may be excised from the full symplectic phase space, defining the contact manifold $C$.
+Upon projecting the primary constraint functions that define $M_{0}$, $C$ is restricted to the pre-contact space $C_{0}$, and the left-hand path depicts the implementation of the pre-contact algorithm.
 Both paths lead to the same final reduced space $C_{\scriptscriptstyle\mathcal{P}}$.
@@
 $W=u^{2}\begin{pmatrix}1&1&0&-1\\ 1&1&0&-1\\ 0&0&4u^{-2}&0\\ -1&-1&0&2\end{pmatrix}$
-which is of constant rank 3, provided $u\neq 0$, which we have assumed. This Lagrangian admits a fairly simple scaling symmetry of the form
-$\Sigma=\frac{1}{2}\left(u\frac{\partial}{\partial u}+\dot{u}\frac{\partial}{\partial\dot{u}}\right)$ (63)
-Making the change of variable $u\mapsto e^{\rho/2}$, the scaling symmetry becomes $\Sigma=\partial_{\rho}$, while the Lagrangian adopts the form
+which is of constant rank 3, provided $u\neq 0$, which we have assumed. Making the change of variable $u\mapsto e^{\rho/2}$, the Lagrangian adopts the form
 $L=e^{\rho}\biggr[\frac{1}{2}\left(\dot{\rho}^{2}+\dot{\phi}^{2}\right)+\frac{1}{2}\left(\dot{x}+\dot{y}-\dot{\phi}\right)^{2}-\frac{1}{2}\left(x^{2}+y^{2}\right)+\phi^{2}\biggr]$ (52)
@@
 $M_{1}=\{p\in M_{0}\;|\;\psi^{2}(p):=e^{\rho}\left(x-y\right)=0\,\}$ (57)
-Any vector field solution $X_{H}$ to the geometrical equation $\iota_{X_{H}}\omega_{0}-dH_{0}=0$, restricted to $M_{1}$, must remain tangent to this surface, in order to be considered physically meaningful.
+Any vector field solution $X_{H}$ to the geometrical equation $\iota_{X_{H}}\omega_{0}=dH_{0}$, restricted to $M_{1}$, must remain tangent to this surface in order to be physically meaningful.
 As we have seen, this requirement is entirely equivalent to imposing that $\langle(dH_{0})_{p},T_{p}M_{1}^{\perp}\rangle=0$ for all points $p\in M_{1}$.
@@
 $J=e^{\rho}\begin{pmatrix}0&-2\\ 2&0\end{pmatrix}$
-Clearly, this is invertible, which indicates that $\psi^{1}$ and $\psi^{2}$ constitute a pair of second class constraints; it is therefore unnecessary to eliminate gauge degrees of freedom, and we henceforth refer to $M_{1}$ as $\mathcal{P}$.
+Clearly, this is invertible, indicating that $\psi^{1}$ and $\psi^{2}$ constitute a pair of second class constraints; it is therefore unnecessary to eliminate gauge degrees of freedom, and we henceforth refer to $M_{1}$ as $\mathcal{P}$.
 The inverse $J^{-1}$ is given by
@@
 $\{f,g\}_{D}=\{f,g\}-\frac{e^{-\rho}}{2}\{f,\psi^{1}\}\{\psi^{2},g\}+\frac{e^{-\rho}}{2}\{f,\psi^{2}\}\{\psi^{1},g\}$ (58)
-In order to deduce the dynamical evolution of phase space functions, we should construct the total Hamiltonian; however, having identified $\psi^{1}$ and $\psi^{2}$ as second class, it follows that the total Hamiltonian coincides with $H_{0}$.
-It is then straightforward to show that the equations of motion for our phase space variables, which follow from $\dot{f}=\{f,H_{0}\}_{D}|_{\scriptscriptstyle\mathcal{P}}$ are
+Since $\psi^{1}$ and $\psi^{2}$ are second class constraints, the total Hamiltonian coincides with $H_{0}$.
+The equations of motion for our phase space variables follow from $\dot{f}=\{f,H_{0}\}_{D}|_{\scriptscriptstyle\mathcal{P}}$, and we find
 $\dot{x}$ $=\frac{e^{-\rho}}{2}\left(2p_{x}+p_{\phi}\right)=\dot{y}$ $\dot{p}_{x}$ $=-e^{\rho}x=\dot{p}_{y}$
@@
 Whilst relatively simple in nature, the system considered has allowed us to present a fully-worked example of the commutativity of contact reduction and phase space restriction, and thus concludes our treatment of particle dynamics.
-In the following sections, we introduce the geometrical framework required to describe singular field theories, which will then allow us to provide a field-theoretic discussion, which parallels the ideas presented thus far.
 IX Multisymplectic Field Theory
 The use of multisymplectic geometry to describe classical field theories is an area of active research interest, as the fibred manifolds it employs provide an arena in which a manifestly covariant formalism may be developed, in a finite-dimensional setting [73, 74, 75].
-Scaling symmetries of classical field theories have been studied in [34], and a mathematical framework for the contact reduction of multisymplectic theories developed in [29].
-The multisymplectic Hamiltonian formalism of singular field theories is still not entirely understood; for example, it is known that there exist certain classes of Lagrangian, for which the construction of the corresponding Hamiltonian is either ambiguous, or simply ill-defined [16].
-Additionally, multisymplectic manifolds are not, in general, equipped with a well-defined bracket structure [86], and the local constraint algorithm employs geometrical principles, that render the use of such structures redundant [21, 22].
-As a result, in contrast to our study of systems of particles, our analysis of singular field theories will favour the Lagrangian formalism, and in what follows, we shall provide a heavily abridged summary of the most pertinent ideas, referring to [29] for more details.
+Scaling symmetries of classical field theories have been studied in [34], and a framework for the contact reduction of multisymplectic theories developed in [29].
+The multisymplectic Hamiltonian formalism of singular field theories is still not well understood; for example, there exist certain classes of Lagrangian, for which the construction of the corresponding Hamiltonian is either ambiguous, or simply ill-defined [16].
+Additionally, multisymplectic manifolds are not, in general, equipped with a well-defined bracket structure [76], and since the local constraint algorithm dispenses with such structures, our analysis of singular field theories will favour the Lagrangian formalism, rather than the Hamiltonian picture employed for particles [21, 22].
+In what follows, we provide a heavily abridged summary of the most pertinent ideas, referring to [29] for further details.
 IX.1 Lagrangian Field Theory
@@
 $\iota_{X_{p}}\Omega_{p}=0\quad\quad\implies\quad\quad X_{p}=0$
-If $\Omega$ is closed but 1-degenerate, we refer to the pair $(\mathcal{M},\Omega)$ as a pre-multisymplectic manifold; such objects will be of particular relevance in our study of singular theories.
-In order to formulate the field equations of a dynamical system in a geometrical fashion, we require a class of objects known as multivector fields.
+If $\Omega$ is closed but 1-degenerate, we refer to the pair $(\mathcal{M},\Omega)$ as a pre-multisymplectic manifold.
+The field equations of a dynamical system are expressed geometrically in terms of multivector fields.
 In general, a multivector field of degree $r$ on $\mathcal{M}$ is a section $\boldsymbol{X}\in\Gamma(\wedge^{r}T\mathcal{M})$ of the $r^{\textrm{th}}$ exterior power of the tangent bundle.
-We denote the space of all such multivector fields $\mathfrak{X}^{r}(\mathcal{M}):=\Gamma(\wedge^{r}\,T\mathcal{M})$, and declare some $\boldsymbol{X}\in\mathfrak{X}^{r}(\mathcal{M})$ to be locally decomposable if, for every point $p\in\mathcal{M}$, there exists an open neighbourhood $\mathcal{U}_{p}\subset\mathcal{M}$, and vector fields $X_{1},\,\cdots\,,X_{r}\in\mathfrak{X}^{\infty}(\mathcal{U}_{p})$, such that
+We denote the space of all such multivector fields $\mathfrak{X}^{r}(\mathcal{M}):=\Gamma(\wedge^{r}\,T\mathcal{M})$, and declare some $\boldsymbol{X}\in\mathfrak{X}^{r}(\mathcal{M})$ to be locally decomposable if, around point $p\in\mathcal{M}$, there exists an open neighbourhood $\mathcal{U}_{p}\subset\mathcal{M}$, and vector fields $X_{1},\,\cdots\,,X_{r}\in\mathfrak{X}^{\infty}(\mathcal{U}_{p})$, such that
 $\boldsymbol{X}|_{\mathcal{U}_{p}}=X_{1}\wedge\,\cdots\,\wedge X_{r}$
@@
 Consider a fibre bundle $\pi:E\rightarrow M$ over the $d$-dimensional spacetime manifold $M$; we shall suppose that $M$ is orientable, with volume form $\omega$.
-Local coordinates on $M$ are written as $(x^{\mu})$, with $\mu=0,\,\cdots,d-1$, so that $\omega=dx^{0}\wedge\,\cdots\wedge dx^{d-1}:=d^{d}x$.
-The $(n+d)$-dimensional manifold $E$ is referred to as the covariant configuration space, and the first jet bundle $\kappa:J^{1}E\rightarrow E$ of sections of $\pi$ is the natural space upon which to introduce a Lagrangian density [87, 88]; local coordinates on $J^{1}E$, adapted to the bundle structure of $E$, are given by $(x^{\mu},y^{a},y^{a}_{\mu})$, with $a=1,\,\cdots,n$.
+Local coordinates on $M$ are denoted $(x^{\mu})$, with $\mu=0,\,\cdots,d-1$, so that $\omega=dx^{0}\wedge\,\cdots\wedge dx^{d-1}:=d^{d}x$.
+The $(n+d)$-dimensional manifold $E$ is referred to as the covariant configuration space, and the first jet bundle $\kappa:J^{1}E\rightarrow E$ of sections of $\pi$ is the natural space upon which to introduce a Lagrangian density [77, 78]; local adapted coordinates on $J^{1}E$ are given by $(x^{\mu},y^{a},y^{a}_{\mu})$, with $a=1,\,\cdots,n$.
 Introducing the bundle projection $\widehat{\pi}:=\pi\circ\kappa:J^{1}E\rightarrow M$, the Lagrangian density $\mathcal{L}$ may be expressed as a $\widehat{\pi}-$semibasic $d$-form on $J^{1}E$
@@
 X Pre-multicontact Systems
-In our treatment of scaling symmetries of pre-symplectic systems, we observed that the elimination of a scaling degree of freedom and phase space restriction were commutative processes; indeed, we dedicated significant effort to the construction of the final symmetry-reduced theory via both paths of the commutative diagram 1.
-When considering field theories, we shall not attempt to retain this same level of generality, and so will restrict our attention to the implementation of a constraint algorithm only after the scaling degree of freedom has been eliminated.
-The symmetry reduction of multisymplectic field theories has been studied in detail in [29]; there, we presented a brief account of multicontact geometry, as pertained to the reduction of regular systems.
-When the theories of interest are singular, far more care must be taken, for it is not guaranteed that the symmetry-reduced form derived from the pre-multisymplectic $\Theta_{L}$ defines a pre-multicontact structure.
-Consequently, as an additional step, we must ensure that the reduced theory possesses a well-defined Reeb distribution; such considerations will require the introduction of additional mathematical structures, that were not required in [29].
-Our discussion of multicontact geometry shall be limited to those concepts which are fundamental to the construction of classical field theories; more details may be found in [26, 25, 27], for example.
+In our treatment of pre-symplectic systems, significant effort was dedicated to studying the commutative relationship between the elimination of a scaling degree of freedom and phase space restriction.
+When considering field theories, we shall not attempt to retain this level of generality, and so will restrict our attention to the implementation of a constraint algorithm only after the scaling degree of freedom has been eliminated.
 X.1 Multicontact Lagrangian Field Theory
@@
 The frictional behaviour of the original system is then palliated by the observation that any apparent losses may simply be attributed to an exchange of energy between the system and environment.
-The philosophical stance which motivates our study of scaling symmetries is diametrically opposed to the notion of creating larger, redundant mathematical structures, with the sole objective of streamlining practical calculations [77, 80, 90].
-We posit that the use of the smallest amount of dynamically-irrelevant structure, even at the expense of superficial mathematical simplicity, should be prioritised, wherever possible [91, 92].
-In excising superfluous scaling degrees of freedom, seeking a minimal ontology, we are forced to work with theories which are inherently frictional in nature; consequently, we shall dedicate this section to developing the geometrical tools required to describe such systems.
+The philosophical stance which motivates our study of scaling symmetries is diametrically opposed to the creation of larger redundant mathematical structures, with the sole objective of streamlining calculations [67, 70, 80].
+We posit that the use of the smallest amount of dynamically-irrelevant structure, even at the expense of superficial mathematical simplicity, should be prioritised, where possible [81, 82].
+In excising superfluous scaling degrees of freedom, seeking a minimal ontology, we are forced to work with theories which are inherently frictional in nature; consequently, we dedicate this section to developing the geometrical tools required to describe such systems.
 The fibred manifolds $\pi:E\rightarrow M$, and $\kappa:J^{1}E\rightarrow E$ (with $\textrm{dim}\,M=d$ and $\textrm{dim}\,E=n+d$) of the multisymplectic formalism continue to assume a prominent role in the description of frictional field theories; in particular, the Lagrangian density is now a $d$-form on the manifold
@@
 does not necessarily define a pre-multicontact structure.
-It is therefore necessary to introduce a number of additional concepts, which allow us to determine when a given Lagrangian $d$-form does in fact give rise to such a structure.
-We begin by considering a generalisation of the annihilator of a given subspace, that we first encountered in section (II); for the purpose of defining the pre-multicontact structures of this section, in addition to the Lagrangian $d$-form $\Theta_{L}$, we also introduce the notation $\Xi:=\beta^{*}\omega\in\Omega^{d}(\mathcal{S})$; this volume form serves as a reference object, while $\Theta_{L}$ endows the space $\mathcal{S}$ with the formal geometrical structure of a pre-multicontact manifold.
+It is therefore necessary that a number of additional criteria be met, which we now introduce.
+To facilitate our analysis, we introduce the notation $\Xi:=\beta^{*}\omega\in\Omega^{d}(\mathcal{S})$.
+While $\Theta_{L}$ endows $\mathcal{S}$ with the formal geometrical structure of a pre-multicontact manifold, $\Xi$ serves as a reference object.
 Given a regular distribution $\mathcal{D}\subset T\mathcal{S}$, we introduce the space of $r$-forms which annihilate all sections of $\mathcal{D}$
@@
 Sections of $\mathcal{D}^{R}$ are the Reeb vector fields; we denote this space $\mathfrak{R}:=\Gamma(\mathcal{D}^{R})$.
-In practice, we shall always work with coordinates on $\mathcal{S}$ adapted to the bundle structure $E\rightarrow M$, and so $\Xi=dx^{0}\wedge\,\cdots\,\wedge dx^{d-1}$ is a closed form, from which it follows that $\textrm{ker}\,\Xi$ consists of those vector fields whose vertical components $V^{\mu}\partial_{x^{\mu}}$ are vanishing.
+In practice, we shall always work in coordinates such that $\Xi=dx^{0}\wedge\,\cdots\,\wedge dx^{d-1}$, from which it follows that $\textrm{ker}\,\Xi$ consists of those vector fields whose vertical components $V^{\mu}\partial_{x^{\mu}}$ are vanishing.
 The space $\textrm{Ann}^{d}(\textrm{ker}\,\Xi)$ then comprises all $d$-forms $\xi\in\Omega^{d}(\mathcal{S})$, for which $\iota_{X}\xi=0$ for any choice of vector field of the form
@@
     $\textrm{Ann}^{d-1}(\textrm{ker}\,\Xi)=\{\iota_{R}\Theta_{L}\;|\;R\in\mathfrak{R}\,\}$
-Note that we have excluded the possibility that $k=0$, for this is precisely the condition required for $(\mathcal{S},\Theta_{L},\Xi)$ to be multicontact, and not pre-multicontact.
-In practice, when working in adapted coordinates, the local Reeb fields are deduced from $\iota_{R_{\mu}}\Theta_{L}=d^{d-1}x_{\mu}$; however, observe that this does not determine the $R_{\mu}$ uniquely, for it is always possible to add to $R_{\mu}$ an element of the characteristic distribution, that is to say $\mathfrak{R}=\textrm{span}\hskip 1.13809pt(R_{\mu})+\mathcal{C}$.
-Given an $r$-form $\xi\in\Omega^{r}(\mathcal{S})$, we introduce the following operator
-$\begin{split}\bar{d}\;:\;\Omega^{r}(\mathcal{S})\;&\longrightarrow\;\Omega^{r+1}(\mathcal{S})\\ \xi\;&\longmapsto\;\bar{d}\xi:=d\xi+\sigma_{\Theta_{L}}\wedge\,\xi\end{split}$
-where $\sigma_{\Theta_{L}}\in\Omega^{1}(\mathcal{S})$ is known as the dissipation form, and is expressed in local coordinates as
+We have excluded the possibility that $k=0$, as this is precisely the condition for $(\mathcal{S},\Theta_{L},\Xi)$ to be multicontact, and not pre-multicontact.
+In practice, when working in adapted coordinates, the local Reeb fields are deduced from $\iota_{R_{\mu}}\Theta_{L}=d^{d-1}x_{\mu}$; however, note that this does not determine the $R_{\mu}$ uniquely, for it is possible to add to $R_{\mu}$ an element of the characteristic distribution, that is to say $\mathfrak{R}=\textrm{span}\hskip 1.13809pt(R_{\mu})+\mathcal{C}$.
+To any action-dependent Lagrangian, we associate a dissipation form $\sigma_{\Theta_{L}}\in\Omega^{1}(\mathcal{S})$, expressed locally as
 $\sigma_{\Theta_{L}}=-\,\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}$ (79)
 The $(d+1)$-form $\Omega_{L}$ is then defined as
-$\Omega_{L}=\bar{d}\Theta_{L}=d\Theta_{L}-\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}\wedge\,\Theta_{L}$ (92)
+$\Omega_{L}:=d\Theta_{L}+\sigma_{\Theta_{L}}\wedge\Theta_{L}=d\Theta_{L}-\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}\wedge\,\Theta_{L}$ (80)
 Given a (pre-)multicontact Lagrangian system $(\mathcal{S},\Theta_{L},\Xi)$, the equations of motion for sections $\Psi\in\Gamma(M,\mathcal{S})$ are given by
@@
 which is clearly an element of $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$.
-Upon substituting this into the dynamical equations $\iota_{\boldsymbol{X}_{L}}\Theta_{L}=0$ and $\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$, we must ensure that the resulting expressions do not contain any inconsistencies.
-Supposing that we do in fact obtain a compatible set of equations, we now restrict $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$ to the subset of semi-holonomic solutions, which amounts to setting $F_{\mu}^{a}=y^{a}_{\mu}$ in the decomposition (96); in general, this step gives rise to further consistency conditions, and possibly to new constraints.
+Upon substituting this into the dynamical equations $\iota_{\boldsymbol{X}_{L}}\Theta_{L}=0$ and $\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$, we must ensure that the resulting expressions do not contain inconsistencies.
+Supposing that we obtain a compatible set of equations, we now restrict $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$ to the subset of semi-holonomic solutions, which amounts to setting $F_{\mu}^{a}=y^{a}_{\mu}$ in the decomposition (84); in general, this step gives rise to further consistency conditions, and possibly to new constraints.
 As was the case for the pre-symplectic algorithm, we must, at every stage, ensure that the dynamics of the system remain confined to the constraint submanifold; thus, for each local constraint function $\Phi$, if $\boldsymbol{X}_{L}$ is expressed as the product $\boldsymbol{X}_{L}=X_{0}\wedge\cdots\wedge X_{d-1}$, we must demand that $\mathfrak{L}_{X_{\mu}}\Phi=0$ for $\mu=0,\,\cdots,d-1$.
@@
 XI Contact Reduction of Singular Field Theories
-We shall now present the field-theoretic construction, which generalises the concepts developed in section (VII); as stated previously, we shall favour the Lagrangian formalism, as opposed to the Hamiltonian framework, as the latter suffers from significantly more pathological features, when considering singular field theories.
-Our exposition follows closely that of [29], and so we shall focus on features of the reduction process that are unique to singular theories.
-Referring to the notation developed throughout section (IX.1), we suppose that $(J^{1}E,\Omega_{L})$ is a pre-multisymplectic Lagrangian system, with corresponding Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$.
+Following closely the framework developed in [29], we now present a field-theoretic generalisation of the ideas of section (VII).
+Let us consider the pre-multisymplectic system $(J^{1}E,\Omega_{L})$, with corresponding Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$.
 From the Cartan $d$-form $\Theta_{L}$, we define
 $\theta_{L}^{\mu}:=-\,\iota_{\partial_{d-1}}\cdots\;\iota_{\partial_{0}}\left(\Theta_{L}\wedge dx^{\mu}\right)=\frac{\partial L}{\partial y^{a}_{\mu}}dy^{a}$ (85)
-The vector field $\Sigma\in\mathfrak{X}^{\infty}(J^{1}E)$ then constitutes a scaling symmetry of degree $\Lambda$ if
-$\mathfrak{L}_{\Sigma}L=\Lambda L\quad\quad\quad\textrm{and}\quad\quad\quad\mathfrak{L}_{\Sigma}\theta_{L}^{\mu}=\theta_{L}^{\mu}\quad\textrm{for all}\;\mu=0,\,\cdots,d-1$
+The vector field $\Sigma\in\mathfrak{X}^{\infty}(J^{1}E)$ is said to constitute a scaling symmetry of degree $\Lambda$ if
+$\mathfrak{L}_{\Sigma}L=\Lambda L\quad\quad\quad\textrm{and}\quad\quad\quad\mathfrak{L}_{\Sigma}\theta_{L}^{\mu}=\theta_{L}^{\mu}\quad\textrm{for all}\;\mu=0,\,\cdots,d-1$ (86)
 Consider a Herglotz (i.e multicontact) Lagrangian $L^{H}$, embedded within a multisymplectic manifold of one dimension higher through the expression
@@
 $\rho_{\mu}=-\,\frac{\partial L^{H}}{\partial s^{\mu}}$ (88)
-It is relatively straightforward to show that the equations of motion of $L$ directly imply the Herglotz-Lagrange field equations for $L^{H}$, when the former is restricted to the subspace upon which $L^{H}$ is defined.
+It is straightforward to verify that the equations of motion derived from $L$ directly imply the Herglotz-Lagrange field equations for $L^{H}$, when the former is restricted to the subspace upon which $L^{H}$ is defined.
 In light of this, suppose that, having identified a scaling symmetry $\Sigma$ of our pre-multisymplectic system, we adopt coordinates on $J^{1}E$ in such a way so as to render this vector field of the form
@@
 $L=e^{\rho}f(\rho_{\mu},\psi^{a},\psi^{a}_{\mu})$ (90)
-The Euler-Lagrange field equation for $\rho$ then implies that the function $f$ may be written as
+for some function $f$. The Euler-Lagrange equation for $\rho$ then implies that $f$ may be written as
 $f=\rho_{\mu}\frac{\partial f}{\partial\rho_{\mu}}+\frac{\partial}{\partial x^{\mu}}\frac{\partial f}{\partial\rho_{\mu}}$
-which, when compared to (98), and recalling that the Herglotz-Lagrange equations require that $\partial_{\mu}s^{\mu}=L^{H}$, suggests we should identify
+which, when compared to (87), and recalling that $\partial_{\mu}s^{\mu}=L^{H}$, suggests we should identify
 $s^{\mu}=\frac{\partial f}{\partial\rho_{\mu}}\quad\quad\quad\quad L^{H}=f-\rho_{\mu}s^{\mu}$ (91)
-All steps outlined above are applicable to regular Lagrangian systems; however, the novelty that arises in the singular case, is that the Lagrangian $d$-form $\Theta_{L^{H}}$ calculated from the symmetry-reduced function $L^{H}$ is not guaranteed to be pre-multicontact.
-Consequently, as a crucial additional step, we must compute the characteristic and Reeb distributions, after making the ‘naïve’ contact reduction; if these objects do not satisfy the conditions given in section (X), the symmetry reduction procedure fails.
+All steps outlined above are applicable to regular systems; the novelty that arises in the singular case, is that the $d$-form $\Theta_{L^{H}}$ calculated from the symmetry-reduced function $L^{H}$ is not guaranteed to be pre-multicontact.
+Consequently, we must compute the characteristic and Reeb distributions, after making the ‘naïve’ contact reduction; if these do not satisfy the conditions given in section (X), the symmetry reduction procedure fails.
 Of course, we shall only consider those physically interesting cases, in which the reduced space does inherit a pre-multicontact structure; nevertheless, it is a non-trivial condition that must be verified.
-Having calculated $\Theta_{L^{H}}$, and confirmed that $\textrm{rank}\,\mathcal{D}^{R}=d+k$ and $\textrm{rank}\,\mathcal{C}=k$, for some $0<k\leqslant n(1+d)$, we have a pre-multicontact manifold $(\mathcal{S}_{0},\Theta_{L^{H}})$, in which $\mathcal{S}_{0}$ denotes the reduced space, obtained as a result of excising the scaling variable $\rho$.
-Geometrically speaking, if the system of interest is such that we may globally choose a single covariant configuration space variable as a scaling function: that is to say, if the quantity $\xi$ in (100) corresponds to a single field variable on $E$, then, upon writing the Lagrangian in the form (101), $E$ separates into two connected pieces $E_{\pm}$, where each of $E_{\pm}$ is the product of a trivial bundle, and a codimension-1 subspace: $E_{\pm}\cong\widetilde{E}\times_{M}(M\times\mathbb{R}_{\pm})$.
-The two components $E_{\pm}$ arise as a result of the change of variable $\xi\rightarrow e^{\rho/\Lambda}$, which selects only the positive component of $\xi$.
-In order to cover the full range of $\xi$, we must also consider a change of variables in which $\xi\rightarrow-\,e^{\rho/\Lambda}$.
-Within each of $E_{\pm}$, it is precisely the trivial bundle $M\times\mathbb{R}_{\pm}\rightarrow M$ that is eliminated by the contact reduction.
-Thus, provided we consider both components, so as not to lose any dynamical information, the reduced space is isomorphic to $J^{1}\widetilde{E}\times\mathbb{R}^{d}$, in which the codimension-1 subspace $\widetilde{E}$ may intuitively be identified with a covariant configuration space comprised of all original field variables, except $\rho$.
-The reduced space does not simply inherit the structure of the jet bundle $J^{1}\widetilde{E}$, since the field variable $\rho$ has been excised, but the velocities $\rho_{\mu}$ have not; indeed these $d$ independent parameters span a copy of $\mathbb{R}^{d}$, and the Cartan form $\Theta_{L^{H}}$ determines whether this defines a Reeb distribution.
-Supposing that $(\mathcal{S}_{0},\Theta_{L^{H}})$ does constitute a pre-multicontact system, with characteristic distribution $\mathcal{C}$, we calculate the $(d+1)$-form $\Omega_{L^{H}}:=\bar{d}\Theta_{L^{H}}$, and introduce a multivector field $\boldsymbol{X}_{L^{H}}\in\mathfrak{X}^{d}(\mathcal{S}_{0})$, with local decomposition
+Having calculated $\Theta_{L^{H}}$, and confirmed that $\textrm{rank}\,\mathcal{D}^{R}=d+k$ and $\textrm{rank}\,\mathcal{C}=k$, for some $0<k\leqslant n(1+d)$, we have a pre-multicontact manifold $(\mathcal{S}_{0},\Theta_{L^{H}})$, in which $\mathcal{S}_{0}$ denotes the reduced space, obtained upon excising the scaling variable $\rho$.
+Geometrically, if our system is such that the quantity $\xi$ in (89) corresponds to a single field variable on $E$, then, upon writing the Lagrangian in the form (90), $E$ separates into two connected pieces $E_{\pm}$, where each of $E_{\pm}$ is the product of a trivial bundle, and a codimension-1 subspace: $E_{\pm}\cong\widetilde{E}\times_{M}(M\times\mathbb{R}_{\pm})$.
+The change of variable $\xi\rightarrow e^{\rho/\Lambda}$ selects only the positive component of $\xi$; in order to cover the full range of $\xi$, we must also consider a change of variables in which $\xi\rightarrow-\,e^{\rho/\Lambda}$.
+Within each of $E_{\pm}$, it is precisely the trivial bundle that is eliminated by the contact reduction.
+Thus, provided we consider both components, so as not to lose any dynamical information, the reduced space is isomorphic to $J^{1}\widetilde{E}\times\mathbb{R}^{d}$, in which the codimension-1 subspace $\widetilde{E}$ may be identified with a configuration space comprised of all original field variables, except $\rho$.
+The reduced space does not simply inherit the structure of the jet bundle $J^{1}\widetilde{E}$, as the velocities $\rho_{\mu}$ have not been eliminated; indeed these coordinates span a copy of $\mathbb{R}^{d}$, and the Cartan form $\Theta_{L^{H}}$ determines whether this defines a Reeb distribution.
+Supposing that $(\mathcal{S}_{0},\Theta_{L^{H}})$ does constitute a pre-multicontact system, we calculate the $(d+1)$-form $\Omega_{L^{H}}$, and introduce a multivector field $\boldsymbol{X}_{L^{H}}\in\mathfrak{X}^{d}(\mathcal{S}_{0})$, with local decomposition
 $\boldsymbol{X}_{L^{H}}=\bigwedge\limits_{\mu=0}^{d-1}\left(\frac{\partial}{\partial x^{\mu}}+F^{a}_{\mu}\frac{\partial}{\partial\psi^{a}}+G^{a}_{\mu\nu}\frac{\partial}{\partial\psi^{a}_{\nu}}+K^{\nu}_{\mu}\frac{\partial}{\partial s^{\nu}}\right)$
@@
 The appearance of scalar fields that directly couple to gauge curvature terms is a phenomenon which arises most notably in string-inspired models [84, 85].
-Indeed, the imprint of dilaton fields on low-energy effective actions has speculatively been regarded as a means to offer novel insight into problems such as colour confinement [96, 97].
-In this section, we present an example of a non-Abelian gauge theory coupled to a dilaton-like scalar field; our goal is to illustrate how the concepts of previous sections are applied in practice, as well as to discuss the physical interpretation of our results.
+Indeed, the imprint of dilaton-like fields on low-energy effective actions has speculatively been regarded as a means to offer novel insight into problems such as colour confinement [86, 87].
+In this section, we present an example of a non-Abelian gauge theory coupled to a scalar field.
+Having implemented the reduction procedure, we discuss the physical implications of our results.
 XII.1 Geometrical Setting
-Let $\pi:P\rightarrow M$ be a principle $G$-bundle over the $d$-dimensional spacetime manifold $M$; on $P$, there exists a unique, Lie algebra valued connection $1$-form $\omega\in\Omega^{1}(P,\mathcal{L}(G))$, which descends to a local form $A:=s^{*}\omega\in\Omega^{1}(U,\mathcal{L}(G))$, in which $s:U\rightarrow P$ is a local trivialising section of $P$, defined on the open subset $U\subset M$.
 We shall suppose that the spacetime manifold $M$ is equipped with a Lorentzian metric $g$ of signature $(+,-,-,\,\cdots\,)$; this metric is considered to be parametric, and non-variational [19], so that, physically, our theory is defined on a curved background, which we may change, but is not coupled to gravity.
@@
 $L=\textrm{Tr}\left[-\frac{\phi^{2}}{2g^{2}}F_{\mu\nu}F^{\mu\nu}+2J_{\mu}A^{\mu}\right]+\frac{1}{2}g^{\mu\nu}\phi_{\mu}\phi_{\nu}-V(\phi)$ (94)
-where the trace is taken over the indices of $\mathcal{L}(G)$, and $F$ refers to the Lie algebra valued curvature 2-form $F:=dA+\frac{1}{2}[A,A]$.
+where the trace is taken over the indices of $\mathcal{L}(G)$, and $F$ refers to the Lie algebra valued curvature 2-form.
 In what follows, we shall adopt group-theoretic conventions more prevalent in the physics literature, taking the generators $T_{a}$ of the Lie algebra to be Hermitian, with $T_{a},T_{b}]=if_{ab}^{\;\;\;c}\,T_{c}$.
-Supposing that $G$ is a semi-simple compact Lie group, the Hermiticity of the $T_{a}$ implies that the Killing form is both non-degenerate, and positive semi-definite; in the fundamental representation, we adopt the following normalisation with respect to the trace
+In the fundamental representation, we adopt the following normalisation with respect to the trace
 $\textrm{Tr}\left(T_{a}T_{b}\right)=\frac{1}{2}\delta_{ab}$ (95)
@@
 For simplicity of exposition, we shall take $V(\phi)$ to be a single mass term $\frac{1}{2}m^{2}\phi^{2}$, and temporarily ignore the source; having carried out the contact reduction, we will briefly comment on the implications of reinstating $J_{\mu}^{a}$.
-We remark that, with the above conventions, we have implicitly assumed that the dilaton transforms trivially under $G$; were $\phi$ to transform in some non-trivial representation $\rho:G\rightarrow GL(V)$, we would instead have to consider the associated bundle $P\times_{\rho}V$, with the result that our covariant configuration space $\mathcal{E}$ would instead be
-$\mathcal{E}=C(P)\times_{M}\left(P\times_{\rho}V\right)\times_{M}\textrm{Sym}_{2}^{1,d-1}(M)$
-where the trivial bundle $M\times\mathbb{R}$ has been replaced with the vector bundle $P\times_{\rho}V$ associated to $P$ via the representation $\rho$.
+Note that we have implicitly assumed that the dilaton transforms trivially under $G$; were $\phi$ to transform in some non-trivial representation $\rho:G\rightarrow GL(V)$, the trivial bundle $M\times\mathbb{R}$ would be replaced with the vector bundle $P\times_{\rho}V$ associated to $P$ via the representation $\rho$.
 In the interest of limiting the number of additional complications, we shall content ourselves with a dilaton that transforms trivially.
@@
 $L=-\,\frac{\phi^{2}}{4g^{2}}F^{a}_{\mu\nu}F_{a}^{\mu\nu}+\frac{1}{2}g^{\mu\nu}\phi_{\mu}\phi_{\nu}-\frac{1}{2}m^{2}\phi^{2}$ (96)
-in which the curvature $F^{a}_{\mu\nu}$ has the following local coordinate expression
-$F_{\mu\nu}^{a}=A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f_{bc}^{\;\;\;a}A^{b}_{\mu}A^{c}_{\nu}$
+in which the curvature is expressed locally as $F_{\mu\nu}^{a}=A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f_{bc}^{\;\;\;a}A^{b}_{\mu}A^{c}_{\nu}$.
 The degeneracy of this Lagrangian arises as a result of the gauge symmetry; indeed, the Hessian matrix with respect to $A^{a}_{\mu,\nu}$ and $\phi_{\mu}$ has the following entries
@@
 XII.2 Contact Reduction
-Upon making the change of variables $\phi\mapsto e^{\rho/2}$, the scaling symmetry vector field $\Sigma$ is now simply $\partial_{\rho}$, and the Lagrangian, with the curvature term expanded in full, reads
-$\begin{split}L=-\,\frac{e^{\rho}}{4g^{2}}\biggr[g^{\mu\lambda}g^{\nu\kappa}\left(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\right)\left(A^{a}_{\kappa,\lambda}-A^{a}_{\lambda,\kappa}+f_{de}^{\;\;\;a}A^{d}_{\lambda}A^{e}_{\kappa}\right)\biggr]\\ +e^{\rho}\left(\frac{1}{8}g^{\mu\nu}\rho_{\mu}\rho_{\nu}-\frac{1}{2}m^{2}\right)\end{split}$ (110)
-Recall from section (XI) that when the Lagrangian is expressed in the form $L=e^{\rho}f(\rho_{\mu},A_{\mu},A_{\mu,\nu})$, the action density is found from $s^{\mu}=\partial f/\partial\rho_{\mu}$, whilst the Herglotz Lagrangian takes the form $L^{H}=f-\rho_{\mu}s^{\mu}$; for the former, we find
+Upon making the change of variables $\phi\mapsto e^{\rho/2}$, the scaling symmetry vector field $\Sigma$ is now simply $\partial_{\rho}$, and the Lagrangia reads
+$\begin{split}L=-\,\frac{e^{\rho}}{4g^{2}}F_{\mu\nu}^{a}F^{\mu\nu}_{a}+e^{\rho}\left(\frac{1}{8}g^{\mu\nu}\rho_{\mu}\rho_{\nu}-\frac{1}{2}m^{2}\right)\end{split}$ (99)
+Recall that when the Lagrangian is expressed in the form $L=e^{\rho}f(\rho_{\mu},A_{\mu},A_{\mu,\nu})$, the action density is found from $s^{\mu}=\partial f/\partial\rho_{\mu}$, whilst the Herglotz Lagrangian takes the form $L^{H}=f-\rho_{\mu}s^{\mu}$; for the former, we find
 $s^{\mu}:=\frac{\partial f}{\partial\rho_{\mu}}=\frac{1}{4}g^{\mu\nu}\rho_{\nu}\quad\quad\implies\quad\quad\rho_{\mu}=4g_{\mu\nu}s^{\nu}$ (100)
-In anticipation of the elimination of $\rho_{\mu}$ in favour of the action density, we have expressed the former in terms of the latter. The Herglotz Lagrangian for this system is then
-$L^{H}=-\,\frac{1}{4g^{2}}\biggr[g^{\mu\lambda}g^{\nu\kappa}\left(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\right)\left(A^{a}_{\kappa,\lambda}-A^{a}_{\lambda,\kappa}+f^{\;\;\;a}_{de}A^{d}_{\lambda}A^{e}_{\kappa}\right)\biggr]-\frac{1}{2}m^{2}-2g_{\mu\nu}s^{\mu}s^{\nu}$ (112)
-Due to the singular nature of the Lagrangian (107) from which $L^{H}$ is constructed, it is not a priori guaranteed that the $d$-form $\Theta_{L^{H}}$ endows the reduced space with a pre-multicontact structure.
-Consequently, prior to implementing the constraint algorithm, we must compute the characteristic and Reeb distributions, verifying that they do indeed satisfy the conditions required for $\Theta_{L^{H}}$ to be pre-multicontact.
+while the Herglotz Lagrangian for this system is
+$L^{H}=-\,\frac{1}{4g^{2}}F_{\mu\nu}^{a}F^{\mu\nu}_{a}-\frac{1}{2}m^{2}-2g_{\mu\nu}s^{\mu}s^{\nu}$ (101)
 Geometrically, referring to the decomposition (92), we see that the excision of the dilaton corresponds to the removal of the trivial bundle $M\times\mathbb{R}\rightarrow M$, defining a reduced covariant configuration space $\mathcal{E}_{\textrm{red}}$.
 More formally, since $\phi$ is a globally-defined scaling function, the symplectification $\widetilde{\mathcal{E}}_{\textrm{red}}$ is a trivial $\mathbb{R}^{\times}$-bundle over $M$, composed of two connected components $\widetilde{\mathcal{E}}_{\textrm{red}}^{\pm}$, which, with the change of variables $\phi\rightarrow e^{\rho/2}$, correspond to symplectification via $\pm e^{\rho/2}$.
-Consequently, with the understanding that care must be taken to ensure that both components are considered, we may somewhat informally speak of a reduced space $\mathcal{E}_{\textrm{red}}$ that is isomorphic to $C(P)\times_{M}\textrm{Sym}^{1,d-1}_{2}(M)$.
+Provided that both components are considered, we may somewhat informally write $\mathcal{E}_{\textrm{red}}\cong C(P)\times_{M}\textrm{Sym}^{1,d-1}_{2}(M)$.
 The Herglotz Lagrangian is defined on the space
@@
 Note that the final quantity on the right-hand side is precisely $L^{H}$, and the bracket $\left(A^{a}_{\nu,\mu}-C^{a}_{\mu\nu}\right)$ vanishes upon imposing semi-holonomy.
-The contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$ is somewhat involved, and the lengthy algebraic manipulations required are not particularly illuminating; for completeness, we provide the details in appendix A.
-Performing the contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$, and setting the result to zero, we obtain the following pair of expressions
+The contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$ produces a pair of expressions, both of which must vanish separately; we provide the details of this calculation in appendix A
 $0$ $=(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu})-(C^{a}_{\mu\nu}-C^{a}_{\nu\mu})$ (108)
@@
 $\begin{split}\partial_{\mu}s^{\mu}=L^{H}\\ \partial_{\mu}F^{a\mu\nu}+f_{bc}^{\;\;\;a}A^{b}_{\mu}F^{c\mu\nu}+4g_{\mu\lambda}s^{\lambda}F^{a\mu\nu}=0\end{split}$ (113)
-The first two terms of the second expression form the familiar covariant derivative $(\mathcal{D}_{\mu}F^{\mu\nu})^{a}$, whilst the final term is a common feature of contact-reduced theories, that arises as a result of eliminating a scaling degree of freedom from our ontology.
-More concretely, in the original Lagrangian (105), there was a coupling between the scalar field $\phi$ and the gauge-kinetic term $\textrm{Tr}\,(F_{\mu\nu}F^{\mu\nu})$; the removal of the dilaton, in favour of a minimal ontology, comes at the cost of introducing a dissipative characteristic into the theory, which is made manifest through the coupling of the gauge field strength to the action density.
 Before concluding, there remains an important matter to be discussed, concerning the physical interpretation of the excision of the scalar field $\phi$.
@@
 However, this is not the case; consider, for instance, an experiment in which one attempts to deduce the expectation value $\langle\phi\rangle$, which, for the current discussion, we shall suppose to fix some coupling parameter $g_{s}$.
-The value of $g_{s}$ is not something directly accessible to an observer, but rather is deduced via comparison to a reference object; this is entirely analogous to how an object of length 1 m is only known to be such because its size is in a 1:1 ratio with that of a previously-standardised instrument.
-Consequently, under changes generated by $\phi$, the measuring apparatus is subjected to precisely the same rescaling as $g_{s}$, with the result that the overall scaling is empirically inaccessible.
-Thus, we are justified in asserting that this degree of freedom is redundant, from the perspective of the algebra of dynamical observables.
+The value of $g_{s}$ is not directly accessible, but must be deduced via comparison to a reference object; this is entirely analogous to how an object of length 1 m is only known to be such because its size is in a 1:1 ratio with that of a previously-standardised instrument.
+Consequently, changes in $\langle\phi\rangle$ are accompanied by simultaneous rescalings of the sensitivity of our apparatus, with which we would like to ascertain the value of $g_{s}$.
+It thus follows that the object responsible for these empirically-inaccessible changes in $g_{s}$ (i.e the field $\phi$) is wholly redundant, from the perspective of the physical dynamics.
 In fact, we make a somewhat bolder claim: if one is able to construct a theory whose empirically-accessible degrees of freedom are invariant under a particular transformation, we assert that it is a necessary condition, for any relational description, that the mathematical artifice generating such transformations be excisable.
@@
 For instance, all lengths are deduced with respect to a standard metre stick, and similarly, all times are measured as multiples of the standardised ‘tick’ of a clock.
-Suppose that we ourselves are extrinsic ($i.e$ external) observers to $\mathcal{S}$, and we note that, under a certain transformation $\mathcal{T}$, the set of variables we use to describe $\mathcal{S}$ suffer a rescaling.
+Suppose that we ourselves are external observers to $\mathcal{S}$, and we note that, under a certain transformation $\mathcal{T}$, the set of variables we use to describe $\mathcal{S}$ suffer a rescaling.
 However, this rescaling is such that the dynamical equations we deduce remain invariant.
 As a trivial example, we consider a pair of variables $(x,t)$, with $\mathcal{T}:(x,t)\mapsto(\lambda x,\lambda t)$; the velocity $\dot{x}$ remains invariant, even though the variables used to deduce $\dot{x}$ do not.
-The generator of the transformation $\mathcal{T}$ must be a redundant component of our description; if it were redundant, but non-excisable, then it would necessarily appear in the ontology of the observer, else the intrinsic and extrinsic theories would not describe the same system.
+The generator of the transformation $\mathcal{T}$ is clearly a redundant part of our description, and must therefore be excisable.
+If this were not the case, then it would necessarily appear in the ontology of the observer, else the intrinsic and extrinsic theories would not describe the same system.
 This contradicts our assumption that the observer’s description is relational.
@@
 Many of the most mathematically-rich models in contemporary theoretical physics possess, as a central component, gauge symmetries; such theories are described by singular Lagrangians, and over the course of this article, we have provided a basic introduction to gauge theories, from the perspective of geometric mechanics.
-For systems of particles, the appropriate formalism was that of pre-symplectic geometry, while a field-theoretic discussion required the introduction of the fibred manifolds of pre-multisymplectic geometry.
-We have demonstrated that singular theories which exhibit invariance under global scaling transformations may be subjected to a reduction process similar to that which has been studied for regular systems, provided that degeneracies are controlled, through use of a suitable constraint algorithm.
-We observed that this constraint procedure could be implemented pre- or post-reduction; from a physical perspective, this is to be expected: if the processes of contact reduction and phase space restriction were non-commutative, this would imply that the constraint functions had, in some sense, information about the global scaling variable.
-Had the reduction of a constrained system led to a theory distinct from that obtained via the constraint of a reduced theory, then any system described by a degenerate Lagrangian would offer a means to detect changes in global scale, which is a highly unphysical outcome.
-Having described the reduction of singular field theories, we then provided an example of an effective non-Abelian gauge theory coupled to a dilaton-like field; it was observed that this real scalar field was in fact redundant, and could therefore be excised.
-The elimination of this degree of freedom was, at first sight, relatively alarming, particularly because the coupling parameters of a theory are often fixed by the expectation value of such scalar fields.
-However, we argued that the strength of an interaction is something that must necessarily be deduced via empirical methods; as such, any hypothetical experiment, conducted with the aim of measuring such parameters will be insensitive to a rescaling of their values, since the apparatus itself will suffer an identical change in scale.
+We have demonstrated that singular theories which exhibit invariance under global scaling transformations may be subjected to a reduction process similar to that which has been studied for regular systems.
+Degeneracies must be controlled, and we found that the constraint procedure required to do so could be implemented pre- or post-reduction.
+From a physical perspective, this is unsurprising: had the processes of contact reduction and phase space restriction been non-commutative, this would imply that the constraint functions had, in some sense, information about the global scaling variable.
+If the reduction of a constrained system led to a theory distinct from that obtained via the constraint of a reduced theory, then any system described by a degenerate Lagrangian would offer a means to detect changes in global scale, which is a highly unphysical outcome.
+Having described the reduction of singular field theories, we provided an example of an effective non-Abelian gauge theory coupled to a dilaton-like field; it was observed that this real scalar field was in fact redundant, and could therefore be excised.
+The elimination of this degree of freedom was, at first sight, relatively alarming, particularly because the coupling parameters of a theory are often fixed by the expectation values of such scalar fields.
+However, we argued that the strength of an interaction is something that is necessarily deduced via empirical methods; as such, any hypothetical experiment, conducted with the aim of measuring such parameters will be insensitive to a rescaling of their values, since the apparatus itself will suffer an identical change in scale.
 The results presented throughout constitute a framework within which to analyse singular theories; it is known that, upon making a conformal decomposition of the spacetime metric, the Einstein-Hilbert Lagrangian possesses a scaling symmetry, corresponding precisely to the conformal factor [34].
@@
 From the results of the present work, we know that it is possible to implement the geometrical constraint algorithm after having eliminated the scaling degree of freedom.
-As such, it will be of great interest to examine any simplifications that arise, as a result of working with a reduced ontology, in which we retain only those degrees of freedom required to close the algebra of dynamical observables.
-In general, we distinguish between two principal approaches to the treatment of classical field theory.
-On the one hand, as is customary when attempting to construct a quantum analogue of the theory, one may consider the temporal evolution of the initial data, defined on a particular choice of Cauchy slice [101].
-This ‘canonical’ formulation studies spatially-dependent fields, defined at fixed instants of time, and has the notably inconvenient features of possessing an infinite-dimensional phase space, and sacrificing manifest Lorentz covariance [102].
-The ADM formalism of General Relativity is a constrained Hamiltonian theory that adopts just such an approach, considering the (globally-hyperbolic) spacetime manifold to be foliated by spacelike hypersurfaces [103, 104].
-The Hamilton De-Donder Weyl formalism we have employed throughout makes no distinction between the spacetime coordinates, which immediately guarantees the Lorentz covariance that was absent in the canonical formulation.
-Having established a framework within which to treat constrained theories, it will be highly illustrative to contrast the constraint of a contact-reduced action for General Relativity with the more familiar canonical approach.
+It will be of great interest, therefore, to examine any simplifications or novelties that arise as a result of working with a reduced ontology, particularly at those points where the standard Einstein-Hilbert action provides ill-defined solutions, where scales become singular.
+Containing no reference to such scales, our contact-reduced action will not suffer the same pathologies, and will therefore provide an arena in which to explore these singular points.
+Finally, more work is required to ascertain how one should construct a field-theoretic generalisation of the promotion of coupling parameters to dynamical variables, discussed at the end of section (XII).
+Having found such a prescription, it will, in principle, be possible to analyse a much broader class of gauge theories, such as the Standard Model.
 Appendix A Multivector Field Manipulations for Non-Abelian Gauge Theory